### v3

# Embed-Search-Align: DNA Sequence Alignment using Transformer models

Abstract
DNA sequence alignment involves assigning short DNA reads to the most probable locations on an extensive reference genome. This process is crucial for various genomic analyses, including variant calling, transcriptomics, and epigenomics. Conventional methods, refined over decades, tackle this challenge in two steps: genome indexing followed by efficient search to locate likely positions for given reads. Building on the success of Large Language Models (LLM) in encoding text into embeddings, where the distance metric captures semantic similarity, recent efforts have explored whether the same Transformer architecture can produce numerical representations for DNA sequences. Such models have shown early promise in tasks involving classification of short DNA sequences, such as the detection of coding- vs non-coding regions, as well as the identification of enhancer and promoter sequences. Performance at sequence classification tasks does not, however, translate to sequence alignment, where it is necessary to conduct a genome-wide search to successfully align every read. We address this open problem by framing it as an “Embed-Search-Align” task. In this framework, a novel encoder model DNA-ESA generates representations of reads and fragments of the reference, which are projected into a shared vector space where the read-fragment distance is used as a surrogate for alignment. In particular, DNA-ESA introduces: (1) Contrastive loss for self-supervised training of DNA sequence representations, facilitating rich sequence-level embeddings, and (2) a DNA vector store to enable search across fragments on a global scale. DNA-ESA is $>97\%$ accurate when aligning $250$-length reads onto a human reference genome of 3 gigabases (single-haploid), far exceeds the performance of $6$ recent DNA-Transformer model baselines and shows task transfer across chromosomes and species.
1 Background
The aim of this paper is to establish a foundation model tailored for DNA sequences, where the vocabulary consists of only a few symbols ($\{A,T,G,C\}$ in this case). Numerous DNA Transformer models (Ji et al., 2021; Zvyagin et al., 2022; Fishman et al., 2023; Dalla-Torre et al., 2023) have emerged recently, mainly designed for classification tasks in downstream applications. However, these models do not explicitly consider a fundamental distinction between Limited Vocabulary Languages (LVL) (such as, genomes) and natural languages with large vocabulary. In LVLs, there is a one-to-one correspondence between the precise symbol ordering and the underlying “meaning”. If a protein is encoded by one amino acid sequence, a slightly different sequence (with a few edits) would encode for a different protein with a distinct functionality in the cells. Existing foundation Transformer models generate sequence embeddings such that their pairwise distances correspond to class separation, thus sequences with very large edit distances end up with representations that are close by. There are however tasks such as Sequence Alignment where the pairwise representation distance has to closely match the sequence edit distance. Indeed as portrayed in Figure 2, current DNA Transformer models fail to perform Sequence Alignment task. This requires a new kind of foundation model which is able to transfer a precise and computationally expensive distance metric over LVLs to a representation space.
The simplest sequence alignment task applies to single-end [^1] A DNA fragment is ligated to an adapter and then sequenced from one end only. reads. Given a reference sequence :=$\mathcal{R}:=\{b_{1},b_{2},\dots,b_{N}\}$ – for the single-haploid human genome (Nurk et al., 2022), $N\approx 3$ gigabases (gb) – the primary objective is to identify the most probable start- and end-positions within this reference for a short DNA read,
[…]
which may contain mutations due to base insertions, deletions, and substitutions. Computational simulators have been developed to generate synthetic reads that have properties of real reads. These simulators mimic the read quality and characteristics produced by actual sequencing machines, thus providing a scalable means for validating new alignment approaches (Huang et al., 2012).
(a) A $20,000$-long nucleotide sequence from Chr. 1, 2, and 3 each is randomly sampled and broken up into $100$ consecutive reference fragments (||||) each of length $1000$ and stride $200$. For each fragment, we sample a read (–) of 100-250 consecutive base pairs. Each sequence is encoded using DNA-ESA and the embeddings are visualized using 2D UMAP (McInnes et al., 2018). The colors identify chromosomes and the position of the fragment is coded by its intensity.
(b) Fragment (||||) and read (–) sequences from five gene regions (listed above in the inset) are similarly projected into a shared latent space. Note that for UMAP projections (both subplots (a) and (b)), the consecutive fragments belonging to the same nucleotide sequence constitute an order-preserving 1D manifold. A successfully aligned read is observed to be co-located with its corresponding fragment in the embedding space while a read further away is observed to be misclassified.
Figure 1: Illustrating DNA-ESA’s Preservation of Sequence Locality in Embedding Space: In DNA-ESA, the reference genome $\mathcal{R}$ is divided into fragments $\mathcal{F}_{i}$, each represented by an embedding $h(\mathcal{F}_{i})$. For effective sequence alignment, specific structures are expected in the embedding space: (1) Overlapping fragments $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$ should have proximate embeddings; (2) Consecutive fragments forming a long sequence should correspond to a distinct manifold in the embedding space. Subfigures (a) and (b) display this emergent geometry, as visualized using DNA-ESA.
Conventional sequence alignment methods, such as those used in BWA-MEM, have evolved significantly from the basic Smith-Waterman (SW) approach, which computes similarity between a read $r$ and reference $\mathcal{R}$, but is computationally intensive for large genomes. Key advancements include: (a) Sharding the reference into smaller fragments for individual searching; (b) Progressive search using phylogenetic trees and distance heuristics for logarithmic complexity scaling; (c) Compression techniques like the Burrows-Wheeler transform (Li & Durbin, 2009), reducing search length; and (d) Multi-core/thread implementations and database instantiations for faster computations and improved data recall (Li, 2018; Vasimuddin et al., 2019; Langmead et al., 2018).
In this work we explore an alternative paradigm for aligning a read to a genome, drawing parallels with advancements in Natural Language Processing (NLP). Traditional NLP relied on rule-based techniques to diagram sentences into grammatical components, using dependency parse trees to discern semantic relationships. However, the introduction the introduction of models like the Transformers (Vaswani et al., 2017) revolutionized this approach. Unlike rule-based frameworks, Transformers implicitly identify syntactic and semantic structures, as evidenced in various NLP tasks like Sentiment Analysis, Entity Recognition, and Question-Answering (Devlin et al., 2019; Reimers & Gurevych, 2019; Qiu et al., 2020). The flexibility of this architecture has facilitated its application beyond language processing, including in vision (Dosovitskiy et al., 2020), auditory (Verma & Berger, 2021), and neurological domains (Wang et al., 2023). In bioinformatics, researchers are now harnessing Transformers to bypass genome-specific hard-coded rules, a development we contribute to with our novel method for Sequence Alignment.
1.1 Transformer models: Written Language to DNA Sequence Alignment
DNA sequences share remarkable similarities with written language, offering a compelling avenue for the application of Transformer models. Like written language, these are sequences generated by a small alphabet of nucleotides $\{A,T,G,C\}$. Classical DNA modeling efforts have already accommodated mature encoding and hashing techniques initially developed for written language – such as Suffix trees/arrays and Huffman coding (Huffman, 1952; Manber & Myers, 1993) – to successfully parse and compress DNA sequences. Furthermore, just as written language contains repeated subsequences (words, phrases) to represent real-world objects, DNA sequences similarly possess repeating “words” and groupings of such words into a “sentence” representing, for example, genes.
Within the last few years, several Transformer-based models have been developed for DNA sequence analysis. Notably, DNABERT-2 (Ji et al., 2021; Zhou et al., 2023), Nucleotide Transformer (Dalla-Torre et al., 2023), GenSLM (Zvyagin et al., 2022), and GENA-LM (Fishman et al., 2023) have been designed to discern relationships between short genetic fragments and their functions. Specifically, Nucleotide Transformer representations have shown utility in classifying key genomic features such as enhancer regions and promoter sequences. Similarly, GENA-LM has proven effective in identifying enhancers and Poly-adenylation sites in Drosophila. In parallel, DNABERT-2 representations have also been found to cluster in the representation space according to genetic function. Given these advances, a natural question arises: Can these Transformer architectures be readily applied to the task of Sequence Alignment? We delineate the associated challenges as follows:
Figure 2: Alignment Recall of Transformer-DNA Baselines by Read Length: Existing Transformer-DNA models were adapted for sequence alignment using mean-/max-pooling. Their performance, measured by recall (top-$K$) over $40K$ reads of varying lengths across the human genome (3gb - single-haploid), is shown. Trendlines represent each baseline, with error bars (Clopper-Pearson Interval (Clopper & Pearson, 1934) @ 95%) in grey. The vertical line at $x=250$ marks a typical read length. Overall, these baselines show suboptimal performance. For more details, see Sec. 4.
  • [L1]
    Two-Stage Training: DNA-based Transformer models typically undergo pretraining via a Next Token/Masked Token Prediction framework, a method originally developed for natural language tasks. To form sequence-level representations, these models often employ pooling techniques that aggregate token-level features into a single feature vector. This approach, however, is sometimes critiqued for yielding suboptimal aggregate features  (Reimers & Gurevych, 2019).
  • [L2]
    Computation Cost: The computational requirements for Transformer models grow quadratically with the length of the input sequence. This is particularly challenging for sequence alignment tasks that necessitate scanning entire genomic reference sequences.
Figure 2 shows the sequence alignment performance (recall) of several Transformer-DNA models. The testing protocols are elaborated in Sec. 4. Notably, these models exhibit subpar recall performance when aligning typical read lengths of 250.
2 Our Contributions
In this paper, we argue that both limitations L1, L2 of Transformer-DNA models can be mitigated by formulating sequence alignment as a vector search-and-retrieval task. Our approach is twofold: (A) We introduce a sequence encoder DNA-ESA, trained through self-supervision, to map DNA reads to relevant fragments in a reference sequence within a shared embedding space. (B) We leverage a specialized data structure, termed a DNA vector store, as a memory bank. This provides efficient access to the entire reference sequence for each read alignment. These strategies have been explored in NLP: (A) Sequence-to-embedding training using contrastive loss has shown improved performance – over explicit pooling methods – at abstractive semantic tasks such as prose summarization and paragraph classification (Gao et al., 2021; Chen et al., 2020a). (B) Specialized data structures, such as “vector stores” or “vector databases” like FAISS (Johnson et al., 2019) and Pinecone, use advanced indexing and retrieval algorithms for scalable numerical representation search.
3 Methods
We formulate the problem of Sequence Alignment as minimizing a sequence alignment function, SA, applied to a read $r$ and a reference sequence $\mathcal{R}$ as
$v^{*}=\min_{q}\texttt{SA}(r,\mathcal{R})$ (2)
where $q\in\mathbb{N}_{0}$ is a candidate reference starting position and $v^{*}$ is the optimal alignment score. Lower scores indicate better alignments. This optimization exhibits the following property:
[P1] Sharding for sequence alignment: for a read segment $r$ of length $Q$ and reference $\mathcal{R}$ of length $N$, the complexity of $\texttt{SA}(r,\mathcal{R})$ scales as $\mathcal{O}(Q)$ when $N\rightarrow Q$.
Using P1, we can simplify the optimization problem by breaking it into sub-tasks with significantly shorter reference sequences. Specifically:
$v^{*}\approx\min_{\mathcal{F}_{i}\in\{\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{K}\}}\texttt{SA}\left(r,\mathcal{F}_{j}\right).$ (3)
Here, each $\mathcal{F}_{j}$ is a fragment of $\mathcal{R}$ (i.e., $\mathcal{F}_{j}\in\mathcal{R}$), and $K$ is the number of these sub-tasks [^2] We denote the relative distance between an alignment and the optimal score at read length $Q$ as $d_{SW}=\frac{|mQ-v^{*}|}{|mQ|}$ where $m(=-2)$ is the match score while computing the SW distance.. This approximation is effective under the conditions:
  • (1)
    Fragment $\mathcal{F}_{j}$ lengths are on the order of the read length ($r$), not the longer reference ($\mathcal{R}$);
  • (2)
    There are enough fragments $\mathcal{F}_{i}$ to cover $\mathcal{R}$, i.e. $\cup\mathcal{F}_{j}=\mathcal{R}$;
  • (3)
    $K$ is significantly smaller than $\frac{N}{Q}$. If $\frac{N}{Q}$ then this amounts to scanning the whole reference.
Conditions (1) and (2) imply that fragments should be short and numerous enough to cover the reference genome. Condition (3) restricts the number of retrieved reference fragments per read — that we deem to be most likely to contain $r$ — to a small value $K$. Analogous methods have shown efficacy in text-based Search-and-Retrieval tasks  (Peng et al., 2023; Dai et al., 2022) on Open-Domain Question-Answering, Ranking among other tasks. Subsequent sections describe a parallel framework for retrieving reference fragments given a read. The pipeline is shown in Figs. 3 and 4.
Figure 3: System Overview [A] - Training Encoder and Populating Vector Store: Reference genome fragments $\mathcal{F}_{i}$ and within them, randomly sampled pure reads $r_{i}$ (positive pairs) are numerically represented via shared encoder $h$. Encoder training follows a contrastive approach as per equation 5. After training, the genome is segmented into overlapping fragments, encoded, and uploaded into the vector store.
3.1 Designing effective sequence representations
An optimal sequence encoder model $h$ is such that the corresponding embeddings of any read $r$ and reference fragment $\mathcal{F}$$h(r),h(\mathcal{F})$ respectively – obey the following constraints over a pre-determined distance metric $d$:
$d\{h(r_{j}),h(\mathcal{F}_{i})\}\geq d\{h(r_{j}),h(\mathcal{F}_{j})\},\quad i\neq j$ (4)
Here $i$ and $j$ serve to distinguish whether a read is aligned to a particular reference fragment, a positive sample $\{r_{j},\mathcal{F}_{j}\}$, or there is a mismatch (negative sample): $\{r_{j},\mathcal{F}_{i}\}$. Observe that these inequalities constitute the only requirements for the encoder. As long as the neighborhood of $r_{j}$ in the representation space contains the representation for $\mathcal{F}_{j}$, it will be recovered in the nearest neighbors (top-K set) and alignment will succeed. Equality is observed when $r_{j}$ is a repeat sequence matched equally well to more than one fragment. This motivates using self-supervision (Hadsell et al., 2006; Chen et al., 2020a; Gao et al., 2021) where we are only concerned about the relative distances between positive and negative (read, reference fragment) pairs.
3.2 Self-supervision and contrastive loss
A popular choice for sequence learning using self-supervision involves a contrastive loss setup described by  Chen et al. (2020a) and Gao et al. (2021): i.e. for a read $r$ aligned to reference fragment $\mathcal{F}_{j}$, the loss $l_{r}$ simultaneously minimizes the distance of $h(r)$ to $h(\mathcal{F}_{j})$ and maximizes the distance to a batch of random fragments of size $B-1$:
$l_{r}=-\log\frac{e^{-d(h(r),h(\mathcal{F}_{j}))/\tau}}{e^{-d(h(r),h(\mathcal{F}_{j}))/\tau}+\sum^{B-1}_{i=1}e^{-d(h(r),h(\mathcal{F}_{i}))/\tau}}.$ (5)
Here $\tau$ is a tuneable temperature parameter. To stabilize the training procedure and reach a non-trivial solution, the encoder applies different dropout masks to the reads and fragments similar to the method described in  Chen et al. (2020b). Similar setups have been shown to work in written language applications, most notably in Sentence Transformers (Reimers & Gurevych, 2019; Gao et al., 2021; Muennighoff et al., 2023), which continue to be a strong benchmark for several downstream tasks requiring pre-trained sequence embeddings.
3.3 Encoder implementation
DNA-ESA uses a Transformer-encoder (Devlin et al., 2019; Vaswani et al., 2017), comprising $12$ heads and $6$-layers of encoder blocks. The size of vocabulary is $10,000$. Batch size $B$ is set to $16$ with gradient accumulation across $16$ steps. The learning rate is annealed using one-cycle cosine annealing (Smith & Topin, 2019), dropout is set to $0.1$, and $\tau=0.05$. Reference fragment $|\mathcal{F}_{i}|\sim\mathcal{U}[800,2000]$ and read $|r_{i}|\sim\mathcal{U}[150,500]$. Shorter sequences were padded to equal the length of the longest sequence in a batch. The distance metric used is Cosine Similarity.
3.4 Search and retrieval
Figure 4: System Overview [B] - Inference on a New Read: A read, as per equation 1 and generated by ART (Huang et al., 2012), is encoded by $h$. This is then compared to reference fragment representations in the vector database. The nearest-K fragments in the embedding space are retrieved for each read, and the optimal alignment is determined using equation 6.
An outline of the search and retrieval process is presented in Fig. 4. Every read is encoded using the trained model and matched to reference fragments in the vector database. The top-K retrieved fragments per read are then aligned using a SW alignment library to find the optimal alignment. The following sections describe the indexing and retrieval part in more detail.
Indexing: For a given reference genome $\mathcal{R}$, we construct a minimal set of reference fragments $\mathcal{F}:={\mathcal{F}_{1},\mathcal{F}_{2}},\dots$ to span $\mathcal{R}$. Note that the fragments overlap at least a read length; i.e. $|\mathcal{F}_{i}\bigcap\mathcal{F}_{i+1}|\geq Q$ to guarantee that every read is fully contained within some fragment in the set. In our experiments with external read generators (Huang et al., 2012), $Q_{max}=250,|\mathcal{F}_{i}|=1250$. Each reference fragment is encoded using the trained DNA-ESA model, and the resulting sequence embeddings ($\in\mathbb{R}^{384}$) – $3$M vectors for a reference of $3$B nucleotides – are inserted into a Pinecone database. Once populated with all the fragments, we are ready to perform the alignment.
Retrieval: Given a read $r$, we project its corresponding DNA-ESA representation into the vector store and retrieve the approximate nearest-$K$ set of reference fragment vectors and the corresponding fragment metadata $\{\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{K}\}$.
Diversity priors: While the top-$K$ retrieved fragments can be drawn from across the entire vector store (genome), contemporary recommendation systems that use the top-$K$ retrieval setup rank and re-rank top search results (Slate Optimization – see Zhu et al. (2007)) to ensure rich and diverse recommendations. Similarly, we apply a uniform prior wherein every retrieval step selects the top-$K$ per Chromosome.
Fine-Alignment: A standard SW distance library (Cock et al., 2009) is used to solve equation 3, which can be executed concurrently across the $K$-reference fragments. Let the optimal fragment be $\mathcal{F}^{*}$. The metadata for each vector includes (a) the raw $\mathcal{F}^{*}$ sequence; (b) the start position of $\mathcal{F}^{*}$ within the reference $\mathcal{R}$, $q_{\mathcal{F}^{*}|\mathcal{R}}$. Upon retrieval of a fragment and fine-alignment to find the fragment-level start index, $*q_{|\mathcal{F}^{*}}$, the global reference start position is obtained as:
$q^{*}=q_{|\mathcal{F}^{*}}+q_{\mathcal{F}^{*}|\mathcal{R}}.$ (6)
4 Transformer-DNA baselines
This section outlines the setup for evaluating Transformer-DNA baselines, with their recall performance depicted in Fig. 2. We selected three architectures modeling nucleotide sequences: [NT] NucleotideTransformer ($\in\mathbb{R}^{1280}$(Dalla-Torre et al., 2023), [DB2] DNABERT-2 ($\in\mathbb{R}^{768}$(Ji et al., 2021), and [HD] HyenaDNA ($\in\mathbb{R}^{256}$(Nguyen et al., 2023). Each model employs mean- and max-pooling of token representations for sequence encoding ($2\times 3=6$ baselines total). Independent vector stores for each baseline encode fragments from the entire $3$gb genome. We sampled $40$K pure reads of varying lengths ($Q\sim\mathcal{U}[25,1000]$) and assessed the average recall for top-$5$, top-$25$, and top-$50$ fragments, as shown in Fig. 2. Overall, while baseline performance is modest, mean-pooling generally outperforms max-pooling, with DB2 (mean-pooled) and HD (max-pooled) as the most effective. These two baselines will be contrasted with DNA-ESA in Table 1.
5 Results and Discussion
DNA-ESA convergence plots are presented in Figs. 4(a) and  4(b). Model checkpoints are available at OSF. In Fig. 1, representations of short $1,000-$length sequences sampled from sequential (in-order) and gene-specific locations in the reference are visualized in a reduced 2D-UMAP (McInnes et al., 2018). The representation space demonstrates desired properties suitable for successfully performing alignment: (a) Sequences sampled in order form a trajectory in the representation space: The loss function described in equation 5 encourages a pair of sequences close to one another to have a short distance between them in the representation space, and pairs further apart to have a larger distance. (b) Representations of sequences drawn from specific gene locations – despite not being close to one another – show gene-centric clustering: The DNA-ESA representation space partially acquires function-level separation as a byproduct of imposing local alignment constraints. Codebase is linked.
5.1 Sequence alignment of ART-simulated reads
The results from Sec. 4 demonstrate that even for pure reads, baseline models do not generate adequate representations to perform sequence alignment. In this section, DNA-ESA and the two best baselines – DB2, mean and HD, max – are evaluated using reads generated from an external read simulator (ART) – see Huang et al. (2012). ART has served as a reliable benchmark for evaluating other contemporary alignment tools and provides controls to model mutations and variations common in reads generated by Illumina machines.
Simulator configurations: The different simulation configuration options and settings are listed: (A) Phred quality score $Q_{PH}\in\{[30,60],[60,90]\}$: the likelihood of errors in base-calls of a generated read; (B) Insertion rate $I\in\{0,10^{-2}\}$: the likelihood of adding a base to a random location in a read; (C) Deletion rate $D\in\{0,10^{-2}\}$: the likelihood of deleting a base in the read; (Others): Simulator system: MSv3 [MiSeq]; Read length: $250$.
Recall configurations: Once the top-K fragments have been retrieved, the first step is to solve equation 3, and for this, we need to compute the SW distance. For all presented results, the settings are: match_score = -2, mismatch_penalty = +1, open_gap_penalty = +0.5, continue_gap_penalty = +0.1. After alignment, we get $q^{*}$ – see equation 6 – as the estimated location of a read in the genome. Let $\hat{q}^{*}$ be its true location. If $q^{*}=\hat{q}^{*}$, it is a perfect match and the recall is successful. In cases where there is a mutation in the first or last position in a read, the fine-alignment will return $q^{*}_{\mathcal{F}^{*}}$ offset by at most $2$ locations, resulting in $\hat{q}^{*}=q^{*}\pm 2$. Hence, the condition for an exact location match: $|q^{*}-\hat{q}^{*}|\leq 2$.
Distance bound, $d_{SW}\in\{\text{None},1\%\}$: It is well known that short fragments frequently repeat in the genome (Li & Freudenberg, 2014) and $q^{*}$ can correspond to the position of the read in a different location than from where it was sampled. In this case, $q^{*}\neq\hat{q}^{*}$, but the SW distance is the minimum possible ($d_{SW}=0$ – see Footnote 2). Moreover, when reads have mutations, the reference sequence corresponding to the read is no longer a perfect match; i.e. $d_{SW}>0$. The best an alignment algorithm could do is to find an exact match for the read leading to an optimal alignment score for that read length ($-2Q$). We consider an alignment (with $Q=250$) to be successful if $d_{SW}<0.01$; i.e. a mismatch of at most $2$ bases.
DNA-ESA (ours)       DB2, mean HD, max
I D Top-$K$ 50 / Chr. $d_{SW}=1\%$ Best Top-K @ 50 / Chr. + $d_{SW}=1\%$
$Q_{PH}\in[60,90]$
0 0 95.2 ± 0.63 + 1.6 96.8 ± 0.52 37.4 ± 1.35 14.6 ± 1.01
0 0.01 95.8 ± 0.59 + 1.3 97.1 ± 0.50 37.4 ± 1.35 14.8 ± 1.01
0.01 0 95.1 ± 0.63 + 2.1 97.2 ± 0.49 35.9 ± 1.35 14.6 ± 1.01
0.01 0.01 95.8 ± 0.59 + 1.2 97.0 ± 0.51 34.9 ± 1.34 14.5 ± 1.01
$Q_{PH}\in[30,60]$
0 0 95.2 ± 0.63 + 1.4 96.6 ± 0.54 35.4 ± 1.34 13.7 ± 0.98
0 0.01 95.1 ± 0.64 + 2.0 97.1 ± 0.49 37.3 ± 1.35 14.2 ± 0.99
0.01 0 94.2 ± 0.68 + 2.1 96.3 ± 0.56 36.0 ± 1.35 13.8 ± 0.98
0.01 0.01 94.9 ± 0.64 + 1.5 96.4 ± 0.55 36.8 ± 1.35 13.5 ± 0.98
Table 1: Performance of DNA-ESA with respect to baselines: (with diversity priors – see Def. 4) The performance of DNA-ESA approeaches that of conventional algorithmic and far exceeds the performance of Transformer-DNA baselines, DB2, mean and HD, max (top performing baselines from Fig. 2). Both DNA-ESA and baselines utilize a dedicated vector store for the entire genome, maintaining a consistent search strategy. For details on $I,D,Q_{PH},K,d_{SW}$, refer to Recall/Simulator Configurations.
Performance: Table 1 contains the result of sweeps along these several parameters in addition to a direct comparison to DB2, mean and HD, max baselines, the best-performing baselines on pure reads in Sec. 4. We observe the following: (A) DNA-ESA demonstrates strong recall of $\sim 97\%$ across a variety of read generation and recall configurations described in Sec. 5.1; (B) Reads with less noise – high $Q_{PH}$, low $I,D$ – are more often correctly aligned; (C) The $d_{SW}$ bound adds $\sim 1.5-2\%$ in recall (to cover $5\%$ of misalignments) – a $\sim 20\%$ boost. This suggests that in the cases of misses, while finding the exact index match, the retrieved reference fragments are still high-quality retrievals that differ by at most $2$ bases; (D) While the precise method of (a) generating reads (ART settings, gapped, long vs. short, paired-end vs. single-end, etc.), (b) accounting for the variation in the reference (with or without mitochondrial DNA, single- or double haploid, etc.), (c) detecting a successful alignment (duplicates, distance criteria) vary considerably, DNA-ESA approaches the performance reported by mature algorithmic methods of StrobeAlign, BWA-Mem2, and Minimap (Sahlin, 2022; Vasimuddin et al., 2019; Li, 2018).
5.2 Task transfer from Chromosome 2
Are these results indicative of DNA-ESA’s adaptability to new genomic sequences, rather than a strict adherence to its training data? This would suggest the model’s learning to solve the sequence alignment task rather than memorizing the genome.
[…]
Acc. $\uparrow$ F1 $\uparrow$ P $\uparrow$ R $\uparrow$
Train Reads 2 (seen in training) 3 (populated) Y (populated)
Chr. 2 Pure 98.0±0.43 98.5±0.39 99.0±0.28 97.5±0.46 98.2±0.41 99.3±0.27
ART 97.9±0.43 97.8±0.43 97.8±0.43 97.9±0.43 97.6±0.44 98.0±0.43
Table 2: DNA-ESA Performance on Chr. 2 and Task Transfer to Chr. 3, Y: DNA-ESA, trained on Chr. 2, is evaluated for aligning positive samples from Chr. 2 and negative samples from Chr. 3. Anticipated failure to align Chr. 3 reads (false positives) is due to significant SW distance. Metrics reported include weighted accuracy (Acc.), precision (P), F1-score (F1), and recall (R), for comparison with Table 1. [Grey] For Chr. 3 and Y, separate vector stores are populated using DNA-ESA representations learned using Chr. 2 {read,fragment} pairs. These results support the generalizing tendency of DNA-ESA to align sequences beyond the training reference.
Species Pure ART
Recall Top-K @50 $\uparrow$
Thermus Aquaticus [All] 100.0 (93.0±0.75 @K=1) 100.0 (95.2±0.63, @K=1)
Acidobacteriota [All] 100.0 (93.2±0.74, @K=1) 100.0 (94.9±0.64, @K=1)
Rattus Norwegicus [Chr. 1] 93.3±0.73 95.4±0.62
Rattus Norwegicus [Chr. 2] 93.2±0.74 96.4±0.55
Pan Troglodytes [Chr. 2A] 94.5±0.67 97.1±0.51
Pan Troglodytes [Chr. 2B] 94.3±0.68 97.2±0.50
Table 3: Cross-Species Task Transfer with DNA-ESA: Training on Human Chr. 2, Testing on Diverse Species: DNA-ESA, trained on human Chr. 2, aligns fragments and reads from different species, including Rattus Norwegicus, Pan Troglodytes, Acidabacteriota, and Thermus Aquaticus. These species, are evaluated for read alignment recall using pure and ART-generated reads. The findings, consistent with Table 1, indicate DNA-ESA’s proficiency in modeling DNA sequence structure, beyond simply memorizing training data. Top-$K=1$ tests are conducted on species achieving 100% recall with $K=50$.
Performance: Details on convergence are in Appendix A. Table 2 shows that the performance on unseen human chromosomes (3, Y) is similar to the performance reported in Table 1, despite only training DNA-ESA on Chr. 2. This suggests DNA-ESA’s ability to generalize sequence alignment across chromosomes with different compositions. High Accuracy, Precision, and F1-scores further confirm the method’s specificity at this high recall. In Table 3, a trend of decreasing recall across species is observed: Human (2, 3, Y) $>$ Chimpanzee (2A, 2B) (Blue) $>$ Rat (1, 2), with chimpanzees likely performing better due to genetic similarities with humans (Ijdo et al., 1991). Even distantly related species like Thermus Aquaticus and Acidobacteriota show significant recall, highlighting DNA-ESA’s task transferability beyond simple data memorization.
5.3 Ablation Studies
We’ve conducted additional experiments on DNA-ESA, detailed in the Appendix: (A) Sec. C.1 covers DNA-ESA’s performance on noisy reads from ART simulators and the PacBio CCS dataset. (B) In Sec. C.2, we explore its performance under various $d_{SW}$ bounds ({5%,10%}) and top-$K$ settings ({10,20}) per chromosome. High performance at lower $K$ values suggests a more effective representation space, while the $d_{SW}$ range assesses fragment retrieval accuracy without exact index matches. (C) Sec. C.3 describes outcomes when omitting diversity priors in searches across the whole genome (top-$K$), resulting in predictably lower performance. (D) Sec. C.4 evaluates DNA-ESA’s effectiveness with different read lengths, including those beyond $Q=250$. Notably, performance improves with longer reads, demonstrating a Zero-shot effect for lengths not included in training.
6 Future Work
While DNA-ESA shows promising recall performance when compared to traditional algorithmic methods, its current computational speed, at 200 reads per minute, is a limiting factor for extensive genomic studies involving millions of reads. Despite this, the alignment task’s inherent parallelizability (outlined in [P1]) offers avenues for efficiency improvements. Our ongoing and future efforts are directed towards overcoming this time efficiency bottleneck, as detailed in Appendix Sec. B. We are considering various optimization strategies, including model compilation to speed up inference and enhanced parallelization in vector store searches and fragment-read alignment. Additionally, we aim to enhance DNA-ESA’s performance with shorter read lengths, exploring alternative training methods, integrating more diverse read/reference features, and employing data augmentation techniques.
7 Concluding Remarks
Current DNA sequence alignment methods rely on algorithmic approaches that have been refined over decades, incorporating DNA-specific enhancements in both indexing and retrieval. We have introduced an alternate data-driven paradigm. Employing a Transformer-based DNA-ESA encoder, our approach performs sequence alignment through self-supervised learning using contrastive loss. While such methods have previously been used for identifying (approximate) semantic similarity between sequences in written language applications, we demonstrate the surprising ability to find exact overlaps among DNA sequences. Our empirical results also show that the model, once trained, embodies the inherent structure of any DNA sequence (up to context length) irrespective of the location or species origin: a foundation model. DNA-ESA’s innovation lies in its ability to represent various sequence lengths within a shared vector space, facilitating a ‘flat’ search approach. Unlike traditional methods that rely on hierarchical search, DNA-ESA treats shorter reads and longer reference fragments equally, using cosine similarity as a symmetric distance metric. This novel approach not only challenges existing paradigms but also enhances the potential for discovering novel DNA sequence representations. Mirroring the impact of Sentence Transformers in addressing sequence-level tasks in NLP, DNA-ESA, combined with traditional techniques, could lead to breakthroughs like aligning reads to the Pan Genome, where variations in the reference sequences model the genetic variations across individuals.
Appendix
Appendix A Convergence
Fig. 5 plot the convergence of the DNA-ESA encoder model discussed in the main text. We see convergence after $\sim$20k steps.
(a) Trained on the Human Genome
(b) Trained on the Human Chromosome 2
Figure 5: DNA-ESA convergence plots. Both plots show the loss pr. step (grey). For clarity, we smooth the loss using a moving average (black).
Appendix B Complexity of computing alignment
B.1 Cost of constructing a new representation
Computing the embedding $\mathcal{E}$ of a sequence of length $F$ using DNA-ESA encoding – a typical Transformer-based attention architecture – has the following computation complexity:
$\mathcal{O}(LH*(F^{2}*d+d^{2}*F))\Rightarrow\mathcal{O}(F^{2}*d+d^{2}*F)$ (7)
Where $d$ is the embedding dimension of the model, $L$ is the number of layers in the Transformer and $H$ is the number of heads per layer. As $d$ is a controllable parameter for the model, we can further simplify:
$\mathcal{O}(F^{2}*d+d^{2}*F)\Rightarrow\mathcal{O}(F^{2})$ (8)
The $F^{2}$ complexity follows the basic implementation of attention in transformers, but recent efforts (Beltagy et al., 2020; Kitaev et al., 2020) have developed shortcuts to reduce the cost. These have already been applied to DNA sequence modeling  (Nguyen et al., 2023).
B.2 Vector Store Upstream
Vector store $\mathcal{D}$ is populated once (in bulk) with encoded fragment-length sequences drawn from the entire genome; constant time complexity $C$ to upload $<10M$ vectors.
B.3 Retrieval Cost
Given a new embedding, ${\epsilon}(G,K)$ is the cost of retrieving top-$K$ nearest neighbors across the fragment embeddings, where $G$ is the length of the reference genome. In modern vector databases, where several hashing techniques such as approximate K-nearest neighbors are used, ${\epsilon}$ scales logarithmically with $G$. This is indeed the key benefit of using such databases instead of a naïve search technique.
B.3.1 Fine-grained Alignment
Existing libraries/algorithms (e.g. the Smith–Waterman algorithm) can identify the alignment between a fragment sequence (of length $F$) and read (of length $Q$) in $\mathcal{O}(FQ)$.
B.3.2 Total Complexity
Total complexity involves (a) constructing the representation of a read; (b) querying the vector store; (c) running fine-grained alignment with respect to the $K$ returned reference fragment sequences:
$\mathcal{O}(F^{2}+FQK+\epsilon(G,K))\Rightarrow\mathcal{O}(F(F+QK)+\epsilon(G,K))\Rightarrow\mathcal{O}(FQK+\epsilon(G,K))$
Appendix C Ablation studies
C.1 Performance on high-noise reads
In Table 4, we report the performance of DNA-ESA on two noisy datasets: (i) ART-generated (MiSeqv3-based) reads of length $Q=250$ with Phred quality score $Q_{PH}\in[10,20],I,D=1\%$ and (ii) reads from PacBio CCN as generated on the Ashkenazim Trio (Son) (retrieved from the Genome in a Bottle resource (GIAB) Zook et al. (2016)). For (ii), the raw reads are 10 kilobases long and for evaluation we consider random subset of $1000$ reads derived to Chr. 2. Reads specific to a chromosome are filtered using the pbmm2 (Minimap) package Li (2018). Reads input into the aligner are cropped to length $250,350$ and $500$. Note that in these experiments, we take care to increase the fragment-to-fragment overlap $|\mathcal{F}_{i}\bigcap\mathcal{F}_{i+1}|$ from $250$ to $500$. DNA-ESA performance is reported with respect to the Bowtie-2 performance. While Bowtie-2 has a $>99.9\%$ recall, DNA-ESA isn’t far behind and is comparable (within $1\%$) to the results on less noisy reads reported in Table 1. While the recall is lower, the quality of the recovered fragments is comparable: the average Smith-Waterman distance for all successfully aligned {read, fragment} pairs are similar.
MiSeqv3 - Q = 250
(via ART, PH in [10,20], I,D:1%)
PacBio, CCS - pbmm2, chr2
(Ashkenazim Trio, Son)
Model
Very-high
Sensitivity
Alignment
SW (500)
Very-high
Specificity
Alignment
SW (500)
Q=250 Q=350 Q=500
Bowtie-2 >99.9 490.2
87.0 ± 0.94
(–mp 1000
–np 1000)
490.4 <+2.0% <+2.8% <+2.5%
DNA-ESA
94.4 ± 0.60
(dSW=30)
488.3
89.0 ± 0.87
(@dSW=15)
490.6 97.5 ± 0.47 96.5 ± 0.55 97.0 ± 0.51
Table 4: DNA-ESA performance on noisy reads: Noisy reads are simulated using ART ($Q_{PH}=20$, $I,D=1\%$) and derived from an external PacBio CCS read dump available in GIAB. The reads from PacBio are filtered to those derived from Chr. 2 as determined using the pbmm package Li (2018). In this evalaution, DNA-ESA is contextualized with Bowtie-2 Langmead et al. (2018), a popular conventional alignment tool. On both datasets, DNA-ESA performs slightly worse than Bowtie-2. Despite this, the quality of the retrieved fragments – as measured using the SW distance – is very similar. Note that in all cases, the performance of DNA-ESA is comparable (within $1\%$) to those reported in Table. 1.
C.2 Sweeping Top-K and SW distance bound
In Table 5, we report the performance of DNA-ESA with larger Smith-Waterman distance bounds $d_{SW}\in\{10\%,20\%\}$ and a smaller number of recalled fragments $K\in\{10,20\}$. Distance bound $d_{SW}<10\%$ is equivalent to an acceptable mismatch of at most $\sim 8$ bases between the read (length $Q=250$) and recalled reference fragments. We observe that with small $K$, the recall is $>91\%$ (exact index match). As the distance bound increases, performance predictably improves; however, the performance gain from $d_{SW}=25$ to $d_{SW}=50$ is small. This implies that many retrieved fragments are high-quality retrievals; i.e. of high-likelihood to align with the read, and do not benefit from a more generous distance bound to improve recall performance significantly. Furthermore, decreasing the top-$K$ per chrosomosome from $20\rightarrow 10$ does not substantially worsen performance ($\sim 2\%$) indicating that the optimal retrievals are usually the closest in the embedding space.
$Q_{PH}\in[30,60]$ $Q_{PH}\in[60,90]$
$I$ $D$ Exact $d_{SW}<5\%$ $d_{SW}<10\%$ Exact $d_{SW}<5\%$ $d_{SW}<10\%$
Top-10 / chromosome
0.0 0.0 91.5±0.81 96.6±0.54 96.6±0.54 91.9±0.79 96.4±0.55 97.3±0.49
0.01 92.0±0.79 96.5±0.55 96.8±0.52 92.0±0.79 96.2±0.57 97.6±0.47
0.01 0.0 91.9±0.79 96.4±0.55 96.5±0.55 92.0±0.79 96.6±0.54 97.4±0.48
0.01 91.8±0.80 96.3±0.56 97.2±0.50 91.5±0.81 96.3±0.56 97.3±0.49
Top-20 / chromosome
0.0 0.0 93.4±0.73 97.5±0.47 98.0±0.43 93.7±0.71 97.7±0.45 97.8±0.44
0.01 93.2±0.74 97.2±0.49 98.2±0.41 93.6±0.72 97.5±0.47 97.8±0.45
0.01 0.0 93.0±0.74 97.9±0.44 97.9±0.43 92.6±0.76 97.4±0.48 97.8±0.44
0.01 93.0±0.74 97.4±0.48 97.8±0.45 93.4±0.72 97.8±0.45 98.1±0.42
Table 5: Sequence alignment recall of DNA-ESA sweeping top-K and $d_{SW}$: The various parameters are described in Sec. 5.1. DNA-ESA presents a recall of $>97\%$ across several read configurations rivaling contemporary algorithmic models. Performance predictably degrades as more noise is introduced into a read. Performance improves with larger search radius (top-$K$), higher quality reads ($Q_{PH}$) and large distance bound $d_{SW}$.
C.3 Without diversity priors in top-K
In Table 6, we report the performance without diversity priors used in the retrieval step: i.e. the nearest-K neighbors in the embedding space are selected from the entire set of fragments spanning the genome rather than uniformly sampling from each chromosome. The performance predicably falls in comparison to those reported in Table 1 since fewer fragments scattered unevenly across the different chromosomes are being retrieved per read.
$Q_{PH}\in[30,60]$ $Q_{PH}\in[60,90]$
$I$ $D$ Exact $d_{SW}<5\%$ $d_{SW}<10\%$ Exact $d_{SW}<5\%$ $d_{SW}<10\%$
Top-100
0.0 0.0 82.6±1.08 92.0±0.79 93.2±0.74 88.5±0.92 91.9±0.79 93.0±0.74
0.01 82.3±1.09 91.8±0.80 92.7±0.76 89.3±0.89 93.2±0.74 93.9±0.70
0.01 0.0 83.3±1.06 92.3±0.77 93.3±0.73 88.7±0.91 92.4±0.77 93.3±0.73
0.01 82.8±1.08 91.6±0.81 92.5±0.76 88.7±0.91 92.4±0.77 93.2±0.73
Top-50
0.0 0.0 81.1±1.11 90.6±0.84 92.0±0.79 86.7±0.97 90.7±0.84 91.9±0.79
0.01 80.5±1.13 90.4±0.85 91.6±0.80 87.4±0.95 91.7±0.80 92.7±0.76
0.01 0.0 81.3±1.11 90.8±0.83 92.3±0.78 86.8±0.97 91.0±0.83 92.2±0.78
0.01 81.2±1.11 90.4±0.85 91.7±0.80 87.1±0.96 91.2±0.82 92.4±0.77
Table 6: Sequence alignment recall of DNA-ESA - without diversity priors: The various parameters are described in Sec. 5.1. DNA-ESA presents a recall of $\approx 90\%$. Similar to the result presented in the main text, performance improves with larger search radius (top-$K$), higher quality reads ($Q_{PH}$) and large distance bound ($d_{SW}$).
C.4 Across read lengths
In Table 7, the performance of DNA-ESA is reported across read lengths. We observe Zero-shot performance at longer read lengths: The model performs better at longer read lengths (even exceeding the read length bound established during training $\mathcal{U}[150,500]$); while evaluating for longer reads, we make sure to guarantee that the reads exist as subsequences of fragments. Improving the performance at shorter read lengths is the subject of future work.
Top-K @ 50
I D $Q=150$ $Q=200$ $Q=250$ $1000500<Q<$ 1000
$Q_{PH}\in[60,90]$ Pure Reads
0 0 84.6 ± 1.03 94.9 ± 0.65 96.8 ± 0.52 >97
0 0.01 85.9 ± 0.99 95.2 ± 0.63 97.1 ± 0.50
0.01 0 85.8 ± 0.99 94.5 ± 0.66 97.2 ± 0.49
0.01 0.01 86.2 ± 0.98 95.3 ± 0.63 97.0 ± 0.51
$Q_{PH}\in[30,60]$
0 0 84.6 ± 1.03 94.6 ± 0.66 96.6 ± 0.54 >96
0 0.01 85.2 ± 1.01 95.0 ± 0.64 97.1 ± 0.49
0.01 0 85.7 ± 1.00 94.5 ± 0.66 96.3 ± 0.56
0.01 0.01 85.8 ± 0.99 94.4 ± 0.67 96.4 ± 0.55
Table 7: DNA-ESA recall performance across read lengths: Performance of DNA-ESA is higher for longer reads including those lengths on which the model was not trained ($Q\in\mathcal{U}[150,500]$). Shorter reads are more challenging for the model potentially due to replicates found across the reference.

### v4

@@
-Embed-Search-Align: DNA Sequence Alignment using Transformer models
+
 Abstract
@@
 Such models have shown early promise in tasks involving classification of short DNA sequences, such as the detection of coding- vs non-coding regions, as well as the identification of enhancer and promoter sequences.
-Performance at sequence classification tasks does not, however, translate to sequence alignment, where it is necessary to conduct a genome-wide search to successfully align every read.
+Performance at sequence classification tasks does not, however, translate to sequence alignment, where it is necessary to conduct a genome-wide search to align every read successfully.
 We address this open problem by framing it as an “Embed-Search-Align” task.
@@
 In particular, DNA-ESA introduces: (1) Contrastive loss for self-supervised training of DNA sequence representations, facilitating rich sequence-level embeddings, and (2) a DNA vector store to enable search across fragments on a global scale.
-DNA-ESA is $>97\%$ accurate when aligning $250$-length reads onto a human reference genome of 3 gigabases (single-haploid), far exceeds the performance of $6$ recent DNA-Transformer model baselines and shows task transfer across chromosomes and species.
-1 Background
-The aim of this paper is to establish a foundation model tailored for DNA sequences, where the vocabulary consists of only a few symbols ($\{A,T,G,C\}$ in this case).
-Numerous DNA Transformer models (Ji et al., 2021; Zvyagin et al., 2022; Fishman et al., 2023; Dalla-Torre et al., 2023) have emerged recently, mainly designed for classification tasks in downstream applications.
-However, these models do not explicitly consider a fundamental distinction between Limited Vocabulary Languages (LVL) (such as, genomes) and natural languages with large vocabulary.
+DNA-ESA is 99% accurate when aligning 250-length reads onto a human reference genome of 3 gigabases (single-haploid), rivaling conventional algorithmic sequence alignment methods such as Bowtie and BWA-Mem.
+DNA-ESA far exceeds the performance of $6$ recent DNA-Transformer model baselines such as Nucleotide Transformer, Hyena-DNA, and shows task transfer across chromosomes and species.
+keywords:
+Transformers, DNA Sequence Alignment, Large Language Models, Vector Stores
+1 Introduction
+Equal Contribution
+Sequence alignment is a central problem in the analysis of sequence data.
+Many DNA sequencers generate short reads that are only a couple of hundred bases long.
+In order to interpret this data, a typical first step is to align the reads to a genome.
+Genomes come in many sizes, but many of the most commonly studied genomes, such as that of humans, are three gigabases long.
+Moreover, most experiments will generate millions of reads.
+Therefore, the resulting task of aligning reads to genomes is computationally expensive, and many decades of work have led to optimized approaches that can perform these tasks with great efficiency.
+However, many limitations remain with existing genome alignment methods.
+One limitation is that a reference genome usually is an approximation of the sequence from a single individual, and does not capture the variability of genomes across populations.
+Another, is that sequence alignment algorithms may not correctly model sequencing errors, and thus may introduce mistakes when trying to score the best match for a sequence against a genome.
+We asked whether some of these limitations could potentially be overcome if we utilize a new paradigm to align reads to genomes by exploiting the Transformer architectures vaswani_attention_2017 adopted for recent language modeling efforts.
+The applications of the Transformer model – and, more generally “Large Language Models” (LLM) – in Bioinformatics applications are still in their infancy, yet hold substantial promise: Transformer models have demonstrated an unprecedented ability to construct powerful numerical representations of sequential data vit ; wav2vec ; gpt3 .
+We establish a foundation model foundation tailored for DNA sequences, where the vocabulary consists of only a few symbols ($\{A,T,G,C\}$ in this case).
+Numerous DNA Transformer models (ji_dnabert_2021, ; zvyagin_genslms_2022, ; fishman_gena-lm_2023, ; nucleotide, ) have emerged recently, mainly designed for classification tasks in downstream applications.
+However, these models do not explicitly consider a fundamental distinction between Limited Vocabulary Languages (LVL) (such as genomes) and natural languages with large vocabulary.
 In LVLs, there is a one-to-one correspondence between the precise symbol ordering and the underlying “meaning”.
 If a protein is encoded by one amino acid sequence, a slightly different sequence (with a few edits) would encode for a different protein with a distinct functionality in the cells.
-Existing foundation Transformer models generate sequence embeddings such that their pairwise distances correspond to class separation, thus sequences with very large edit distances end up with representations that are close by.
-There are however tasks such as Sequence Alignment where the pairwise representation distance has to closely match the sequence edit distance.
-Indeed as portrayed in Figure 2, current DNA Transformer models fail to perform Sequence Alignment task.
-This requires a new kind of foundation model which is able to transfer a precise and computationally expensive distance metric over LVLs to a representation space.
-The simplest sequence alignment task applies to single-end [^1] A DNA fragment is ligated to an adapter and then sequenced from one end only.
-reads.
-Given a reference sequence :=$\mathcal{R}:=\{b_{1},b_{2},\dots,b_{N}\}$ – for the single-haploid human genome (Nurk et al., 2022), $N\approx 3$ gigabases (gb) – the primary objective is to identify the most probable start- and end-positions within this reference for a short DNA read,
+Existing foundation Transformer models, trained on classification tasks, generate sequence embeddings such that their pairwise distances correspond to class separation.
+As a result, pairs of sequences with very large edit distances between them are mapped to numerical representations that are close by.
+There are, however, tasks such as Sequence Alignment where the objective is quite different: the pairwise representation distance has to closely match the sequence edit distance.
+Indeed as portrayed in Figure 1, current DNA Transformer models without any further refinement are unable to perform Sequence Alignment tasks: the representation space is simply not designed for such an objective.
+This demonstrates the need for a new model which is able to transfer a precise and computationally expensive distance metric, such as the edit distance, over LVLs to a continuous representation space.
+The simplest sequence alignment task applies to single-end [^1] A DNA fragment is ligated to an adapter and then sequenced from one end only. reads, where a sequencer generates a read of length $Q$
+$r:=(\tilde{b}_{1},\tilde{b}_{2},\dots,\tilde{b}_{Q}),$ (1)
+where $\tilde{b}_{i}\in\{A,T,G,C\}$.
+In practice, these reads come from individual genomes that do not necessarily match the reference and may contain mutations due to base insertions, deletions, and substitutions.
+Thus, it is assumed that this read is a noisy substring taken from a reference genome sequence $\mathcal{R}:=(b_{1},b_{2},\dots,b_{N}),\,b_{i}\in\{A,T,G,C\}$ and $N\gg Q$; for example, for the single-haploid human genome (chm_consort, ), $N\approx 3$ gigabases (gb), and $Q\approx 250$ though reads of much longer lengths are becoming increasingly affordable and accurate giab .
+The alignment task is to find a substring in $\mathcal{R}$
 […]
-which may contain mutations due to base insertions, deletions, and substitutions.
+such that the edit distance $d(r,\tilde{r})$ – for example the Smith-Waterman (SW) distance – is minimized.
+The primary objective is to identify the most probable location, $q$, of this read within the reference genome.
+Figure 1: Alignment Recall of Transformer-DNA Baselines by Read Length: Existing Transformer-DNA models were adapted for sequence alignment using mean-/max-pooling.
+Their performance, measured by recall (top-$K$) over $40K$ reads of varying lengths across the human genome (3gb - single-haploid), is shown.
+Trendlines represent each baseline, with error bars (Clopper-Pearson Interval (cps, ) @ 95%) in grey.
+The vertical line at $x=250$ marks a typical read length.
+Overall, these baselines show suboptimal performance.
+For more details, see Sec.
+3.
 Computational simulators have been developed to generate synthetic reads that have properties of real reads.
-These simulators mimic the read quality and characteristics produced by actual sequencing machines, thus providing a scalable means for validating new alignment approaches (Huang et al., 2012).
-(a) A $20,000$-long nucleotide sequence from Chr.
-1, 2, and 3 each is randomly sampled and broken up into $100$ consecutive reference fragments (||||) each of length $1000$ and stride $200$.
-For each fragment, we sample a read (–) of 100-250 consecutive base pairs.
-Each sequence is encoded using DNA-ESA and the embeddings are visualized using 2D UMAP (McInnes et al., 2018).
-The colors identify chromosomes and the position of the fragment is coded by its intensity.
-(b) Fragment (||||) and read (–) sequences from five gene regions (listed above in the inset) are similarly projected into a shared latent space.
-Note that for UMAP projections (both subplots (a) and (b)), the consecutive fragments belonging to the same nucleotide sequence constitute an order-preserving 1D manifold.
-A successfully aligned read is observed to be co-located with its corresponding fragment in the embedding space while a read further away is observed to be misclassified.
+These simulators mimic the read quality and characteristics produced by actual sequencing machines, thus providing a scalable means for validating new alignment approaches (huang_art_2012, ).
+((a)) Three $20,000$-long nucleotide sequences were selected, one each from Chr.
+1, 2, and 3.
+Each sequence is broken up into $100$ consecutive reference fragments, each of length $1000$.
+Consecutive fragments have an overlap of $800$ bases (stride $200$).
+Additionally, reads of length $Q\sim\mathcal{U}[100,250]$ are sampled from within each fragment.
+The colors correspond to the respective chromosomes and the position of a particular fragment along the 20K-nucleotide sequence is coded by its color intensity.
+((b)) Five gene regions (listed above in the inset) of different sizes are selected from the reference genome.
+Each gene is broken up into consecutive fragments of length $1000$ with a stride of $200$ (similar to subfigure (a)).
+Additionally, reads of length $Q\sim\mathcal{U}[100,250]$ are sampled from within each fragment.
+The colors correspond to genes and the position of a particular fragment along a specific gene is coded by its color intensity.
 Figure 2: Illustrating DNA-ESA’s Preservation of Sequence Locality in Embedding Space: In DNA-ESA, the reference genome $\mathcal{R}$ is divided into fragments $\mathcal{F}_{i}$, each represented by an embedding $h(\mathcal{F}_{i})$.
@@
 Subfigures (a) and (b) display this emergent geometry, as visualized using DNA-ESA.
-Conventional sequence alignment methods, such as those used in BWA-MEM, have evolved significantly from the basic Smith-Waterman (SW) approach, which computes similarity between a read $r$ and reference $\mathcal{R}$, but is computationally intensive for large genomes.
-Key advancements include: (a) Sharding the reference into smaller fragments for individual searching; (b) Progressive search using phylogenetic trees and distance heuristics for logarithmic complexity scaling; (c) Compression techniques like the Burrows-Wheeler transform (Li & Durbin, 2009), reducing search length; and (d) Multi-core/thread implementations and database instantiations for faster computations and improved data recall (Li, 2018; Vasimuddin et al., 2019; Langmead et al., 2018).
-In this work we explore an alternative paradigm for aligning a read to a genome, drawing parallels with advancements in Natural Language Processing (NLP).
+In both subfigures, the fragment and read representations are jointly visualized in a 2D low-dimensional space constructed by applying UMAP umap to the original embeddings.
+Vertical marker lines (||||) refer to fragments, “–” markers correspond to reads.
+We observe that the consecutive fragments belonging to the same nucleotide (subfigure (a)) or gene (subfigure (b)) sequence constitute an order-preserving 1D manifold.
+Additionally, almost all of the reads are close to their corresponding fragments, making this space viable for the task of Sequence Alignment: by searching for fragments in the neighborhood of an external read represented in the embedding space, one is likely to retrieve a fragment most responsible for the read.
+Conventional sequence alignment methods, such as those used in BWA-MEM, have evolved significantly from the basic SW approach, which computes the similarity between a read $r$ and reference $\mathcal{R}$; using the SW distance on its own, however, is computationally intensive for large genomes.
+Some key advances include: (a) Sharding the reference into smaller fragments for individual searching; (b) Progressive search using phylogenetic trees and distance heuristics for logarithmic complexity scaling; (c) Compression techniques like the Burrows-Wheeler transform (bw_transform, ), reducing the effective search length; and (d) Multi-core/thread implementations and database instantiations for faster computations and improved data recall (minimap, ; bwamem2, ; bowtie, ).
+In this work, we explore an alternative paradigm for aligning a read to a genome, drawing parallels with advancements in Natural Language Processing (NLP).
 Traditional NLP relied on rule-based techniques to diagram sentences into grammatical components, using dependency parse trees to discern semantic relationships.
-However, the introduction the introduction of models like the Transformers (Vaswani et al., 2017) revolutionized this approach.
-Unlike rule-based frameworks, Transformers implicitly identify syntactic and semantic structures, as evidenced in various NLP tasks like Sentiment Analysis, Entity Recognition, and Question-Answering (Devlin et al., 2019; Reimers & Gurevych, 2019; Qiu et al., 2020).
-The flexibility of this architecture has facilitated its application beyond language processing, including in vision (Dosovitskiy et al., 2020), auditory (Verma & Berger, 2021), and neurological domains (Wang et al., 2023).
+However, the introduction of models like the Transformers (vaswani_attention_2017, ) revolutionized this approach.
+Unlike rule-based frameworks, Transformers implicitly identify syntactic and semantic structures tenney , as evidenced by their performance at various NLP tasks like Sentiment Analysis, Entity Recognition, and Question-Answering (devlin_bert_2019, ; reimers_sentence-bert_2019, ; reviewllm, ).
+The flexibility of this architecture has facilitated its application beyond language processing, including in vision (dosovitskiy_image_2020, ), auditory (audiotransformer, ), and neurological domains (wang_brainbert_2023, ).
 In bioinformatics, researchers are now harnessing Transformers to bypass genome-specific hard-coded rules, a development we contribute to with our novel method for Sequence Alignment.
@@
 Like written language, these are sequences generated by a small alphabet of nucleotides $\{A,T,G,C\}$.
-Classical DNA modeling efforts have already accommodated mature encoding and hashing techniques initially developed for written language – such as Suffix trees/arrays and Huffman coding (Huffman, 1952; Manber & Myers, 1993) – to successfully parse and compress DNA sequences.
-Furthermore, just as written language contains repeated subsequences (words, phrases) to represent real-world objects, DNA sequences similarly possess repeating “words” and groupings of such words into a “sentence” representing, for example, genes.
+Classical DNA modeling efforts have already accommodated mature encoding and hashing techniques initially developed for written language – such as Suffix trees/arrays and Huffman coding (huffman_method_1952, ; manber_suffix_1993, ) – to successfully parse and compress DNA sequences.
+Furthermore, just as written language contains repeated subsequences (words, phrases) to represent real-world objects, DNA sequences similarly possess repeating “words” and groupings of such words into a “sentence” representing, for example, subsets of genes.
 Within the last few years, several Transformer-based models have been developed for DNA sequence analysis.
-Notably, DNABERT-2 (Ji et al., 2021; Zhou et al., 2023), Nucleotide Transformer (Dalla-Torre et al., 2023), GenSLM (Zvyagin et al., 2022), and GENA-LM (Fishman et al., 2023) have been designed to discern relationships between short genetic fragments and their functions.
+Notably, DNABERT-2 (ji_dnabert_2021, ; dna2bert, ), Nucleotide Transformer (nucleotide, ), GenSLM (zvyagin_genslms_2022, ), HyenaDNA (nguyen_hyenadna_2023, ) and GENA-LM (fishman_gena-lm_2023, ) have been designed to discern relationships between short genetic fragments and their functions.
 Specifically, Nucleotide Transformer representations have shown utility in classifying key genomic features such as enhancer regions and promoter sequences.
 Similarly, GENA-LM has proven effective in identifying enhancers and Poly-adenylation sites in Drosophila.
-In parallel, DNABERT-2 representations have also been found to cluster in the representation space according to genetic function.
+In parallel, DNABERT-2 representations have also been found to cluster in the representation space according to certain types of genetic function.
 Given these advances, a natural question arises: Can these Transformer architectures be readily applied to the task of Sequence Alignment?
 We delineate the associated challenges as follows:
-Figure 2: Alignment Recall of Transformer-DNA Baselines by Read Length: Existing Transformer-DNA models were adapted for sequence alignment using mean-/max-pooling.
-Their performance, measured by recall (top-$K$) over $40K$ reads of varying lengths across the human genome (3gb - single-haploid), is shown.
-Trendlines represent each baseline, with error bars (Clopper-Pearson Interval (Clopper & Pearson, 1934) @ 95%) in grey.
-The vertical line at $x=250$ marks a typical read length.
-Overall, these baselines show suboptimal performance.
-For more details, see Sec.
-4.
   • [L1]
@@
 To form sequence-level representations, these models often employ pooling techniques that aggregate token-level features into a single feature vector.
-This approach, however, is sometimes critiqued for yielding suboptimal aggregate features  (Reimers & Gurevych, 2019).
+This approach, however, is sometimes critiqued for yielding suboptimal aggregate features  (reimers_sentence-bert_2019, ).
   • [L2]
@@
 The testing protocols are elaborated in Sec.
-4.
+3.
 Notably, these models exhibit subpar recall performance when aligning typical read lengths of 250.
@@
 Our approach is twofold: (A) We introduce a sequence encoder DNA-ESA, trained through self-supervision, to map DNA reads to relevant fragments in a reference sequence within a shared embedding space.
-(B) We leverage a specialized data structure, termed a DNA vector store, as a memory bank.
+(B) We leverage a specialized data structure, termed a DNA vector store, as a memory bank [^2] Codebase available here: https://anonymous.4open.science/r/dna2vec-7E4E/.
 This provides efficient access to the entire reference sequence for each read alignment.
-These strategies have been explored in NLP: (A) Sequence-to-embedding training using contrastive loss has shown improved performance – over explicit pooling methods – at abstractive semantic tasks such as prose summarization and paragraph classification (Gao et al., 2021; Chen et al., 2020a).
-(B) Specialized data structures, such as “vector stores” or “vector databases” like FAISS (Johnson et al., 2019) and Pinecone, use advanced indexing and retrieval algorithms for scalable numerical representation search.
-3 Methods
+These strategies have been explored in NLP: (A) Sequence-to-embedding training using contrastive loss has shown improved performance – over explicit pooling methods – at abstractive semantic tasks such as evidence retrieval lewis and semantic text similarity (gao_simcse_2021, ; chen_simple_2020, ).
+(B) Specialized data structures, such as “vector stores” or “vector databases” like FAISS (johnson_billion-scale_2019, ) and Pinecone, use advanced indexing and retrieval algorithms for scalable numerical representation search.
+2.1 Task Definition
 We formulate the problem of Sequence Alignment as minimizing a sequence alignment function, SA, applied to a read $r$ and a reference sequence $\mathcal{R}$ as
 $v^{*}=\min_{q}\texttt{SA}(r,\mathcal{R})$ (3)
-where $q\in\mathbb{N}_{0}$ is a candidate reference starting position and $v^{*}$ is the optimal alignment score.
+where $q\in\{1,2,3,\dots\}$ is a candidate reference starting position and $v^{*}$ is the optimal alignment score.
 Lower scores indicate better alignments.
 This optimization exhibits the following property:
-[P1] Sharding for sequence alignment: for a read segment $r$ of length $Q$ and reference $\mathcal{R}$ of length $N$, the complexity of $\texttt{SA}(r,\mathcal{R})$ scales as $\mathcal{O}(Q)$ when $N\rightarrow Q$.
-Using P1, we can simplify the optimization problem by breaking it into sub-tasks with significantly shorter reference sequences. Specifically:
+Sharding for sequence alignment: For a read segment $r$ of length $Q$ and reference $\mathcal{R}$ of length $N$, the complexity of $\texttt{SA}(r,\mathcal{R})$ scales as $\mathcal{O}(NQ)$, which is expensive.
+Ideally, we would like to use $\mathcal{R}$ to get a set of fragments $\{\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{K}\}$ that are a subset of the original reference. Then:
 $v^{*}\approx\min_{\mathcal{F}_{i}\in\{\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{K}\}}\texttt{SA}\left(r,\mathcal{F}_{j}\right).$ (4)
-Here, each $\mathcal{F}_{j}$ is a fragment of $\mathcal{R}$ (i.e., $\mathcal{F}_{j}\in\mathcal{R}$), and $K$ is the number of these sub-tasks [^2] We denote the relative distance between an alignment and the optimal score at read length $Q$ as $d_{SW}=\frac{|mQ-v^{*}|}{|mQ|}$ where $m(=-2)$ is the match score while computing the SW distance..
-This approximation is effective under the conditions:
+Here, each $\mathcal{F}_{j}$ is a fragment of $\mathcal{R}$ (i.e., $\mathcal{F}_{j}\subset\mathcal{R}$), and $K$ is the number of these sub-tasks. This approximation is effective under the conditions:
   • (1)
-    Fragment $\mathcal{F}_{j}$ lengths are on the order of the read length ($r$), not the longer reference ($\mathcal{R}$);
+    Fragment $\mathcal{F}_{j}$ lengths are on the order of the read length ($Q$), and not the length of the longer reference $\mathcal{R}$;
   • (2)
@@
 Condition (3) restricts the number of retrieved reference fragments per read — that we deem to be most likely to contain $r$ — to a small value $K$.
-Analogous methods have shown efficacy in text-based Search-and-Retrieval tasks  (Peng et al., 2023; Dai et al., 2022) on Open-Domain Question-Answering, Ranking among other tasks.
+Analogous methods have shown efficacy in text-based Search-and-Retrieval tasks  (peng2023, ; dai2022, ) on Open-Domain Question-Answering, Ranking among other tasks.
 Subsequent sections describe a parallel framework for retrieving reference fragments given a read.
@@
 Figure 3: System Overview [A] - Training Encoder and Populating Vector Store: Reference genome fragments $\mathcal{F}_{i}$ and within them, randomly sampled pure reads $r_{i}$ (positive pairs) are numerically represented via shared encoder $h$.
-Encoder training follows a contrastive approach as per equation 5.
+Encoder training follows a contrastive approach as per (6).
 After training, the genome is segmented into overlapping fragments, encoded, and uploaded into the vector store.
@@
 Equality is observed when $r_{j}$ is a repeat sequence matched equally well to more than one fragment.
-This motivates using self-supervision (Hadsell et al., 2006; Chen et al., 2020a; Gao et al., 2021) where we are only concerned about the relative distances between positive and negative (read, reference fragment) pairs.
+This motivates using self-supervision (hadsell_dimensionality_2006, ; chen_simple_2020, ; gao_simcse_2021, ) where we are only concerned about the relative distances between positive and negative (read, reference fragment) pairs.
 2.3 Self-supervision and contrastive loss
-A popular choice for sequence learning using self-supervision involves a contrastive loss setup described by  Chen et al.
-(2020a) and Gao et al.
-(2021): i.e.
+A popular choice for sequence learning using self-supervision involves a contrastive loss setup described by  chen_simple_2020 and gao_simcse_2021 : i.e.
 for a read $r$ aligned to reference fragment $\mathcal{F}_{j}$, the loss $l_{r}$ simultaneously minimizes the distance of $h(r)$ to $h(\mathcal{F}_{j})$ and maximizes the distance to a batch of random fragments of size $B-1$:
@@
 Here $\tau$ is a tuneable temperature parameter.
-To stabilize the training procedure and reach a non-trivial solution, the encoder applies different dropout masks to the reads and fragments similar to the method described in  Chen et al.
-(2020b).
-Similar setups have been shown to work in written language applications, most notably in Sentence Transformers (Reimers & Gurevych, 2019; Gao et al., 2021; Muennighoff et al., 2023), which continue to be a strong benchmark for several downstream tasks requiring pre-trained sequence embeddings.
+To stabilize the training procedure and reach a non-trivial solution, the encoder applies different dropout masks to the reads and fragments similar to the method described in prior work simclr .
+Similar setups have been shown to work in written language applications, most notably in Sentence Transformers (reimers_sentence-bert_2019, ; gao_simcse_2021, ; muennighoff_mteb_2023, ), which continue to be a strong benchmark for several downstream tasks requiring pre-trained sequence embeddings.
 2.4 Encoder implementation
-DNA-ESA uses a Transformer-encoder (Devlin et al., 2019; Vaswani et al., 2017), comprising $12$ heads and $6$-layers of encoder blocks.
-The size of vocabulary is $10,000$.
+DNA-ESA uses a Transformer-encoder (devlin_bert_2019, ; vaswani_attention_2017, ), comprising $12$ heads and $6$-layers of encoder blocks.
+The size of the vocabulary is $10,000$.
 Batch size $B$ is set to $16$ with gradient accumulation across $16$ steps.
-The learning rate is annealed using one-cycle cosine annealing (Smith & Topin, 2019), dropout is set to $0.1$, and $\tau=0.05$.
-Reference fragment $|\mathcal{F}_{i}|\sim\mathcal{U}[800,2000]$ and read $|r_{i}|\sim\mathcal{U}[150,500]$.
+Generated numerical representations for each read, $h(r)\in\mathbb{R}^{1020}$.
+The learning rate is annealed using one-cycle cosine annealing (smith_super-convergence_2019, ), dropout is set to $0.1$, and $\tau=0.05$.
+During training, the reference fragment $|\mathcal{F}_{i}|\sim\mathcal{U}([800,2000])$ and read $|r_{i}|\sim\mathcal{U}([150,500])$ have variable lengths sampled from a uniform distribution.
+Here, $|x|$ denotes the length of $x$.
+To improve model performance on noisy reads, $1-5\%$ of bases are replaced with another random base in $40\%$ of the reads in a batch.
 Shorter sequences were padded to equal the length of the longest sequence in a batch.
@@
 2.5 Search and retrieval
-Figure 4: System Overview [B] - Inference on a New Read: A read, as per equation 1 and generated by ART (Huang et al., 2012), is encoded by $h$.
+Figure 4: System Overview [B] - Inference on a New Read: A read, as per (1) and generated by ART (huang_art_2012, ), is encoded by $h$.
 This is then compared to reference fragment representations in the vector database.
-The nearest-K fragments in the embedding space are retrieved for each read, and the optimal alignment is determined using equation 6.
+The nearest-$K$ fragments in the embedding space are retrieved for each read, and the optimal alignment is determined using (7).
 An outline of the search and retrieval process is presented in Fig.
@@
 The following sections describe the indexing and retrieval part in more detail.
-Indexing: For a given reference genome $\mathcal{R}$, we construct a minimal set of reference fragments $\mathcal{F}:={\mathcal{F}_{1},\mathcal{F}_{2}},\dots$ to span $\mathcal{R}$.
+Indexing: For a given reference genome $\mathcal{R}$, we construct a minimal set of reference fragments :=$\mathcal{F}:=\{\mathcal{F}_{1},\mathcal{F}_{2},\dots\}$ to span $\mathcal{R}$.
 Note that the fragments overlap at least a read length; i.e.
 $|\mathcal{F}_{i}\bigcap\mathcal{F}_{i+1}|\geq Q$ to guarantee that every read is fully contained within some fragment in the set.
-In our experiments with external read generators (Huang et al., 2012), $Q_{max}=250,|\mathcal{F}_{i}|=1250$.
-Each reference fragment is encoded using the trained DNA-ESA model, and the resulting sequence embeddings ($\in\mathbb{R}^{384}$) – $3$M vectors for a reference of $3$B nucleotides – are inserted into a Pinecone database.
+In our experiments with external read generators (huang_art_2012, ), $Q_{max}=250,|\mathcal{F}_{i}|=1250$.
+Each reference fragment is encoded using the trained DNA-ESA model, and the resulting sequence embeddings ($\in\mathbb{R}^{1020}$) – $3$M vectors for a reference of $3$B nucleotides – are inserted into a Pinecone database.
 Once populated with all the fragments, we are ready to perform the alignment.
 Retrieval: Given a read $r$, we project its corresponding DNA-ESA representation into the vector store and retrieve the approximate nearest-$K$ set of reference fragment vectors and the corresponding fragment metadata $\{\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{K}\}$.
-Diversity priors: While the top-$K$ retrieved fragments can be drawn from across the entire vector store (genome), contemporary recommendation systems that use the top-$K$ retrieval setup rank and re-rank top search results (Slate Optimization – see Zhu et al.
-(2007)) to ensure rich and diverse recommendations.
+Diversity priors: While the top-$K$ retrieved fragments can be drawn from across the entire vector store (genome), contemporary recommendation systems that use the top-$K$ retrieval setup rank and re-rank top search results (Slate Optimization – see grasshopper ) to ensure rich and diverse recommendations.
 Similarly, we apply a uniform prior wherein every retrieval step selects the top-$K$ per chromosome.
-Fine-Alignment: A standard SW distance library (Cock et al., 2009) is used to solve equation 3, which can be executed concurrently across the $K$-reference fragments.
+Fine-Alignment: A standard SW distance library (biopython, ) is used to solve (4), which can be executed concurrently across the $K$-reference fragments.
 Let the optimal fragment be $\mathcal{F}^{*}$.
 The metadata for each vector includes (a) the raw $\mathcal{F}^{*}$ sequence; (b) the start position of $\mathcal{F}^{*}$ within the reference $\mathcal{R}$, $q_{\mathcal{F}^{*}|\mathcal{R}}$.
-Upon retrieval of a fragment and fine-alignment to find the fragment-level start index, $*q_{|\mathcal{F}^{*}}$, the global reference start position is obtained as:
+Upon retrieval of a fragment and fine-alignment to find the fragment-level start index, $q_{|\mathcal{F}^{*}}$, the global reference start position is obtained as:
 $q^{*}=q_{|\mathcal{F}^{*}}+q_{\mathcal{F}^{*}|\mathcal{R}}.$ (7)
@@
 This section outlines the setup for evaluating Transformer-DNA baselines, with their recall performance depicted in Fig.
-2.
-We selected three architectures modeling nucleotide sequences: [NT] NucleotideTransformer ($\in\mathbb{R}^{1280}$(Dalla-Torre et al., 2023), [DB2] DNABERT-2 ($\in\mathbb{R}^{768}$(Ji et al., 2021), and [HD] HyenaDNA ($\in\mathbb{R}^{256}$(Nguyen et al., 2023).
+1.
+We selected three architectures modeling nucleotide sequences: [NT] NucleotideTransformer ($\in\mathbb{R}^{1280}$(nucleotide, ), [DB2] DNABERT-2 ($\in\mathbb{R}^{768}$(ji_dnabert_2021, ), and [HD] HyenaDNA ($\in\mathbb{R}^{256}$(nguyen_hyenadna_2023, ).
 Each model employs mean- and max-pooling of token representations for sequence encoding ($2\times 3=6$ baselines total).
 Independent vector stores for each baseline encode fragments from the entire $3$gb genome.
-We sampled $40$K pure reads of varying lengths ($Q\sim\mathcal{U}[25,1000]$) and assessed the average recall for top-$5$, top-$25$, and top-$50$ fragments, as shown in Fig.
-2.
+We sampled $40$K pure reads of varying lengths ($Q\sim\mathcal{U}([25,1000])$) and assessed the average recall for top-$5$, top-$25$, and top-$50$ fragments, as shown in Fig.
+1.
 Overall, while baseline performance is modest, mean-pooling generally outperforms max-pooling, with DB2 (mean-pooled) and HD (max-pooled) as the most effective.
 These two baselines will be contrasted with DNA-ESA in Table 1.
-5 Results and Discussion
 DNA-ESA convergence plots are presented in Figs.
-4(a) and  4(b).
+6(a) and  6(b).
 Model checkpoints are available at OSF.
 In Fig.
-1, representations of short $1,000-$length sequences sampled from sequential (in-order) and gene-specific locations in the reference are visualized in a reduced 2D-UMAP (McInnes et al., 2018).
-The representation space demonstrates desired properties suitable for successfully performing alignment: (a) Sequences sampled in order form a trajectory in the representation space: The loss function described in equation 5 encourages a pair of sequences close to one another to have a short distance between them in the representation space, and pairs further apart to have a larger distance.
+2, representations of short $1,000-$length sequences sampled from sequential (in-order) and gene-specific locations in the reference are visualized in a reduced 2D-UMAP (umap, ).
+The representation space demonstrates desired properties suitable for successfully performing alignment: (a) Sequences sampled in order form a trajectory in the representation space: The loss function described in (6) encourages a pair of sequences close to one another to have a short distance between them in the representation space, and pairs further apart to have a larger distance.
 (b) Representations of sequences drawn from specific gene locations – despite not being close to one another – show gene-centric clustering: The DNA-ESA representation space partially acquires function-level separation as a byproduct of imposing local alignment constraints.
-Codebase is linked.
-5.1 Sequence alignment of ART-simulated reads
+The codebase is linked.
+4 Sequence alignment of ART-simulated reads
 The results from Sec.
 3 demonstrate that even for pure reads, baseline models do not generate adequate representations to perform sequence alignment.
-In this section, DNA-ESA and the two best baselines – DB2, mean and HD, max – are evaluated using reads generated from an external read simulator (ART) – see Huang et al.
-(2012).
+In this section, DNA-ESA and the two best baselines – DB2, mean and HD, max – are evaluated using reads generated from an external read simulator (ART) – see huang_art_2012 .
 ART has served as a reliable benchmark for evaluating other contemporary alignment tools and provides controls to model mutations and variations common in reads generated by Illumina machines.
-Simulator configurations: The different simulation configuration options and settings are listed: (A) Phred quality score $Q_{PH}\in\{[30,60],[60,90]\}$: the likelihood of errors in base-calls of a generated read; (B) Insertion rate $I\in\{0,10^{-2}\}$: the likelihood of adding a base to a random location in a read; (C) Deletion rate $D\in\{0,10^{-2}\}$: the likelihood of deleting a base in the read; (Others): Simulator system: MSv3 [MiSeq]; Read length: $250$.
-Recall configurations: Once the top-K fragments have been retrieved, the first step is to solve equation 3, and for this, we need to compute the SW distance.
+Simulator configurations: The different simulation configuration options and settings are listed: (A) Phred quality score $Q_{PH}$ in one of three $\{[10,30],[30,60],[60,90]\}$ ranges: the likelihood of errors in base-calls of a generated read; (B) Insertion rate $I\in\{0,10^{-2}\}$: the likelihood of adding a base to a random location in a read; (C) Deletion rate $D\in\{0,10^{-2}\}$: the likelihood of deleting a base in the read; (Others): Simulator system: MSv3 [MiSeq]; Read length: $250$.
+Recall configurations: Once the top-K fragments have been retrieved, the first step is to solve (4), and for this, we need to compute the SW distance.
 For all presented results, the settings are: match_score = -2, mismatch_penalty = +1, open_gap_penalty = +0.5, continue_gap_penalty = +0.1.
-After alignment, we get $q^{*}$ – see equation 6 – as the estimated location of a read in the genome.
+After alignment, we get $q^{*}$ – see (7) – as the estimated location of a read in the genome.
 Let $\hat{q}^{*}$ be its true location.
@@
 Hence, the condition for an exact location match: $|q^{*}-\hat{q}^{*}|\leq 2$.
-Distance bound, $d_{SW}\in\{\text{None},1\%\}$: It is well known that short fragments frequently repeat in the genome (Li & Freudenberg, 2014) and $q^{*}$ can correspond to the position of the read in a different location than from where it was sampled.
-In this case, $q^{*}\neq\hat{q}^{*}$, but the SW distance is the minimum possible ($d_{SW}=0$ – see Footnote 2).
-Moreover, when reads have mutations, the reference sequence corresponding to the read is no longer a perfect match; i.e.
-$d_{SW}>0$.
-The best an alignment algorithm could do is to find an exact match for the read leading to an optimal alignment score for that read length ($-2Q$).
-We consider an alignment (with $Q=250$) to be successful if $d_{SW}<0.01$; i.e.
-a mismatch of at most $2$ bases.
-DNA-ESA (ours)       DB2, mean HD, max
-I D Top-$K$ 50 / Chr. $d_{SW}=1\%$ Best Top-K @ 50 / Chr. + $d_{SW}=1\%$
-$Q_{PH}\in[60,90]$
-0 0 95.2 ± 0.63 + 1.6 96.8 ± 0.52 37.4 ± 1.35 14.6 ± 1.01
-0 0.01 95.8 ± 0.59 + 1.3 97.1 ± 0.50 37.4 ± 1.35 14.8 ± 1.01
-0.01 0 95.1 ± 0.63 + 2.1 97.2 ± 0.49 35.9 ± 1.35 14.6 ± 1.01
-0.01 0.01 95.8 ± 0.59 + 1.2 97.0 ± 0.51 34.9 ± 1.34 14.5 ± 1.01
-$Q_{PH}\in[30,60]$
-0 0 95.2 ± 0.63 + 1.4 96.6 ± 0.54 35.4 ± 1.34 13.7 ± 0.98
-0 0.01 95.1 ± 0.64 + 2.0 97.1 ± 0.49 37.3 ± 1.35 14.2 ± 0.99
-0.01 0 94.2 ± 0.68 + 2.1 96.3 ± 0.56 36.0 ± 1.35 13.8 ± 0.98
-0.01 0.01 94.9 ± 0.64 + 1.5 96.4 ± 0.55 36.8 ± 1.35 13.5 ± 0.98
-Table 1: Performance of DNA-ESA with respect to baselines: (with diversity priors – see Def.
-4) The performance of DNA-ESA approeaches that of conventional algorithmic and far exceeds the performance of Transformer-DNA baselines, DB2, mean and HD, max (top performing baselines from Fig.
-2).
-Both DNA-ESA and baselines utilize a dedicated vector store for the entire genome, maintaining a consistent search strategy.
+Distance bound, $d_{SW}\in\{1\%,2\%,5\%\}$: It is well known that short fragments frequently repeat in the genome and $q^{*}$ can correspond to the position of the read in a different location than from where it was sampled mappable ; reinventwillis .
+In this case, $q^{*}\neq\hat{q}^{*}$, but the SW distance is the minimum possible.
+Moreover, when reads have mutations, the reference sequence corresponding to the read is no longer a perfect match.
+We define a bound $d_{SW}$ for classifying whether a (read, retrieved fragment) pair is a successful alignment based on the SW distance between the pair ($v^{*}$ – see Eq.
+4) and the optimal SW distance as a particular read length $Q$:
+$d_{SW}=\frac{v^{*}-mQ}{(n-m)Q},$
+where $m(=-2)$ is the match score and $n(=+1)$ is the mismatch penalty, hyperparameters in the computation of the SW distance.
+We consider an alignment (with $Q=250$) to be successful if the resulting SW distance between the read and the top returned fragment is within $d_{SW}$ of the optimal for that read length.
+A $d_{SW}=2\%$ (default) is equivalent to a mismatch of around $4$ bases in a read of length $250$.
+Model Settings
+ART
+PacBio CCS
+($\mu(v^{*}),\sigma(v^{*})$)
+$I,D=0.01,Q=250$,
+$d_{SW}=2\%$
+Search [10,30] [30,60] [60,90]
+Chr. 2, pbmm2
+$d_{SW}=5\%$
+Transformer Architectures HD (max) K=50 13.20 $\pm$ 1.81 17.01 $\pm$ 2.01 18.03 $\pm$ 2.05 -
+DB2 (mean) 30.21 $\pm$ 2.44 39.19 $\pm$ 2.58 39.20 $\pm$ 2.58 -
+DNA-ESA (ours) 98.40 $\pm$ 0.71 98.64 $\pm$ 0.67 98.60 $\pm$ 0.67 97.5 $\pm$ 0.82
+K=75 98.80 $\pm$ 0.62 99.28 $\pm$ 0.52 98.88 $\pm$ 0.60
+97.5 $\pm$ 0.82
+(-498.0, 3.78)
+Bowtie-2
+- 99.80 $\pm$ 0.19
+99.50$\pm$0.43
+(-497.9, 3.78)
+diff. - <1% <2%
+Table 1: Performance of DNA-ESA with respect to baselines: The performance of DNA-ESA is compared to the DNA-BERT 2 and Hyena-DNA Transformer-based baseline models (with no additional finetuning) – the top performing baselines from Fig.
+1.
+In addition, DNA-ESA is also compared with Bowtie-2 bowtie , a classical aligner.
+Comparisons are conducted across varying qualities (Phred score) of reads generated by ART huang_art_2012 .
+Additionally, an external PacBio CCS dataset of ($Q=250$ length-) reads from Chr.
+2 of the Ashkenazim Trio - Son sample (as determined by pbmm2 minimap ) are aligned using both Bowtie-2 and DNA-ESA.
+All baseline models utilize a dedicated vector store.
 For details on $I,D,Q_{PH},K,d_{SW}$, refer to Recall/Simulator Configurations.
-Performance: Table 1 contains the result of sweeps along these several parameters in addition to a direct comparison to DB2, mean and HD, max baselines, the best-performing baselines on pure reads in Sec.
+Across models and ART-generated datasets, the performance of DNA-ESA supersedes other Transformer-based models and is comparable to Bowtie-2 to within $1\%$.
+With respect to reads from PacBio CCS, DNA-ESA performs with $2\%$ of Bowtie-2 after controlling for the intrinsic quality of retrieved fragments according to the SW distance.
+Mean ($\mu$) and standard deviation ($\sigma$) across the top SW distance values ($v^{*}$) of reads successfully aligned in the case of DNA-ESA and Bowtie-2 models are reported.
+4.1 Performance
+Table 1 reports the performance of DNA-ESA across several read generation configurations – we choose from one of three ranges of Phred score $Q_{PH}$, $\{[10,30],[30,60],[60,90]\}$$I,D=0.01$, $d_{SW}=2\%$ – see Sec.
+4), in addition to a direct comparison to DB2, mean and HD, max baselines (without any further finetuning), the best-performing baselines on pure reads identified in Sec.
+3.
+In addition, we also compare DNA-ESA to Bowtie-2 bowtie , a conventional algorithmic aligner.
+Additionally, models are also evaluated on an external set of reads from PacBio CCN generated on the Ashkenazim Trio (Son) (retrieved from the Genome in a Bottle resource (GIAB) giab ).
+For (ii), the raw reads are $10$ kilobases long and for evaluation we consider random subset of $5000$ reads associated to Chr.
+2333Reads specific to a chromosome are filtered using the pbmm2 (Minimap) package minimap .
+Reads input into the aligner are cropped randomly to a length of $250$ to satisfy the computational requirements of DNA-ESA ($|\mathcal{F}_{i}\bigcap\mathcal{F}_{i+1}|=250$).
+We observe that: (A) DNA-ESA shows strong recall performance of $>99\%$ across a variety of read generation and recall configurations; (B) On cleaner reads, DNA-ESA and Bowtie-2 are within $<1\%$ of each other in recall while controlling for the quality of the retrieved alignments (as determined by the SW distance).
+The results indicate that this constitutes a new state-of-the-art model for Transformer-based sequence alignment.
+Sweeping Top-K and the SW distance bound In Table 2, we report the performance of DNA-ESA across distance bounds $d_{SW}\in\{1\%,2\%,5\%\}$ and a number of recalled fragment settings $K\in\{25,50,75\}$.
+Distance bound $d_{SW}<5\%$ is equivalent to an acceptable mismatch of at most $\sim 8$ bases between the read (length $Q=250$) and recalled reference fragments.
+Recall rates are comparable with different $d_{SW}$ values suggesting that when a read is successfully aligned, it is usually aligned to an objectively best match.
+Decreasing the top-$K$ per chromosome from $50\rightarrow 25$ does not substantially worsen performance ($<1\%$), indicating that the optimal retrievals are usually the closest in the embedding space.
+Additional ablation studies are presented in the Appendix.
+$Q_{PH}\in[30,60]$ $Q_{PH}\in[60,90]$
+$I$ $D$ $d_{SW}<1\%$ $d_{SW}<2\%$ $d_{SW}<5\%$ $d_{SW}<1\%$ $d_{SW}<2\%$ $d_{SW}<5\%$
+Top-25 / chromosome
+0.0 0.0 97±0.94 98.12±0.76 98.56±0.69 97.24±0.91 97.96±0.0.80 98.6±0.69
+0.01 96.88±0.97 98±0.78 98.36±0.73 97.24±0.91 97.96±0.80 98.76±0.64
+0.01 0.0 97.08±0.93 98.04±0.78 98.56±0.69 97±0.94 98.24±0.75 98.92±0.59
+0.01 97.16±0.80 97.96±0.56 99.2±0.52 97.6±0.85 98.12±0.76 98.76±0.62
+Top-50 / chromosome
+0.0 0.0 97.52±0.86 99.04±0.57 99.08±0.57 98.2±0.75 98.88±0.59 99.16±0.52
+0.01 97.48±0.87 99.16±0.52 99.08±0.55 97.8±0.82 98.76±0.62 99.04±0.57
+0.01 0.0 98.24±0.75 98.64±0.67 99.12±0.55 98.28±0.75 98.92±0.59 99.28±0.49
+0.01 97.72±0.83 98.64±0.67 99.36±0.46 97.84±0.92 98.6±0.66 99.0±0.57
+Top-75/ chromosome
+0.0 0.0 98.04±0.78 99.12±0.55 99.0±0.57 98.4±0.71 99.12±0.55 98.4±0.71
+0.01 98.4±0.71 99.04±0.56 99.4±0.46 98.44±0.71 98.84±0.62 99.24±0.52
+0.01 0.0 98.64±0.67 99.0±0.57 99.2±0.52 98.0±0.78 98.72±0.64 99.48±0.43
+0.01 98.68±0.64 99.28±0.49 99.4±0.46 98.48±0.69 98.88±0.59 99.4±0.46
+Table 2: Sequence alignment recall of DNA-ESA sweeping top-K and $d_{SW}$: The various parameters are described in Sec.
 4.
-We observe the following: (A) DNA-ESA demonstrates strong recall of $\sim 97\%$ across a variety of read generation and recall configurations described in Sec.
-5.1; (B) Reads with less noise – high $Q_{PH}$, low $I,D$ – are more often correctly aligned; (C) The $d_{SW}$ bound adds $\sim 1.5-2\%$ in recall (to cover $5\%$ of misalignments) – a $\sim 20\%$ boost.
-This suggests that in the cases of misses, while finding the exact index match, the retrieved reference fragments are still high-quality retrievals that differ by at most $2$ bases; (D) While the precise method of (a) generating reads (ART settings, gapped, long vs.
-short, paired-end vs.
-single-end, etc.), (b) accounting for the variation in the reference (with or without mitochondrial DNA, single- or double haploid, etc.), (c) detecting a successful alignment (duplicates, distance criteria) vary considerably, DNA-ESA approaches the performance reported by mature algorithmic methods of StrobeAlign, BWA-Mem2, and Minimap (Sahlin, 2022; Vasimuddin et al., 2019; Li, 2018).
+DNA-ESA presents a recall of $>99\%$ across several read configurations rivaling contemporary algorithmic models such as Bowtie.
+As expected, performance improves with larger search radius (top-$K$), higher quality reads ($Q_{PH}$) and large distance bound $d_{SW}$.
 4.2 Task transfer from Chromosome 2
-Are these results indicative of DNA-ESA’s adaptability to new genomic sequences, rather than a strict adherence to its training data?
+Are these results indicative of DNA-ESA’s adaptability to new genomic sequences rather than strict adherence to its training data?
 This would suggest the model’s learning to solve the sequence alignment task rather than memorizing the genome.
@@
 Acc. $\uparrow$ F1 $\uparrow$ P $\uparrow$ R $\uparrow$
-Train Reads 2 (seen in training) 3 (populated) Y (populated)
-Chr. 2 Pure 98.0±0.43 98.5±0.39 99.0±0.28 97.5±0.46 98.2±0.41 99.3±0.27
-ART 97.9±0.43 97.8±0.43 97.8±0.43 97.9±0.43 97.6±0.44 98.0±0.43
+Train ART Reads 2 (seen in training) 3 (populated) Y (populated)
+Chr. 2 $Q_{PH}\in[30,60]$ 99.74 $\pm$ 0.29 99.74 $\pm$ 0.29 $9$$\pm$ 0.01 99.49±0.46 98.87±0.59 99.99±0.01
+Chr. 2 $Q_{PH}\in[60,90]$ 99.37 $\pm$ 0.49 99.36 $\pm$ 0.49 $9$$\pm$ 0.01 98.74±0.64 99.12±0.55 99.5±0.43
 Table 3: DNA-ESA Performance on Chr.
@@
 These results support the generalizing tendency of DNA-ESA to align sequences beyond the training reference.
-Species Pure ART
+Species $Q_{PH}\in[30,60]$ $Q_{PH}\in[60,90]$
 Recall Top-K @50 $\uparrow$
-Thermus Aquaticus [All] 100.0 (93.0±0.75 @K=1) 100.0 (95.2±0.63, @K=1)
-Acidobacteriota [All] 100.0 (93.2±0.74, @K=1) 100.0 (94.9±0.64, @K=1)
-Rattus Norwegicus [Chr. 1] 93.3±0.73 95.4±0.62
-Rattus Norwegicus [Chr. 2] 93.2±0.74 96.4±0.55
-Pan Troglodytes [Chr. 2A] 94.5±0.67 97.1±0.51
-Pan Troglodytes [Chr. 2B] 94.3±0.68 97.2±0.50
+Thermus Aquaticus [All] 99.9 $\pm$ 0.01 99.9 $\pm$ 0.01
+Acidobacteriota [All] 99.9 $\pm$ 0.01 99.9 $\pm$ 0.01
+Rattus Norwegicus [Chr. 1] 95.98±1.07 96.48±1.01
+Rattus Norwegicus [Chr. 2] 97.99±0.78 96.49±1.01
+Pan Troglodytes [Chr. 2A] 98.5±0.69 95.74±1.11
+Pan Troglodytes [Chr. 2B] 95.99±1.08 96.25±1.05
 Table 4: Cross-Species Task Transfer with DNA-ESA: Training on Human Chr.
@@
 2, aligns fragments and reads from different species, including Rattus Norwegicus, Pan Troglodytes, Acidabacteriota, and Thermus Aquaticus.
-These species, are evaluated for read alignment recall using pure and ART-generated reads.
+These species, are evaluated for read alignment recall using ART-generated reads.
 The findings, consistent with Table 1, indicate DNA-ESA’s proficiency in modeling DNA sequence structure, beyond simply memorizing training data.
-Top-$K=1$ tests are conducted on species achieving 100% recall with $K=50$.
-Performance: Details on convergence are in Appendix A.
+Performance: Details on convergence are in Appendix 8.
 Table 3 shows that the performance on unseen human chromosomes (3, Y) is similar to the performance reported in Table 1, despite only training DNA-ESA on Chr.
@@
 High Accuracy, Precision, and F1-scores further confirm the method’s specificity at this high recall.
-In Table 3, a trend of decreasing recall across species is observed: Human (2, 3, Y) $>$ Chimpanzee (2A, 2B) (Blue) $>$ Rat (1, 2), with chimpanzees likely performing better due to genetic similarities with humans (Ijdo et al., 1991).
+In Table 4, a trend of decreasing recall across species is observed: Human (2, 3, Y) $>$ Chimpanzee (2A, 2B) (Blue) $>$ Rat (1, 2), with chimpanzees likely performing better due to genetic similarities with humans (originstory, ).
 Even distantly related species like Thermus aquaticus and Acidobacteriota show significant recall, highlighting DNA-ESA’s task transferability beyond simple data memorization.
-5.3 Ablation Studies
-We’ve conducted additional experiments on DNA-ESA, detailed in the Appendix: (A) Sec.
-C.1 covers DNA-ESA’s performance on noisy reads from ART simulators and the PacBio CCS dataset.
-(B) In Sec.
-C.2, we explore its performance under various $d_{SW}$ bounds ({5%,10%}) and top-$K$ settings ({10,20}) per chromosome.
-High performance at lower $K$ values suggests a more effective representation space, while the $d_{SW}$ range assesses fragment retrieval accuracy without exact index matches.
-(C) Sec.
-C.3 describes outcomes when omitting diversity priors in searches across the whole genome (top-$K$), resulting in predictably lower performance.
-(D) Sec.
-C.4 evaluates DNA-ESA’s effectiveness with different read lengths, including those beyond $Q=250$.
-Notably, performance improves with longer reads, demonstrating a Zero-shot effect for lengths not included in training.
-6 Future Work
-While DNA-ESA shows promising recall performance when compared to traditional algorithmic methods, its current computational speed, at 200 reads per minute, is a limiting factor for extensive genomic studies involving millions of reads.
-Despite this, the alignment task’s inherent parallelizability (outlined in [P1]) offers avenues for efficiency improvements.
+5 Limitations and Future Work
+DNA-ESA has the following challenges that we hope to overcome in future generations of the system: the computational speed, at 10K reads per minute, is a limiting factor for extensive genomic studies involving millions of reads.
+Despite this, the alignment task’s inherent parallelizability offers avenues for efficiency improvements.
 Our ongoing and future efforts are directed towards overcoming this time efficiency bottleneck, as detailed in Appendix Sec.
-B.
-We are considering various optimization strategies, including model compilation to speed up inference and enhanced parallelization in vector store searches and fragment-read alignment.
-Additionally, we aim to enhance DNA-ESA’s performance with shorter read lengths, exploring alternative training methods, integrating more diverse read/reference features, and employing data augmentation techniques.
+9.
+We are considering various optimization strategies, including model compilation, to speed up inference and enhanced parallelization in vector store searches and fragment-read alignment.
+Additionally, we aim to enhance DNA-ESA’s performance at shorter read lengths, exploring alternative training methods and integrating more diverse read/reference features.
 6 Concluding Remarks
-Current DNA sequence alignment methods rely on algorithmic approaches that have been refined over decades, incorporating DNA-specific enhancements in both indexing and retrieval.
-We have introduced an alternate data-driven paradigm.
-Employing a Transformer-based DNA-ESA encoder, our approach performs sequence alignment through self-supervised learning using contrastive loss.
-While such methods have previously been used for identifying (approximate) semantic similarity between sequences in written language applications, we demonstrate the surprising ability to find exact overlaps among DNA sequences.
-Our empirical results also show that the model, once trained, embodies the inherent structure of any DNA sequence (up to context length) irrespective of the location or species origin: a foundation model.
-DNA-ESA’s innovation lies in its ability to represent various sequence lengths within a shared vector space, facilitating a ‘flat’ search approach.
-Unlike traditional methods that rely on hierarchical search, DNA-ESA treats shorter reads and longer reference fragments equally, using cosine similarity as a symmetric distance metric.
-This novel approach not only challenges existing paradigms but also enhances the potential for discovering novel DNA sequence representations.
-Mirroring the impact of Sentence Transformers in addressing sequence-level tasks in NLP, DNA-ESA, combined with traditional techniques, could lead to breakthroughs like aligning reads to the Pan Genome, where variations in the reference sequences model the genetic variations across individuals.
-Appendix
-Appendix A Convergence
+Current DNA sequence alignment methods, honed over decades, incorporate DNA-specific enhancements in both indexing and retrieval, relying on algorithmic approaches.
+However, we introduce an alternative data-driven paradigm, employing a Transformer-based DNA-ESA encoder.
+Our method performs sequence alignment through self-supervised learning, utilizing contrastive loss.
+While similar methods have previously found application in identifying approximate semantic similarity in written language, our approach surprisingly excels at identifying exact overlaps among DNA sequences.
+Empirical results underscore the model’s ability to embody the inherent structure of any DNA sequence, irrespective of location or species origin.
+A distinguishing feature of DNA-ESA is its ability to map both shorter reads and longer reference fragments into the same space by employing cosine similarity as a symmetric distance metric.
+This complements existing paradigms that have traditionally used a hierarchical and explicit comparison of base distributions.
+DNA-ESA representations would find utility in other related tasks.
+If, for example, one had to align reads to the Pan Genome liao instead of a single reference, one can populate a single vector store with fragments from all $47$ genome assemblies.
+Then given a read, the nearest fragment to the read can belong to any reference variant.
+Note that in this case, we avoid having to construct a separate index for each of the references as is usually done in conventional methods.
+One can similarly repurpose the representation space – originally intended for Sequence Alignment – to address the complementary task of Genome Assembly.
+We conclude this paper with a setup that provides initial proofpoints under reasonable simplifying assumptions about read lengths and overlaps among the reads.
+The resulting framework will be formalized and benchmarked with reads from actual experiments in our future work.
+6.1 A setup for de-novo Genome Assembly using DNA-ESA
+Consider a set of reads, $R:=\{r_{1},r_{2},\dots,r_{C}\}$ from which we would like to estimate a reference $\mathcal{R}$.
+The de-novo assembly task is to estimate an ordering of the reads in $R$, $\hat{R}:=(r_{i},\dots,r_{j},\dots,r_{k})$ such that a concatenation (after removing overlaps) is approximately the original reference.
+In typical setups, all pairs of reads would need to be compared in order to find those that overlap and construct longer chains of the assembly in an incremental fashion.
+This may be visualized as a complete graph $K_{C}$ with $C$ nodes (reads) and $C^{2}$ edges (constraints).
+To avoid the quadratic complexity, modern techniques such as Overlap-Layout-Consensus (OLC) or de Bruijn Graph (DBG) approaches assembly have adopted hot-start heuristics such as base overlap and shared K-mer counts to reduce the number of pairwise read comparisons.
+The representation space modeled by DNA-ESA admits a comparable assembly setup.
+Fig.
+2 suggested emergent 1D manifolds in the representation space – comprising the embeddings of the reads – $h(r_{1}),h(r_{2}),\dots,h(r_{C})$ – such that the relative positions of fragments along the manifold correlated to the relative locations of the fragments along the reference.
+Therefore, a pairwise read-read comparison within a smaller radius in the embedding space would be sufficient to find reads that overlap.
+And more generally, a walk along the 1D manifold would constitute a near-optimal assembly.
+Figure 5: Motivating the task of de-novo assembly using the distance adjacency matrix of the UMAP-generated $k$-nearest neighbor ($k=10$) network of reads from Thermus Aquaticus: We assume that reads $r_{1},r_{2},\dots,r_{C}$ belonging to a species are embedded into the DNA-ESA embedding space $h(r_{1}),h(r_{2}),\dots,h(r_{C})$ using a model trained on a known reference genome belonging to a different species.
+For this illustrative numerical example we ensure that reads are created with the following properties: first, no read is a substring of any other read (efficient algorithms for the case where reads can be substrings of others will be covered in our future work); second every read has a corresponding read that it overlaps with for each end, and finally, the union of the reads cover the ground-truth reference genome.
+Thus, for every read $r_{i}$ there exist reads $r_{k}$ and $r_{j}$ such that the ordered triplet $(r_{j},r_{i},r_{k})$ constitutes a fragment of the reference.
+In a perfect embedding scenario, the closest neighbors of $r_{i}$ (i.e.
+$k=1$ in an $k$-NN search) in the embedding space should be $r_{j}$ ad $r_{k}$.
+For this ideal case, if a network is constructed where each read (node) is connected to its two nearest neighbors, and the nodes are arranged based on the order in which they appear in the reference chromosome, one would get a bi-diagonal adjacency matrix.
+When ground truth is not known, the adjacency matrix will be a permuted version of a bi-diagonal matrix, but a complete assembly can still be performed by a greedy algorithm that starts with any node and pieces together the neighboring reads in the embedded space in time that grows linearly in the number of reads.
+Simulated reads, each of length $500$ and with random overlaps of lengths drawn uniformly $\mathcal{U}([150,250])$ with its neighboring reads at each end, are sampled from $4$ of the chromosomes in Thermus Aquaticus, and are embedded into the DNA-ESA embedding space trained with a reference human genome.
+The sorted adjacency matrix of the UMAP umap -generated $k$-NN ($k=10$) network is presented above for each chromosome.
+The dark off-diagonal distance values in the $k$-NN adjacency graph with $k=10$ show that the distance properties of actual embeddings of simulated reads from Thermus Aquaticus are almost coincident with that of the ideal case discussed above.
+Moreover, just as in the ideal case, by finding an approximate solution to the Traveling Salesman Problem in the unsorted network, one can find a walk (the assembly) that is close to the original reference.
+The results are presented in Table 5.
+We provide an initial computational implementation of this idea: (a) The k-nearest neighbor (kNN) graph $G(R,d)$, representing the distances between close-by reads, is computed using the density-aware corrections provided by the UMAP library umap : graph adjacency matrices for each of the chromosomes is visualized in Fig.
+5.
+For this figure, the several reads are ordered manually as they appear in the assembly in order to highlight the connectivity pattern; the bright off-diagonal values indicate a chain-like structure corresponding to the observed manifold; (b) A Hamiltonian cycle computed through the graph that has the lowest total distance corresponds to an assembly, $\hat{R}$.
+This involves solving the Traveling Salesman Problem (TSP) – we use the 3/2-approximate Christofides algorithm christofides .
+After a walk across the reads is generated, a second pass through the walk greedily concatenates adjacent reads $(r_{i},r_{i+1})$ by computing the alignment using the SW distance and removing duplicates and overlaps.
+In cases where adjacent reads do not have any alignment to one another – the walk isn’t guaranteed to be perfect since the DNA-ESA model is heuristical and the TSP algorithm is approximate – the current assembly part is stowed away as a contig contig , and a new contig is instantiated with the incoming read.
+The fewer the contigs, the better the manifold in the representation space correlates to the ordering of the true reference.
+Assembly is conducted across $4$ of the shorter chromosomes in the Thermus Aquaticus genome.
+We consider (long) reads of length $Q=500$, which have no indels, and the entire corpus of reads is ensured to have full coverage across each chromosome.
+Each read has an overlap of $\mathcal{U}([150,250])$ bases with another read in the corpus of reads.
+QUAST quast metrics are reported in Table 5.
+Size (in kb) $\uparrow$ No. of Contigs $\downarrow$ N50 (in kb) $\uparrow$ NGA50 (in kb) $\uparrow$
+Thermus Aquaticus NZ_CP010823.1 14.4 1 14.4 14.4
+NZ_CP010824.1 16.6 1 16.6 16.6
+NZ_CP010825.1 69.9 22 6.0 4.4
+NZ_CP010826.1 78.7 15 10.3 10.3
+Table 5: DNA-ESA performance on de-novo assembly of the Thermus Aquaticus genome: The solution to the approximate TSP applied on the DNA-ESA read-read network generates a set of contigs that can be evaluated with respect to the original reference using QUAST quast .
+For the two shorter chrosomomes, the resulting assembly was perfect; the entire reference is encapsulated by one contig.
+Even in the case with longer chromosomes, the number of contigs are still few and of significant length (as indicated by the high N(G)50 score) suggesting a strong correlation between the distance metric in the embedding space and the ordering of the corresponding reads in the assembly.
+N50 refers to the length of a contig (in bases) such that the contigs that exceed this length span 50% of the reference genome.
+NGA50 is a reference-aware N50 metric.
+The fewer contigs, large N(GA)50 scores in comparison to the large size of the chromosomes, when considered together, demonstrate that the reads are indeed ordered in a relative position-aware manner in the DNA-ESA embedding space with respect to a reference making the de-novo assembly task viable.
+{appendices}
+7 Appendix
+8 A. Convergence
 Fig. 6 plot the convergence of the DNA-ESA encoder model discussed in the main text. We see convergence after $\sim$20k steps.
-(a) Trained on the Human Genome
-(b) Trained on the Human Chromosome 2
+((a)) Trained on the Human Genome
+((b)) Trained on the Human Chromosome 2
 Figure 6: DNA-ESA convergence plots. Both plots show the loss pr. step (grey). For clarity, we smooth the loss using a moving average (black).
-Appendix B Complexity of computing alignment
-B.1 Cost of constructing a new representation
+9 B. Complexity of computing alignment
+9.1 Cost of constructing a new representation
 Computing the embedding $\mathcal{E}$ of a sequence of length $F$ using DNA-ESA encoding – a typical Transformer-based attention architecture – has the following computation complexity:
@@
 $\mathcal{O}(F^{2}*d+d^{2}*F)\Rightarrow\mathcal{O}(F^{2})$ (9)
-The $F^{2}$ complexity follows the basic implementation of attention in transformers, but recent efforts (Beltagy et al., 2020; Kitaev et al., 2020) have developed shortcuts to reduce the cost.
-These have already been applied to DNA sequence modeling  (Nguyen et al., 2023).
-B.2 Vector Store Upstream
+The $F^{2}$ complexity follows the basic implementation of attention in transformers, but recent efforts (beltagy_longformer_2020, ; kitaev_reformer_2020, ) have developed shortcuts to reduce the cost.
+These have already been applied to DNA sequence modeling  (nguyen_hyenadna_2023, ).
+9.2 Vector Store Upstream
 Vector store $\mathcal{D}$ is populated once (in bulk) with encoded fragment-length sequences drawn from the entire genome; constant time complexity $C$ to upload $<10M$ vectors.
-B.3 Retrieval Cost
+9.3 Retrieval Cost
 Given a new embedding, $\epsilon(G,K)$ is the cost of retrieving top-$K$ nearest neighbors across the fragment embeddings, where $G$ is the length of the reference genome.
 In modern vector databases, where several hashing techniques such as approximate K-nearest neighbors are used, $\epsilon$ scales logarithmically with $G$.
-This is indeed the key benefit of using such databases instead of a naïve search technique.
-B.3.1 Fine-grained Alignment
+This is indeed the key benefit of using such databases.
+9.4 Fine-grained Alignment
 Existing libraries/algorithms (e.g. the Smith–Waterman algorithm) can identify the alignment between a fragment sequence (of length $F$) and read (of length $Q$) in $\mathcal{O}(FQ)$.
-B.3.2 Total Complexity
+9.5 Total Complexity
 Total complexity involves (a) constructing the representation of a read; (b) querying the vector store; (c) running fine-grained alignment with respect to the $K$ returned reference fragment sequences:
-$\mathcal{O}(F^{2}+FQK+\epsilon(G,K))\Rightarrow\mathcal{O}(F(F+QK)+\epsilon(G,K))\Rightarrow\mathcal{O}(FQK+\epsilon(G,K))$
-Appendix C Ablation studies
-C.1 Performance on high-noise reads
-In Table 4, we report the performance of DNA-ESA on two noisy datasets: (i) ART-generated (MiSeqv3-based) reads of length $Q=250$ with Phred quality score $Q_{PH}\in[10,20],I,D=1\%$ and (ii) reads from PacBio CCN as generated on the Ashkenazim Trio (Son) (retrieved from the Genome in a Bottle resource (GIAB) Zook et al.
-(2016)).
-For (ii), the raw reads are 10 kilobases long and for evaluation we consider random subset of $1000$ reads derived to Chr.
-2.
-Reads specific to a chromosome are filtered using the pbmm2 (Minimap) package Li (2018).
-Reads input into the aligner are cropped to length $250,350$ and $500$.
-Note that in these experiments, we take care to increase the fragment-to-fragment overlap $|\mathcal{F}_{i}\bigcap\mathcal{F}_{i+1}|$ from $250$ to $500$.
-DNA-ESA performance is reported with respect to the Bowtie-2 performance.
-While Bowtie-2 has a $>99.9\%$ recall, DNA-ESA isn’t far behind and is comparable (within $1\%$) to the results on less noisy reads reported in Table 1.
-While the recall is lower, the quality of the recovered fragments is comparable: the average Smith-Waterman distance for all successfully aligned {read, fragment} pairs are similar.
-MiSeqv3 - Q = 250
-(via ART, PH in [10,20], I,D:1%)
-PacBio, CCS - pbmm2, chr2
-(Ashkenazim Trio, Son)
-Model
-Very-high
-Sensitivity
-Alignment
-SW (500)
-Very-high
-Specificity
-Alignment
-SW (500)
-Q=250 Q=350 Q=500
-Bowtie-2 >99.9 490.2
-87.0 ± 0.94
-(–mp 1000
-–np 1000)
-490.4 <+2.0% <+2.8% <+2.5%
-DNA-ESA
-94.4 ± 0.60
-(dSW=30)
-488.3
-89.0 ± 0.87
-(@dSW=15)
-490.6 97.5 ± 0.47 96.5 ± 0.55 97.0 ± 0.51
-Table 4: DNA-ESA performance on noisy reads: Noisy reads are simulated using ART ($Q_{PH}=20$, $I,D=1\%$) and derived from an external PacBio CCS read dump available in GIAB.
-The reads from PacBio are filtered to those derived from Chr.
-2 as determined using the pbmm package Li (2018).
-In this evalaution, DNA-ESA is contextualized with Bowtie-2 Langmead et al.
-(2018), a popular conventional alignment tool.
-On both datasets, DNA-ESA performs slightly worse than Bowtie-2.
-Despite this, the quality of the retrieved fragments – as measured using the SW distance – is very similar.
-Note that in all cases, the performance of DNA-ESA is comparable (within $1\%$) to those reported in Table.
-1.
-C.2 Sweeping Top-K and SW distance bound
-In Table 5, we report the performance of DNA-ESA with larger Smith-Waterman distance bounds $d_{SW}\in\{10\%,20\%\}$ and a smaller number of recalled fragments $K\in\{10,20\}$.
-Distance bound $d_{SW}<10\%$ is equivalent to an acceptable mismatch of at most $\sim 8$ bases between the read (length $Q=250$) and recalled reference fragments.
-We observe that with small $K$, the recall is $>91\%$ (exact index match).
-As the distance bound increases, performance predictably improves; however, the performance gain from $d_{SW}=25$ to $d_{SW}=50$ is small.
-This implies that many retrieved fragments are high-quality retrievals; i.e.
-of high-likelihood to align with the read, and do not benefit from a more generous distance bound to improve recall performance significantly.
-Furthermore, decreasing the top-$K$ per chrosomosome from $20\rightarrow 10$ does not substantially worsen performance ($\sim 2\%$) indicating that the optimal retrievals are usually the closest in the embedding space.
-$Q_{PH}\in[30,60]$ $Q_{PH}\in[60,90]$
-$I$ $D$ Exact $d_{SW}<5\%$ $d_{SW}<10\%$ Exact $d_{SW}<5\%$ $d_{SW}<10\%$
-Top-10 / chromosome
-0.0 0.0 91.5±0.81 96.6±0.54 96.6±0.54 91.9±0.79 96.4±0.55 97.3±0.49
-0.01 92.0±0.79 96.5±0.55 96.8±0.52 92.0±0.79 96.2±0.57 97.6±0.47
-0.01 0.0 91.9±0.79 96.4±0.55 96.5±0.55 92.0±0.79 96.6±0.54 97.4±0.48
-0.01 91.8±0.80 96.3±0.56 97.2±0.50 91.5±0.81 96.3±0.56 97.3±0.49
-Top-20 / chromosome
-0.0 0.0 93.4±0.73 97.5±0.47 98.0±0.43 93.7±0.71 97.7±0.45 97.8±0.44
-0.01 93.2±0.74 97.2±0.49 98.2±0.41 93.6±0.72 97.5±0.47 97.8±0.45
-0.01 0.0 93.0±0.74 97.9±0.44 97.9±0.43 92.6±0.76 97.4±0.48 97.8±0.44
-0.01 93.0±0.74 97.4±0.48 97.8±0.45 93.4±0.72 97.8±0.45 98.1±0.42
-Table 5: Sequence alignment recall of DNA-ESA sweeping top-K and $d_{SW}$: The various parameters are described in Sec.
-5.1.
-DNA-ESA presents a recall of $>97\%$ across several read configurations rivaling contemporary algorithmic models.
-Performance predictably degrades as more noise is introduced into a read.
-Performance improves with larger search radius (top-$K$), higher quality reads ($Q_{PH}$) and large distance bound $d_{SW}$.
-C.3 Without diversity priors in top-K
+$\mathcal{O}(F^{2}+FQK+\epsilon(G,K))\Rightarrow\mathcal{O}(F(F+QK)+\epsilon(G,K))$
+$\Rightarrow\mathcal{O}(FQK+\epsilon(G,K))$
+10 C. Ablation studies
+10.1 Without diversity priors in top-K
 In Table 6, we report the performance without diversity priors used in the retrieval step: i.e.
@@
 $Q_{PH}\in[30,60]$ $Q_{PH}\in[60,90]$
-$I$ $D$ Exact $d_{SW}<5\%$ $d_{SW}<10\%$ Exact $d_{SW}<5\%$ $d_{SW}<10\%$
-Top-100
-0.0 0.0 82.6±1.08 92.0±0.79 93.2±0.74 88.5±0.92 91.9±0.79 93.0±0.74
-0.01 82.3±1.09 91.8±0.80 92.7±0.76 89.3±0.89 93.2±0.74 93.9±0.70
-0.01 0.0 83.3±1.06 92.3±0.77 93.3±0.73 88.7±0.91 92.4±0.77 93.3±0.73
-0.01 82.8±1.08 91.6±0.81 92.5±0.76 88.7±0.91 92.4±0.77 93.2±0.73
-Top-50
-0.0 0.0 81.1±1.11 90.6±0.84 92.0±0.79 86.7±0.97 90.7±0.84 91.9±0.79
-0.01 80.5±1.13 90.4±0.85 91.6±0.80 87.4±0.95 91.7±0.80 92.7±0.76
-0.01 0.0 81.3±1.11 90.8±0.83 92.3±0.78 86.8±0.97 91.0±0.83 92.2±0.78
-0.01 81.2±1.11 90.4±0.85 91.7±0.80 87.1±0.96 91.2±0.82 92.4±0.77
+$I$ $D$ $d_{SW}<1\%$ $d_{SW}<2\%$ $d_{SW}<5\%$ $d_{SW}<1\%$ $d_{SW}<2\%$ $d_{SW}<5\%$
+[…]
+0.0 0.0 98.56 $\pm$ 0.67 97.60 $\pm$ 0.85 98.56 $\pm$ 0.67 97.76 $\pm$ 0.82 97.76 $\pm$ 0.82 97.44 $\pm$ 0.88
+0.01 0.01 98.8 $\pm$ 0.62 98.4 $\pm$ 0.71 98.8 $\pm$ 0.62 99.2 $\pm$ 0.52 98.4 $\pm$ 0.71 98.4 $\pm$ 0.71
+Top-50 global
+0.0 0.0 91.6 $\pm$ 1.49 92.4 $\pm$ 1.43 93.8 $\pm$ 1.31 90.4 $\pm$ 1.58 93.8 $\pm$ 1.31 94.8 $\pm$ 1.21
+0.01 0.01 89.6 $\pm$ 1.64 91.6 $\pm$ 1.49 96 $\pm$ 1.07 94.4 $\pm$ 1.25 94.4 $\pm$ 1.25 93.2 $\pm$ 1.36
 Table 6: Sequence alignment recall of DNA-ESA - without diversity priors: The various parameters are described in Sec.
-5.1.
-DNA-ESA presents a recall of $\approx 90\%$.
-Similar to the result presented in the main text, performance improves with larger search radius (top-$K$), higher quality reads ($Q_{PH}$) and large distance bound ($d_{SW}$).
-C.4 Across read lengths
+4.
+When compared to the main results presented in Table 1, it is evident that the addition of diversity priors results in an improvement of $\sim$ in recall.
+Similar to the result presented in the main text, performance improves with larger search radius in the vector store (top-$K$), higher quality reads ($Q_{PH}$) and large distance bound ($d_{SW}$).
+10.2 Across read lengths
 In Table 7, the performance of DNA-ESA is reported across read lengths.
@@
 Improving the performance at shorter read lengths is the subject of future work.
-Top-K @ 50
-I D $Q=150$ $Q=200$ $Q=250$ $1000500<Q<$ 1000
-$Q_{PH}\in[60,90]$ Pure Reads
-0 0 84.6 ± 1.03 94.9 ± 0.65 96.8 ± 0.52 >97
-0 0.01 85.9 ± 0.99 95.2 ± 0.63 97.1 ± 0.50
-0.01 0 85.8 ± 0.99 94.5 ± 0.66 97.2 ± 0.49
-0.01 0.01 86.2 ± 0.98 95.3 ± 0.63 97.0 ± 0.51
+I D $Q=200$ $Q=225$ $Q=250$
 $Q_{PH}\in[30,60]$
-0 0 84.6 ± 1.03 94.6 ± 0.66 96.6 ± 0.54 >96
-0 0.01 85.2 ± 1.01 95.0 ± 0.64 97.1 ± 0.49
-0.01 0 85.7 ± 1.00 94.5 ± 0.66 96.3 ± 0.56
-0.01 0.01 85.8 ± 0.99 94.4 ± 0.67 96.4 ± 0.55
-Table 7: DNA-ESA recall performance across read lengths: Performance of DNA-ESA is higher for longer reads including those lengths on which the model was not trained ($Q\in\mathcal{U}[150,500]$).
+0 0 99 $\pm$ 0.57 98.6 $\pm$ 0.67 99 $\pm$ 0.57
+0.01 0.01 97.4 $\pm$ 0.88 98.8 $\pm$ 0.62 98.6 $\pm$ 0.67
+$Q_{PH}\in[60,90]$
+0 0 98.6 $\pm$ 0.67 98.6 $\pm$ 0.67 98.9 $\pm$ 0.59
+0.01 0.01 98.6 $\pm$ 0.67 99 $\pm$ 0.57 98.6 $\pm$ 0.67
+Table 7: DNA-ESA recall performance across read lengths: Performance of DNA-ESA is higher for longer reads including those lengths on which the model was not trained ($Q\in\mathcal{U}([150,500])$).
 Shorter reads are more challenging for the model potentially due to replicates found across the reference.
+11 Competing interests
+No competing interest is declared.
+12 Author contributions statement
+VR devised the task and framework and assisted with writing.
+MP aided in writing and bridging NLP and Bioinformatics communities.
+LSB contributed to writing and brainstorming.
+PH and KCE realized the framework and assisted with writing.
+SR ran several experiments to improve model performance.
+LM and TG reviewed existing conventional baselines.

### v5

@@
 Such models have shown early promise in tasks involving classification of short DNA sequences, such as the detection of coding- vs non-coding regions, as well as the identification of enhancer and promoter sequences.
-Performance at sequence classification tasks does not, however, translate to sequence alignment, where it is necessary to conduct a genome-wide search to align every read successfully.
-We address this open problem by framing it as an “Embed-Search-Align” task.
-In this framework, a novel encoder model DNA-ESA generates representations of reads and fragments of the reference, which are projected into a shared vector space where the read-fragment distance is used as a surrogate for alignment.
-In particular, DNA-ESA introduces: (1) Contrastive loss for self-supervised training of DNA sequence representations, facilitating rich sequence-level embeddings, and (2) a DNA vector store to enable search across fragments on a global scale.
-DNA-ESA is 99% accurate when aligning 250-length reads onto a human reference genome of 3 gigabases (single-haploid), rivaling conventional algorithmic sequence alignment methods such as Bowtie and BWA-Mem.
-DNA-ESA far exceeds the performance of $6$ recent DNA-Transformer model baselines such as Nucleotide Transformer, Hyena-DNA, and shows task transfer across chromosomes and species.
+Performance at sequence classification tasks does not, however, translate to sequence alignment, where it is necessary to conduct a genome-wide search to align every read successfully, a significantly longer-range task by comparison.
+We bridge this gap by framing the Sequence Alignment task for Transformer models as an “Embed-Search-Align” task.
+In this framework, a novel Reference-Free DNA Embedding RDE model generates representations of reads and fragments of the reference, which are projected into a shared vector space where the read-fragment distance is used as a surrogate for alignment.
+In particular, we introduce: (1) Contrastive loss for self-supervised training of DNA sequence representations, facilitating rich reference-free, sequence-level embeddings, and (2) a DNA vector store to enable search across fragments on a global scale.
+RDE is 99% accurate when aligning 250-length reads onto a human reference genome of 3 gigabases (single-haploid), rivaling conventional algorithmic sequence alignment methods such as Bowtie and BWA-Mem.
+RDE far exceeds the performance of $6$ recent DNA-Transformer model baselines on ESA such as Nucleotide Transformer, Hyena-DNA, and shows task transfer across chromosomes and species.
 keywords:
 Transformers, DNA Sequence Alignment, Large Language Models, Vector Stores
-1 Introduction
+1 1. Introduction
 Equal Contribution
 Sequence alignment is a central problem in the analysis of sequence data.
+It is used in various genomic analyses, including variant calling, transcriptomics, and epigenomics.
 Many DNA sequencers generate short reads that are only a couple of hundred bases long.
@@
 Moreover, most experiments will generate millions of reads.
-Therefore, the resulting task of aligning reads to genomes is computationally expensive, and many decades of work have led to optimized approaches that can perform these tasks with great efficiency.
-However, many limitations remain with existing genome alignment methods.
-One limitation is that a reference genome usually is an approximation of the sequence from a single individual, and does not capture the variability of genomes across populations.
-Another, is that sequence alignment algorithms may not correctly model sequencing errors, and thus may introduce mistakes when trying to score the best match for a sequence against a genome.
-We asked whether some of these limitations could potentially be overcome if we utilize a new paradigm to align reads to genomes by exploiting the Transformer architectures vaswani_attention_2017 adopted for recent language modeling efforts.
+The resulting task of aligning short reads to large-size genomes is computationally challenging –akin to finding a pin in a haystack– and many decades of work have led to optimized approaches that can perform these tasks with great efficiency.
+Since genomes are sequences where the alphabet consists of only a few symbols ($\{A,T,G,C\}$) they can be considered as Limited Alphabet Languages (LAL).
+We ask whether one can design a new paradigm to align reads to genomes by exploiting the Transformer architectures vaswani_attention_2017 adopted for recent language modeling efforts.
 The applications of the Transformer model – and, more generally “Large Language Models” (LLM) – in Bioinformatics applications are still in their infancy, yet hold substantial promise: Transformer models have demonstrated an unprecedented ability to construct powerful numerical representations of sequential data vit ; wav2vec ; gpt3 .
-We establish a foundation model foundation tailored for DNA sequences, where the vocabulary consists of only a few symbols ($\{A,T,G,C\}$ in this case).
 Numerous DNA Transformer models (ji_dnabert_2021, ; zvyagin_genslms_2022, ; fishman_gena-lm_2023, ; nucleotide, ) have emerged recently, mainly designed for classification tasks in downstream applications.
-However, these models do not explicitly consider a fundamental distinction between Limited Vocabulary Languages (LVL) (such as genomes) and natural languages with large vocabulary.
-In LVLs, there is a one-to-one correspondence between the precise symbol ordering and the underlying “meaning”.
+However, these models do not explicitly consider a fundamental distinction between Limited Alphabet Languages (LAL) (such as genomes) and natural languages with large vocabulary.
+In LALs, there is a one-to-one correspondence between the precise symbol ordering and the underlying “meaning”.
 If a protein is encoded by one amino acid sequence, a slightly different sequence (with a few edits) would encode for a different protein with a distinct functionality in the cells.
-Existing foundation Transformer models, trained on classification tasks, generate sequence embeddings such that their pairwise distances correspond to class separation.
+These models, trained on classification tasks, generate sequence embeddings such that their pairwise distances correspond to class separation.
 As a result, pairs of sequences with very large edit distances between them are mapped to numerical representations that are close by.
-There are, however, tasks such as Sequence Alignment where the objective is quite different: the pairwise representation distance has to closely match the sequence edit distance.
+However, in tasks such as Sequence Alignment the objective is quite different: the pairwise representation distance has to closely match the sequence edit distance.
 Indeed as portrayed in Figure 1, current DNA Transformer models without any further refinement are unable to perform Sequence Alignment tasks: the representation space is simply not designed for such an objective.
-This demonstrates the need for a new model which is able to transfer a precise and computationally expensive distance metric, such as the edit distance, over LVLs to a continuous representation space.
+This demonstrates the need for a new model which is able to transfer a precise and computationally expensive distance metric, such as the edit distance, over LALs to a continuous representation space.
+We establish a foundation model foundation Reference-free DNA Embedding (RDE) – tailored for embedding DNA sequences.
+The model $h_{\Theta}$ (i.e., parameterized by weights $\Theta$) maps any genome subsequence $F$ into an embedding $h_{\Theta}(F)\in\mathcal{R}^{n}$ (e.g., $n=768)$, such that if $F_{1}$ and $F_{2}$ are similar subsequences of differing lengths – for example, $F_{1}$ is any noisy subsequence of $F_{2}$ or $F_{1}$ and $F_{2}$ have significant overlaps – then the $d(h_{\Theta}(F_{1}),h_{\Theta}(F_{2}))$ is small; where $d(\cdot)$ is a distance metric in the embedding space, such as the cosine distance between two vectors.
+Another requirement of such RDE models is that they should generate embeddings that are reference free, i.e.
+once trained, it creates semantically-aware embedding (i.e.
+sequences that are close in edit distance are mapped close to each other) of any given genome subsequences irrespective of the reference genome from which it is sampled.
+Given such a model, alignment of reads can be achieved by a near-neighbor search in the embedding space: only the fragments of genome with embeddings nearest to the embedding of the given read need to be considered.
+Thus, a global search in a giga-bases long sequence is reduced to a local search in a vector space.
 The simplest sequence alignment task applies to single-end [^1] A DNA fragment is ligated to an adapter and then sequenced from one end only. reads, where a sequencer generates a read of length $Q$
@@
 The colors correspond to genes and the position of a particular fragment along a specific gene is coded by its color intensity.
-Figure 2: Illustrating DNA-ESA’s Preservation of Sequence Locality in Embedding Space: In DNA-ESA, the reference genome $\mathcal{R}$ is divided into fragments $\mathcal{F}_{i}$, each represented by an embedding $h(\mathcal{F}_{i})$.
+Figure 2: Illustrating RDE’s Preservation of Sequence Locality in Embedding Space: The reference genome $\mathcal{R}$ is divided into fragments $\mathcal{F}_{i}$, each represented by an embedding $h(\mathcal{F}_{i})$.
 For effective sequence alignment, specific structures are expected in the embedding space: (1) Overlapping fragments $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$ should have proximate embeddings; (2) Consecutive fragments forming a long sequence should correspond to a distinct manifold in the embedding space.
-Subfigures (a) and (b) display this emergent geometry, as visualized using DNA-ESA.
+Subfigures (a) and (b) display this emergent geometry, as visualized using embeddings produced by RDE.
 In both subfigures, the fragment and read representations are jointly visualized in a 2D low-dimensional space constructed by applying UMAP umap to the original embeddings.
@@
 Unlike rule-based frameworks, Transformers implicitly identify syntactic and semantic structures tenney , as evidenced by their performance at various NLP tasks like Sentiment Analysis, Entity Recognition, and Question-Answering (devlin_bert_2019, ; reimers_sentence-bert_2019, ; reviewllm, ).
-The flexibility of this architecture has facilitated its application beyond language processing, including in vision (dosovitskiy_image_2020, ), auditory (audiotransformer, ), and neurological domains (wang_brainbert_2023, ).
+The flexibility of this architecture has facilitated its application beyond language processing, including in vision (dosovitskiy_image_2020), auditory (audiotransformer, ), and neurological domains (wang_brainbert_2023, ).
 In bioinformatics, researchers are now harnessing Transformers to bypass genome-specific hard-coded rules, a development we contribute to with our novel method for Sequence Alignment.
-1.1 Transformer models: Written Language to DNA Sequence Alignment
+1.1 1.1 Transformer models: Written Language to DNA Sequence Alignment
 DNA sequences share remarkable similarities with written language, offering a compelling avenue for the application of Transformer models.
 Like written language, these are sequences generated by a small alphabet of nucleotides $\{A,T,G,C\}$.
-Classical DNA modeling efforts have already accommodated mature encoding and hashing techniques initially developed for written language – such as Suffix trees/arrays and Huffman coding (huffman_method_1952, ; manber_suffix_1993, ) – to successfully parse and compress DNA sequences.
-Furthermore, just as written language contains repeated subsequences (words, phrases) to represent real-world objects, DNA sequences similarly possess repeating “words” and groupings of such words into a “sentence” representing, for example, subsets of genes.
-Within the last few years, several Transformer-based models have been developed for DNA sequence analysis.
+Traditional DNA modeling efforts have already accommodated mature encoding and hashing techniques initially developed for written language – such as Suffix trees/arrays and Huffman coding (huffman_method_1952, ; manber_suffix_1993, ) – to successfully parse and compress DNA sequences.
+Recognizing such similarities, within the last few years, several Transformer-based models have been developed for DNA sequence analysis.
 Notably, DNABERT-2 (ji_dnabert_2021, ; dna2bert, ), Nucleotide Transformer (nucleotide, ), GenSLM (zvyagin_genslms_2022, ), HyenaDNA (nguyen_hyenadna_2023, ) and GENA-LM (fishman_gena-lm_2023, ) have been designed to discern relationships between short genetic fragments and their functions.
@@
 In parallel, DNABERT-2 representations have also been found to cluster in the representation space according to certain types of genetic function.
-Given these advances, a natural question arises: Can these Transformer architectures be readily applied to the task of Sequence Alignment?
+These models, trained on classification tasks, generate sequence embeddings such that their pairwise distances correspond to class separation.
+As a result, pairs of sequences with very large edit distances between them are mapped to numerical representations that are close by.
+However, in tasks such as Sequence Alignment the objective is quite different: the pairwise representation distance has to closely match the sequence edit distance.
+A natural question arises: Can these Transformer architectures be readily applied to the task of Sequence Alignment?
 We delineate the associated challenges as follows:
@@
 Notably, these models exhibit subpar recall performance when aligning typical read lengths of 250.
-2 Our Contributions
+2 2. Our Contributions
 In this paper, we argue that both limitations L1, L2 of Transformer-DNA models can be mitigated by formulating sequence alignment as a vector search-and-retrieval task.
-Our approach is twofold: (A) We introduce a sequence encoder DNA-ESA, trained through self-supervision, to map DNA reads to relevant fragments in a reference sequence within a shared embedding space.
-(B) We leverage a specialized data structure, termed a DNA vector store, as a memory bank [^2] Codebase available here: https://anonymous.4open.science/r/dna2vec-7E4E/.
-This provides efficient access to the entire reference sequence for each read alignment.
+Our approach is twofold: (A) We introduce a sequence encoder Reference-Free DNA Embedding (RDE), trained through self-supervision, to embed DNA subsequences to vectors.
+(B) Next, we formulate a framework, Embed-Search-Align (ESA), that maps reads to reference genome sequence.
+We leverage a specialized data structure, termed a DNA vector store, that enables efficient storage and search in an embedding space [^2] Codebase available here: https://anonymous.4open.science/r/dna2vec-7E4E/: the entire reference genome is sharded into equal-length overlapping fragments whose embeddings are then uploaded to the vector store.
+For each read, top-K closest fragments can then be searched efficiently in time that scales only logarithimically in the number of fragments.
 These strategies have been explored in NLP: (A) Sequence-to-embedding training using contrastive loss has shown improved performance – over explicit pooling methods – at abstractive semantic tasks such as evidence retrieval lewis and semantic text similarity (gao_simcse_2021, ; chen_simple_2020, ).
 (B) Specialized data structures, such as “vector stores” or “vector databases” like FAISS (johnson_billion-scale_2019, ) and Pinecone, use advanced indexing and retrieval algorithms for scalable numerical representation search.
-2.1 Task Definition
+2.1 2.1 RDE: Designing effective sequence representations
+An optimal sequence encoder model $h$ is such that the corresponding embeddings of any read $r$ and reference fragment $\mathcal{F}$$h(r),h(\mathcal{F})$ respectively – obey the following constraints over a pre-determined distance metric $d$:
+$d\{h(r_{j}),h(\mathcal{F}_{i})\}\geq d\{h(r_{j}),h(\mathcal{F}_{j})\},\quad i\neq j$ (3)
+Subscript $j$ serves to indicate that the read $r_{j}$ is correctly aligned to a fragment reference, $\mathcal{F}_{j}$, (that we call a positive {read, fragment} pair).
+Given another arbitrary fragment $\mathcal{F}_{i}$ to which $r_{j}$ has poor alignment, $\{r_{j},\mathcal{F}_{i}\}$ is a negative {read, fragment} pair.
+Observe that these inequalities constitute the only requirements for the encoder.
+As long as the neighborhood of $r_{j}$ in the representation space contains the representation for $\mathcal{F}_{j}$, it will be recovered in the nearest neighbors (top-$K$ set) and alignment will succeed.
+Equality is observed when $r_{j}$ is a repeat sequence matched equally well to more than one fragment.
+This motivates using self-supervision (hadsell_dimensionality_2006, ; chen_simple_2020, ; gao_simcse_2021, ) where we are only concerned about the relative distances between positive and negative (read, reference fragment) pairs.
+2.2 2.2 RDE: Self-supervision and contrastive loss
+A popular choice for sequence learning using self-supervision involves a contrastive loss setup described by  chen_simple_2020 and gao_simcse_2021 : i.e.
+for a read $r$ aligned to reference fragment $\mathcal{F}_{j}$, the loss $l_{r}$ simultaneously minimizes the distance of $h(r)$ to $h(\mathcal{F}_{j})$ and maximizes the distance to a batch of random fragments of size $B-1$:
+$l_{r}=-\log\frac{e^{-d(h(r),h(\mathcal{F}_{j}))/\tau}}{e^{-d(h(r),h(\mathcal{F}_{j}))/\tau}+\sum^{B-1}_{i=1}e^{-d(h(r),h(\mathcal{F}_{i}))/\tau}}.$ (4)
+Here $\tau$ is a tuneable temperature parameter.
+To stabilize the training procedure and reach a non-trivial solution, the encoder applies different dropout masks to the reads and fragments similar to the method described in prior work chen_simple_2020 .
+Similar setups have been shown to work in written language applications, most notably in Sentence Transformers (reimers_sentence-bert_2019, ; gao_simcse_2021, ; muennighoff_mteb_2023, ), which continue to be a strong benchmark for several downstream tasks requiring pre-trained sequence embeddings.
+2.3 2.3 RDE: Encoder implementation
+RDE uses a Transformer-encoder (devlin_bert_2019, ; vaswani_attention_2017, ), comprising $12$ heads and $6$-layers of encoder blocks.
+The size of the vocabulary is $10,000$.
+Batch size $B$ is set to $16$ with gradient accumulation across $16$ steps.
+Generated numerical representations for each read (and fragment) are projected into a high-dimensional vector space, $h(r)\in\mathbb{R}^{1020}$.
+The learning rate is annealed using one-cycle cosine annealing (smith_super-convergence_2019, ), dropout is set to $0.1$, and $\tau=0.05$.
+During training, the reference fragment $|\mathcal{F}_{i}|\sim\mathcal{U}([800,2000])$ and read $|r_{i}|\sim\mathcal{U}([150,500])$ have variable lengths sampled from a uniform distribution.
+Here, $|x|$ denotes the length of $x$.
+To improve model performance on noisy reads, $1-5\%$ of bases are replaced with another random base in $40\%$ of the reads in a batch.
+Shorter sequences were padded to equal the length of the longest sequence in a batch.
+The distance metric used is Cosine Similarity.
+2.4 2.4 ESA: Task Definition
 We formulate the problem of Sequence Alignment as minimizing a sequence alignment function, SA, applied to a read $r$ and a reference sequence $\mathcal{R}$ as
@@
 Ideally, we would like to use $\mathcal{R}$ to get a set of fragments $\{\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{K}\}$ that are a subset of the original reference. Then:
-$v^{*}\approx\min_{\mathcal{F}_{i}\in\{\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{K}\}}\texttt{SA}\left(r,\mathcal{F}_{j}\right).$ (4)
+$v^{*}\approx\min_{\mathcal{F}_{j}\in\{\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{K}\}}\texttt{SA}\left(r,\mathcal{F}_{j}\right).$ (6)
 Here, each $\mathcal{F}_{j}$ is a fragment of $\mathcal{R}$ (i.e., $\mathcal{F}_{j}\subset\mathcal{R}$), and $K$ is the number of these sub-tasks. This approximation is effective under the conditions:
@@
   • (2)
-    There are enough fragments $\mathcal{F}_{i}$ to cover $\mathcal{R}$, i.e. $\cup\mathcal{F}_{j}=\mathcal{R}$;
+    Since a read can originate from any position along the reference, there must be sufficient fragments $\mathcal{F}_{i}$ to cover $\mathcal{R}$, i.e.
+$\cup\mathcal{F}_{j}=\mathcal{R}$.
+From within this set of fragments, we hope to retrieve a subset of fragments containing the optimal fragment for the read;
   • (3)
-    $K$ is significantly smaller than $\frac{N}{Q}$. If $\frac{N}{Q}$ then this amounts to scanning the whole reference.
+    While retrieving a subset of fragments of size $K$, we require $K$ to be significantly smaller than $\frac{N}{Q}$. If $\frac{N}{Q}$, then this amounts to scanning the whole reference.
 Conditions (1) and (2) imply that fragments should be short and numerous enough to cover the reference genome.
@@
 After training, the genome is segmented into overlapping fragments, encoded, and uploaded into the vector store.
-2.2 Designing effective sequence representations
-An optimal sequence encoder model $h$ is such that the corresponding embeddings of any read $r$ and reference fragment $\mathcal{F}$$h(r),h(\mathcal{F})$ respectively – obey the following constraints over a pre-determined distance metric $d$:
-$d\{h(r_{j}),h(\mathcal{F}_{i})\}\geq d\{h(r_{j}),h(\mathcal{F}_{j})\},\quad i\neq j$ (5)
-Here $i$ and $j$ serve to distinguish whether a read is aligned to a particular reference fragment, a positive sample $\{r_{j},\mathcal{F}_{j}\}$, or there is a mismatch (negative sample): $\{r_{j},\mathcal{F}_{i}\}$.
-Observe that these inequalities constitute the only requirements for the encoder.
-As long as the neighborhood of $r_{j}$ in the representation space contains the representation for $\mathcal{F}_{j}$, it will be recovered in the nearest neighbors (top-$K$ set) and alignment will succeed.
-Equality is observed when $r_{j}$ is a repeat sequence matched equally well to more than one fragment.
-This motivates using self-supervision (hadsell_dimensionality_2006, ; chen_simple_2020, ; gao_simcse_2021, ) where we are only concerned about the relative distances between positive and negative (read, reference fragment) pairs.
-2.3 Self-supervision and contrastive loss
-A popular choice for sequence learning using self-supervision involves a contrastive loss setup described by  chen_simple_2020 and gao_simcse_2021 : i.e.
-for a read $r$ aligned to reference fragment $\mathcal{F}_{j}$, the loss $l_{r}$ simultaneously minimizes the distance of $h(r)$ to $h(\mathcal{F}_{j})$ and maximizes the distance to a batch of random fragments of size $B-1$:
-$l_{r}=-\log\frac{e^{-d(h(r),h(\mathcal{F}_{j}))/\tau}}{e^{-d(h(r),h(\mathcal{F}_{j}))/\tau}+\sum^{B-1}_{i=1}e^{-d(h(r),h(\mathcal{F}_{i}))/\tau}}.$ (6)
-Here $\tau$ is a tuneable temperature parameter.
-To stabilize the training procedure and reach a non-trivial solution, the encoder applies different dropout masks to the reads and fragments similar to the method described in prior work simclr .
-Similar setups have been shown to work in written language applications, most notably in Sentence Transformers (reimers_sentence-bert_2019, ; gao_simcse_2021, ; muennighoff_mteb_2023, ), which continue to be a strong benchmark for several downstream tasks requiring pre-trained sequence embeddings.
-2.4 Encoder implementation
-DNA-ESA uses a Transformer-encoder (devlin_bert_2019, ; vaswani_attention_2017, ), comprising $12$ heads and $6$-layers of encoder blocks.
-The size of the vocabulary is $10,000$.
-Batch size $B$ is set to $16$ with gradient accumulation across $16$ steps.
-Generated numerical representations for each read, $h(r)\in\mathbb{R}^{1020}$.
-The learning rate is annealed using one-cycle cosine annealing (smith_super-convergence_2019, ), dropout is set to $0.1$, and $\tau=0.05$.
-During training, the reference fragment $|\mathcal{F}_{i}|\sim\mathcal{U}([800,2000])$ and read $|r_{i}|\sim\mathcal{U}([150,500])$ have variable lengths sampled from a uniform distribution.
-Here, $|x|$ denotes the length of $x$.
-To improve model performance on noisy reads, $1-5\%$ of bases are replaced with another random base in $40\%$ of the reads in a batch.
-Shorter sequences were padded to equal the length of the longest sequence in a batch.
-The distance metric used is Cosine Similarity.
-2.5 Search and retrieval
+2.5 2.5 ESA: Search and retrieval
 Figure 4: System Overview [B] - Inference on a New Read: A read, as per (1) and generated by ART (huang_art_2012, ), is encoded by $h$.
@@
 In our experiments with external read generators (huang_art_2012, ), $Q_{max}=250,|\mathcal{F}_{i}|=1250$.
-Each reference fragment is encoded using the trained DNA-ESA model, and the resulting sequence embeddings ($\in\mathbb{R}^{1020}$) – $3$M vectors for a reference of $3$B nucleotides – are inserted into a Pinecone database.
+Each reference fragment is encoded using the trained RDE model, and the resulting sequence embeddings ($\in\mathbb{R}^{1020}$) – $3$M vectors for a reference of $3$B nucleotides – are inserted into a Pinecone database.
 Once populated with all the fragments, we are ready to perform the alignment.
-Retrieval: Given a read $r$, we project its corresponding DNA-ESA representation into the vector store and retrieve the approximate nearest-$K$ set of reference fragment vectors and the corresponding fragment metadata $\{\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{K}\}$.
+Retrieval: Given a read $r$, we project its corresponding RDE representation into the vector store and retrieve the approximate nearest-$K$ set of reference fragment vectors and the corresponding fragment metadata $\{\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{K}\}$.
 Diversity priors: While the top-$K$ retrieved fragments can be drawn from across the entire vector store (genome), contemporary recommendation systems that use the top-$K$ retrieval setup rank and re-rank top search results (Slate Optimization – see grasshopper ) to ensure rich and diverse recommendations.
 Similarly, we apply a uniform prior wherein every retrieval step selects the top-$K$ per chromosome.
-Fine-Alignment: A standard SW distance library (biopython, ) is used to solve (4), which can be executed concurrently across the $K$-reference fragments.
+Fine-Alignment: A standard SW score library (biopython, ) is used to solve (6), which can be executed concurrently across the $K$-reference fragments.
 Let the optimal fragment be $\mathcal{F}^{*}$.
@@
 $q^{*}=q_{|\mathcal{F}^{*}}+q_{\mathcal{F}^{*}|\mathcal{R}}.$ (7)
-3 Transformer-DNA baselines
+3 3. Transformer-DNA baselines
 This section outlines the setup for evaluating Transformer-DNA baselines, with their recall performance depicted in Fig.
@@
 Independent vector stores for each baseline encode fragments from the entire $3$gb genome.
-We sampled $40$K pure reads of varying lengths ($Q\sim\mathcal{U}([25,1000])$) and assessed the average recall for top-$5$, top-$25$, and top-$50$ fragments, as shown in Fig.
+We sampled $40$K pure reads (reads without noise in comparison to the reference) of varying lengths ($Q\sim\mathcal{U}([25,1000])$) and assessed the average recall for top-$5$, top-$25$, and top-$50$ fragments, as shown in Fig.
 1.
 Overall, while baseline performance is modest, mean-pooling generally outperforms max-pooling, with DB2 (mean-pooled) and HD (max-pooled) as the most effective.
-These two baselines will be contrasted with DNA-ESA in Table 1.
-DNA-ESA convergence plots are presented in Figs.
-6(a) and  6(b).
-Model checkpoints are available at OSF.
+These two baselines will be contrasted with RDE on ESA in Table 1.
+RDE convergence plots are presented in the SM Sec.
+A.
+Model checkpoints are available in OSF osf .
 In Fig.
@@
 The representation space demonstrates desired properties suitable for successfully performing alignment: (a) Sequences sampled in order form a trajectory in the representation space: The loss function described in (4) encourages a pair of sequences close to one another to have a short distance between them in the representation space, and pairs further apart to have a larger distance.
-(b) Representations of sequences drawn from specific gene locations – despite not being close to one another – show gene-centric clustering: The DNA-ESA representation space partially acquires function-level separation as a byproduct of imposing local alignment constraints.
+(b) Representations of sequences drawn from specific gene locations – despite not being close to one another – show gene-centric clustering: The RDE representation space partially acquires function-level separation as a byproduct of imposing local alignment constraints.
 The codebase is linked.
-4 Sequence alignment of ART-simulated reads
+4 4. Sequence alignment of ART-simulated reads
 The results from Sec.
 3 demonstrate that even for pure reads, baseline models do not generate adequate representations to perform sequence alignment.
-In this section, DNA-ESA and the two best baselines – DB2, mean and HD, max – are evaluated using reads generated from an external read simulator (ART) – see huang_art_2012 .
+In this section, RDE and the two best baselines – DB2, mean and HD, max – are evaluated on ESA using reads generated from an external read simulator (ART) – see huang_art_2012 .
 ART has served as a reliable benchmark for evaluating other contemporary alignment tools and provides controls to model mutations and variations common in reads generated by Illumina machines.
 Simulator configurations: The different simulation configuration options and settings are listed: (A) Phred quality score $Q_{PH}$ in one of three $\{[10,30],[30,60],[60,90]\}$ ranges: the likelihood of errors in base-calls of a generated read; (B) Insertion rate $I\in\{0,10^{-2}\}$: the likelihood of adding a base to a random location in a read; (C) Deletion rate $D\in\{0,10^{-2}\}$: the likelihood of deleting a base in the read; (Others): Simulator system: MSv3 [MiSeq]; Read length: $250$.
-Recall configurations: Once the top-K fragments have been retrieved, the first step is to solve (4), and for this, we need to compute the SW distance.
+Recall configurations: Once the top-K fragments have been retrieved, the first step is to solve (6), and for this, we need to compute the SW score.
 For all presented results, the settings are: match_score = -2, mismatch_penalty = +1, open_gap_penalty = +0.5, continue_gap_penalty = +0.1.
@@
 Distance bound, $d_{SW}\in\{1\%,2\%,5\%\}$: It is well known that short fragments frequently repeat in the genome and $q^{*}$ can correspond to the position of the read in a different location than from where it was sampled mappable ; reinventwillis .
-In this case, $q^{*}\neq\hat{q}^{*}$, but the SW distance is the minimum possible.
+In this case, $q^{*}\neq\hat{q}^{*}$, but the SW score is the minimum possible.
 Moreover, when reads have mutations, the reference sequence corresponding to the read is no longer a perfect match.
-We define a bound $d_{SW}$ for classifying whether a (read, retrieved fragment) pair is a successful alignment based on the SW distance between the pair ($v^{*}$ – see Eq.
-4) and the optimal SW distance as a particular read length $Q$:
+We define a bound $d_{SW}$ for classifying whether a (read, retrieved fragment) pair is a successful alignment based on the SW score between the pair ($v^{*}$ – see Eq.
+6) and the optimal SW score as a particular read length $Q$:
 $d_{SW}=\frac{v^{*}-mQ}{(n-m)Q},$
-where $m(=-2)$ is the match score and $n(=+1)$ is the mismatch penalty, hyperparameters in the computation of the SW distance.
-We consider an alignment (with $Q=250$) to be successful if the resulting SW distance between the read and the top returned fragment is within $d_{SW}$ of the optimal for that read length.
+where $m(=-2)$ is the match score and $n(=+1)$ is the mismatch penalty, hyperparameters in the computation of the SW score.
+We consider an alignment (with $Q=250$) to be successful if the resulting SW score between the read and the top returned fragment is within $d_{SW}$ of the optimal for that read length.
 A $d_{SW}=2\%$ (default) is equivalent to a mismatch of around $4$ bases in a read of length $250$.
@@
 DB2 (mean) 30.21 $\pm$ 2.44 39.19 $\pm$ 2.58 39.20 $\pm$ 2.58 -
-DNA-ESA (ours) 98.40 $\pm$ 0.71 98.64 $\pm$ 0.67 98.60 $\pm$ 0.67 97.5 $\pm$ 0.82
+RDE (ours) 98.40 $\pm$ 0.71 98.64 $\pm$ 0.67 98.60 $\pm$ 0.67 97.5 $\pm$ 0.82
 K=75 98.80 $\pm$ 0.62 99.28 $\pm$ 0.52 98.88 $\pm$ 0.60
@@
 diff. - <1% <2%
-Table 1: Performance of DNA-ESA with respect to baselines: The performance of DNA-ESA is compared to the DNA-BERT 2 and Hyena-DNA Transformer-based baseline models (with no additional finetuning) – the top performing baselines from Fig.
+Table 1: Performance of RDE with respect to baselines: The performance of RDE on ESA is compared to the DNA-BERT 2 and Hyena-DNA Transformer-based baseline models (with no additional finetuning) – the top performing baselines from Fig.
 1.
-In addition, DNA-ESA is also compared with Bowtie-2 bowtie , a classical aligner.
+In addition, RDE is also compared with Bowtie-2 bowtie , a classical aligner.
 Comparisons are conducted across varying qualities (Phred score) of reads generated by ART huang_art_2012 .
 Additionally, an external PacBio CCS dataset of ($Q=250$ length-) reads from Chr.
-2 of the Ashkenazim Trio - Son sample (as determined by pbmm2 minimap ) are aligned using both Bowtie-2 and DNA-ESA.
+2 of the Ashkenazim Trio - Son sample (as determined by pbmm2 minimap ) are aligned using both Bowtie-2 and RDE.
 All baseline models utilize a dedicated vector store.
 For details on $I,D,Q_{PH},K,d_{SW}$, refer to Recall/Simulator Configurations.
-Across models and ART-generated datasets, the performance of DNA-ESA supersedes other Transformer-based models and is comparable to Bowtie-2 to within $1\%$.
-With respect to reads from PacBio CCS, DNA-ESA performs with $2\%$ of Bowtie-2 after controlling for the intrinsic quality of retrieved fragments according to the SW distance.
-Mean ($\mu$) and standard deviation ($\sigma$) across the top SW distance values ($v^{*}$) of reads successfully aligned in the case of DNA-ESA and Bowtie-2 models are reported.
-4.1 Performance
-Table 1 reports the performance of DNA-ESA across several read generation configurations – we choose from one of three ranges of Phred score $Q_{PH}$, $\{[10,30],[30,60],[60,90]\}$$I,D=0.01$, $d_{SW}=2\%$ – see Sec.
+Across models and ART-generated datasets, the performance of RDE supersedes other Transformer-based models and is comparable to Bowtie-2 to within $1\%$.
+With respect to reads from PacBio CCS, RDE performs with $2\%$ of Bowtie-2 after controlling for the intrinsic quality of retrieved fragments according to the SW score.
+Mean ($\mu$) and standard deviation ($\sigma$) across the top SW scores($v^{*}$) of reads successfully aligned in the case of RDE and Bowtie-2 models are reported.
+4.1 4.1 Performance
+Table 1 reports the performance of RDE on ESA across several read generation configurations – we choose from one of three ranges of Phred score $Q_{PH}$, $\{[10,30],[30,60],[60,90]\}$$I,D=0.01$, $d_{SW}=2\%$ – see Sec.
 4), in addition to a direct comparison to DB2, mean and HD, max baselines (without any further finetuning), the best-performing baselines on pure reads identified in Sec.
 3.
-In addition, we also compare DNA-ESA to Bowtie-2 bowtie , a conventional algorithmic aligner.
-Additionally, models are also evaluated on an external set of reads from PacBio CCN generated on the Ashkenazim Trio (Son) (retrieved from the Genome in a Bottle resource (GIAB) giab ).
+In addition, we also compare RDE to Bowtie-2 bowtie , a conventional algorithmic aligner.
+Additionally, models are also evaluated on an external set of reads from PacBio CCS generated on the Ashkenazim Trio (Son) (retrieved from the Genome in a Bottle resource (GIAB) giab ).
 For (ii), the raw reads are $10$ kilobases long and for evaluation we consider random subset of $5000$ reads associated to Chr.
-2333Reads specific to a chromosome are filtered using the pbmm2 (Minimap) package minimap .
-Reads input into the aligner are cropped randomly to a length of $250$ to satisfy the computational requirements of DNA-ESA ($|\mathcal{F}_{i}\bigcap\mathcal{F}_{i+1}|=250$).
-We observe that: (A) DNA-ESA shows strong recall performance of $>99\%$ across a variety of read generation and recall configurations; (B) On cleaner reads, DNA-ESA and Bowtie-2 are within $<1\%$ of each other in recall while controlling for the quality of the retrieved alignments (as determined by the SW distance).
+2.
+Reads input into the aligner are cropped randomly to a length of $250$ to satisfy the computational requirements of RDE ($|\mathcal{F}_{i}\bigcap\mathcal{F}_{i+1}|=250$).
+We observe that: (A) RDE shows strong recall performance of $>99\%$ across a variety of read generation and recall configurations; (B) On cleaner reads, RDE and Bowtie-2 are within $<1\%$ of each other in recall while controlling for the quality of the retrieved alignments (as determined by the SW score).
 The results indicate that this constitutes a new state-of-the-art model for Transformer-based sequence alignment.
-Sweeping Top-K and the SW distance bound In Table 2, we report the performance of DNA-ESA across distance bounds $d_{SW}\in\{1\%,2\%,5\%\}$ and a number of recalled fragment settings $K\in\{25,50,75\}$.
+Sweeping Top-K and the SW distance bound In Table 2, we report the performance of RDE across distance bounds $d_{SW}\in\{1\%,2\%,5\%\}$ and a number of recalled fragment settings $K\in\{25,50,75\}$.
 Distance bound $d_{SW}<5\%$ is equivalent to an acceptable mismatch of at most $\sim 8$ bases between the read (length $Q=250$) and recalled reference fragments.
@@
 Decreasing the top-$K$ per chromosome from $50\rightarrow 25$ does not substantially worsen performance ($<1\%$), indicating that the optimal retrievals are usually the closest in the embedding space.
-Additional ablation studies are presented in the Appendix.
+Several additional experiments are presented in the SI.
 $Q_{PH}\in[30,60]$ $Q_{PH}\in[60,90]$
@@
 0.01 98.68±0.64 99.28±0.49 99.4±0.46 98.48±0.69 98.88±0.59 99.4±0.46
-Table 2: Sequence alignment recall of DNA-ESA sweeping top-K and $d_{SW}$: The various parameters are described in Sec.
+Table 2: Sequence alignment recall of RDE sweeping top-K and $d_{SW}$: The various parameters are described in Sec.
 4.
-DNA-ESA presents a recall of $>99\%$ across several read configurations rivaling contemporary algorithmic models such as Bowtie.
+RDE presents a recall of $>99\%$ across several read configurations rivaling contemporary algorithmic models such as Bowtie.
 As expected, performance improves with larger search radius (top-$K$), higher quality reads ($Q_{PH}$) and large distance bound $d_{SW}$.
-4.2 Task transfer from Chromosome 2
-Are these results indicative of DNA-ESA’s adaptability to new genomic sequences rather than strict adherence to its training data?
-This would suggest the model’s learning to solve the sequence alignment task rather than memorizing the genome.
-[…]
-Acc. $\uparrow$ F1 $\uparrow$ P $\uparrow$ R $\uparrow$
-Train ART Reads 2 (seen in training) 3 (populated) Y (populated)
-Chr. 2 $Q_{PH}\in[30,60]$ 99.74 $\pm$ 0.29 99.74 $\pm$ 0.29 $9$$\pm$ 0.01 99.49±0.46 98.87±0.59 99.99±0.01
-Chr. 2 $Q_{PH}\in[60,90]$ 99.37 $\pm$ 0.49 99.36 $\pm$ 0.49 $9$$\pm$ 0.01 98.74±0.64 99.12±0.55 99.5±0.43
-Table 3: DNA-ESA Performance on Chr.
-2 and Task Transfer to Chr.
-3, Y: DNA-ESA, trained on Chr.
-2, is evaluated for aligning positive samples from Chr.
-2 and negative samples from Chr.
-3.
-Anticipated failure to align Chr.
-3 reads (false positives) is due to significant SW distance.
-Metrics reported include weighted accuracy (Acc.), precision (P), F1-score (F1), and recall (R), for comparison with Table 1.
-[Grey] For Chr.
-3 and Y, separate vector stores are populated using DNA-ESA representations learned using Chr.
-2 {read,fragment} pairs.
-These results support the generalizing tendency of DNA-ESA to align sequences beyond the training reference.
-Species $Q_{PH}\in[30,60]$ $Q_{PH}\in[60,90]$
-Recall Top-K @50 $\uparrow$
-Thermus Aquaticus [All] 99.9 $\pm$ 0.01 99.9 $\pm$ 0.01
-Acidobacteriota [All] 99.9 $\pm$ 0.01 99.9 $\pm$ 0.01
-Rattus Norwegicus [Chr. 1] 95.98±1.07 96.48±1.01
-Rattus Norwegicus [Chr. 2] 97.99±0.78 96.49±1.01
-Pan Troglodytes [Chr. 2A] 98.5±0.69 95.74±1.11
-Pan Troglodytes [Chr. 2B] 95.99±1.08 96.25±1.05
-Table 4: Cross-Species Task Transfer with DNA-ESA: Training on Human Chr.
-2, Testing on Diverse Species: DNA-ESA, trained on human Chr.
-2, aligns fragments and reads from different species, including Rattus Norwegicus, Pan Troglodytes, Acidabacteriota, and Thermus Aquaticus.
-These species, are evaluated for read alignment recall using ART-generated reads.
-The findings, consistent with Table 1, indicate DNA-ESA’s proficiency in modeling DNA sequence structure, beyond simply memorizing training data.
-Performance: Details on convergence are in Appendix 8.
-Table 3 shows that the performance on unseen human chromosomes (3, Y) is similar to the performance reported in Table 1, despite only training DNA-ESA on Chr.
-2.
-This suggests DNA-ESA’s ability to generalize sequence alignment across chromosomes with different compositions.
-High Accuracy, Precision, and F1-scores further confirm the method’s specificity at this high recall.
-In Table 4, a trend of decreasing recall across species is observed: Human (2, 3, Y) $>$ Chimpanzee (2A, 2B) (Blue) $>$ Rat (1, 2), with chimpanzees likely performing better due to genetic similarities with humans (originstory, ).
-Even distantly related species like Thermus aquaticus and Acidobacteriota show significant recall, highlighting DNA-ESA’s task transferability beyond simple data memorization.
-5 Limitations and Future Work
-DNA-ESA has the following challenges that we hope to overcome in future generations of the system: the computational speed, at 10K reads per minute, is a limiting factor for extensive genomic studies involving millions of reads.
-Despite this, the alignment task’s inherent parallelizability offers avenues for efficiency improvements.
-Our ongoing and future efforts are directed towards overcoming this time efficiency bottleneck, as detailed in Appendix Sec.
-9.
-We are considering various optimization strategies, including model compilation, to speed up inference and enhanced parallelization in vector store searches and fragment-read alignment.
-Additionally, we aim to enhance DNA-ESA’s performance at shorter read lengths, exploring alternative training methods and integrating more diverse read/reference features.
-6 Concluding Remarks
-Current DNA sequence alignment methods, honed over decades, incorporate DNA-specific enhancements in both indexing and retrieval, relying on algorithmic approaches.
-However, we introduce an alternative data-driven paradigm, employing a Transformer-based DNA-ESA encoder.
-Our method performs sequence alignment through self-supervised learning, utilizing contrastive loss.
-While similar methods have previously found application in identifying approximate semantic similarity in written language, our approach surprisingly excels at identifying exact overlaps among DNA sequences.
-Empirical results underscore the model’s ability to embody the inherent structure of any DNA sequence, irrespective of location or species origin.
-A distinguishing feature of DNA-ESA is its ability to map both shorter reads and longer reference fragments into the same space by employing cosine similarity as a symmetric distance metric.
-This complements existing paradigms that have traditionally used a hierarchical and explicit comparison of base distributions.
-DNA-ESA representations would find utility in other related tasks.
-If, for example, one had to align reads to the Pan Genome liao instead of a single reference, one can populate a single vector store with fragments from all $47$ genome assemblies.
-Then given a read, the nearest fragment to the read can belong to any reference variant.
-Note that in this case, we avoid having to construct a separate index for each of the references as is usually done in conventional methods.
-One can similarly repurpose the representation space – originally intended for Sequence Alignment – to address the complementary task of Genome Assembly.
-We conclude this paper with a setup that provides initial proofpoints under reasonable simplifying assumptions about read lengths and overlaps among the reads.
-The resulting framework will be formalized and benchmarked with reads from actual experiments in our future work.
-6.1 A setup for de-novo Genome Assembly using DNA-ESA
-Consider a set of reads, $R:=\{r_{1},r_{2},\dots,r_{C}\}$ from which we would like to estimate a reference $\mathcal{R}$.
-The de-novo assembly task is to estimate an ordering of the reads in $R$, $\hat{R}:=(r_{i},\dots,r_{j},\dots,r_{k})$ such that a concatenation (after removing overlaps) is approximately the original reference.
-In typical setups, all pairs of reads would need to be compared in order to find those that overlap and construct longer chains of the assembly in an incremental fashion.
-This may be visualized as a complete graph $K_{C}$ with $C$ nodes (reads) and $C^{2}$ edges (constraints).
-To avoid the quadratic complexity, modern techniques such as Overlap-Layout-Consensus (OLC) or de Bruijn Graph (DBG) approaches assembly have adopted hot-start heuristics such as base overlap and shared K-mer counts to reduce the number of pairwise read comparisons.
-The representation space modeled by DNA-ESA admits a comparable assembly setup.
-Fig.
-2 suggested emergent 1D manifolds in the representation space – comprising the embeddings of the reads – $h(r_{1}),h(r_{2}),\dots,h(r_{C})$ – such that the relative positions of fragments along the manifold correlated to the relative locations of the fragments along the reference.
-Therefore, a pairwise read-read comparison within a smaller radius in the embedding space would be sufficient to find reads that overlap.
-And more generally, a walk along the 1D manifold would constitute a near-optimal assembly.
-Figure 5: Motivating the task of de-novo assembly using the distance adjacency matrix of the UMAP-generated $k$-nearest neighbor ($k=10$) network of reads from Thermus Aquaticus: We assume that reads $r_{1},r_{2},\dots,r_{C}$ belonging to a species are embedded into the DNA-ESA embedding space $h(r_{1}),h(r_{2}),\dots,h(r_{C})$ using a model trained on a known reference genome belonging to a different species.
-For this illustrative numerical example we ensure that reads are created with the following properties: first, no read is a substring of any other read (efficient algorithms for the case where reads can be substrings of others will be covered in our future work); second every read has a corresponding read that it overlaps with for each end, and finally, the union of the reads cover the ground-truth reference genome.
-Thus, for every read $r_{i}$ there exist reads $r_{k}$ and $r_{j}$ such that the ordered triplet $(r_{j},r_{i},r_{k})$ constitutes a fragment of the reference.
-In a perfect embedding scenario, the closest neighbors of $r_{i}$ (i.e.
-$k=1$ in an $k$-NN search) in the embedding space should be $r_{j}$ ad $r_{k}$.
-For this ideal case, if a network is constructed where each read (node) is connected to its two nearest neighbors, and the nodes are arranged based on the order in which they appear in the reference chromosome, one would get a bi-diagonal adjacency matrix.
-When ground truth is not known, the adjacency matrix will be a permuted version of a bi-diagonal matrix, but a complete assembly can still be performed by a greedy algorithm that starts with any node and pieces together the neighboring reads in the embedded space in time that grows linearly in the number of reads.
-Simulated reads, each of length $500$ and with random overlaps of lengths drawn uniformly $\mathcal{U}([150,250])$ with its neighboring reads at each end, are sampled from $4$ of the chromosomes in Thermus Aquaticus, and are embedded into the DNA-ESA embedding space trained with a reference human genome.
-The sorted adjacency matrix of the UMAP umap -generated $k$-NN ($k=10$) network is presented above for each chromosome.
-The dark off-diagonal distance values in the $k$-NN adjacency graph with $k=10$ show that the distance properties of actual embeddings of simulated reads from Thermus Aquaticus are almost coincident with that of the ideal case discussed above.
-Moreover, just as in the ideal case, by finding an approximate solution to the Traveling Salesman Problem in the unsorted network, one can find a walk (the assembly) that is close to the original reference.
-The results are presented in Table 5.
-We provide an initial computational implementation of this idea: (a) The k-nearest neighbor (kNN) graph $G(R,d)$, representing the distances between close-by reads, is computed using the density-aware corrections provided by the UMAP library umap : graph adjacency matrices for each of the chromosomes is visualized in Fig.
-5.
-For this figure, the several reads are ordered manually as they appear in the assembly in order to highlight the connectivity pattern; the bright off-diagonal values indicate a chain-like structure corresponding to the observed manifold; (b) A Hamiltonian cycle computed through the graph that has the lowest total distance corresponds to an assembly, $\hat{R}$.
-This involves solving the Traveling Salesman Problem (TSP) – we use the 3/2-approximate Christofides algorithm christofides .
-After a walk across the reads is generated, a second pass through the walk greedily concatenates adjacent reads $(r_{i},r_{i+1})$ by computing the alignment using the SW distance and removing duplicates and overlaps.
-In cases where adjacent reads do not have any alignment to one another – the walk isn’t guaranteed to be perfect since the DNA-ESA model is heuristical and the TSP algorithm is approximate – the current assembly part is stowed away as a contig contig , and a new contig is instantiated with the incoming read.
-The fewer the contigs, the better the manifold in the representation space correlates to the ordering of the true reference.
-Assembly is conducted across $4$ of the shorter chromosomes in the Thermus Aquaticus genome.
-We consider (long) reads of length $Q=500$, which have no indels, and the entire corpus of reads is ensured to have full coverage across each chromosome.
-Each read has an overlap of $\mathcal{U}([150,250])$ bases with another read in the corpus of reads.
-QUAST quast metrics are reported in Table 5.
-Size (in kb) $\uparrow$ No. of Contigs $\downarrow$ N50 (in kb) $\uparrow$ NGA50 (in kb) $\uparrow$
-Thermus Aquaticus NZ_CP010823.1 14.4 1 14.4 14.4
-NZ_CP010824.1 16.6 1 16.6 16.6
-NZ_CP010825.1 69.9 22 6.0 4.4
-NZ_CP010826.1 78.7 15 10.3 10.3
-Table 5: DNA-ESA performance on de-novo assembly of the Thermus Aquaticus genome: The solution to the approximate TSP applied on the DNA-ESA read-read network generates a set of contigs that can be evaluated with respect to the original reference using QUAST quast .
-For the two shorter chrosomomes, the resulting assembly was perfect; the entire reference is encapsulated by one contig.
-Even in the case with longer chromosomes, the number of contigs are still few and of significant length (as indicated by the high N(G)50 score) suggesting a strong correlation between the distance metric in the embedding space and the ordering of the corresponding reads in the assembly.
-N50 refers to the length of a contig (in bases) such that the contigs that exceed this length span 50% of the reference genome.
-NGA50 is a reference-aware N50 metric.
-The fewer contigs, large N(GA)50 scores in comparison to the large size of the chromosomes, when considered together, demonstrate that the reads are indeed ordered in a relative position-aware manner in the DNA-ESA embedding space with respect to a reference making the de-novo assembly task viable.
-{appendices}
-7 Appendix
-8 A. Convergence
-Fig. 6 plot the convergence of the DNA-ESA encoder model discussed in the main text. We see convergence after $\sim$20k steps.
-((a)) Trained on the Human Genome
-((b)) Trained on the Human Chromosome 2
-Figure 6: DNA-ESA convergence plots. Both plots show the loss pr. step (grey). For clarity, we smooth the loss using a moving average (black).
-9 B. Complexity of computing alignment
-9.1 Cost of constructing a new representation
-Computing the embedding $\mathcal{E}$ of a sequence of length $F$ using DNA-ESA encoding – a typical Transformer-based attention architecture – has the following computation complexity:
-$\mathcal{O}(LH*(F^{2}*d+d^{2}*F))\Rightarrow\mathcal{O}(F^{2}*d+d^{2}*F)$ (8)
-Where $d$ is the embedding dimension of the model, $L$ is the number of layers in the Transformer and $H$ is the number of heads per layer.
-As $d$ is a controllable parameter for the model, we can further simplify:
-$\mathcal{O}(F^{2}*d+d^{2}*F)\Rightarrow\mathcal{O}(F^{2})$ (9)
-The $F^{2}$ complexity follows the basic implementation of attention in transformers, but recent efforts (beltagy_longformer_2020, ; kitaev_reformer_2020, ) have developed shortcuts to reduce the cost.
-These have already been applied to DNA sequence modeling  (nguyen_hyenadna_2023, ).
-9.2 Vector Store Upstream
-Vector store $\mathcal{D}$ is populated once (in bulk) with encoded fragment-length sequences drawn from the entire genome; constant time complexity $C$ to upload $<10M$ vectors.
-9.3 Retrieval Cost
-Given a new embedding, $\epsilon(G,K)$ is the cost of retrieving top-$K$ nearest neighbors across the fragment embeddings, where $G$ is the length of the reference genome.
-In modern vector databases, where several hashing techniques such as approximate K-nearest neighbors are used, $\epsilon$ scales logarithmically with $G$.
-This is indeed the key benefit of using such databases.
-9.4 Fine-grained Alignment
-Existing libraries/algorithms (e.g. the Smith–Waterman algorithm) can identify the alignment between a fragment sequence (of length $F$) and read (of length $Q$) in $\mathcal{O}(FQ)$.
-9.5 Total Complexity
-Total complexity involves (a) constructing the representation of a read; (b) querying the vector store; (c) running fine-grained alignment with respect to the $K$ returned reference fragment sequences:
-$\mathcal{O}(F^{2}+FQK+\epsilon(G,K))\Rightarrow\mathcal{O}(F(F+QK)+\epsilon(G,K))$
-$\Rightarrow\mathcal{O}(FQK+\epsilon(G,K))$
-10 C. Ablation studies
-10.1 Without diversity priors in top-K
-In Table 6, we report the performance without diversity priors used in the retrieval step: i.e.
-the nearest-K neighbors in the embedding space are selected from the entire set of fragments spanning the genome rather than uniformly sampling from each chromosome.
-The performance predicably falls in comparison to those reported in Table 1 since fewer fragments scattered unevenly across the different chromosomes are being retrieved per read.
-$Q_{PH}\in[30,60]$ $Q_{PH}\in[60,90]$
-$I$ $D$ $d_{SW}<1\%$ $d_{SW}<2\%$ $d_{SW}<5\%$ $d_{SW}<1\%$ $d_{SW}<2\%$ $d_{SW}<5\%$
-[…]
-0.0 0.0 98.56 $\pm$ 0.67 97.60 $\pm$ 0.85 98.56 $\pm$ 0.67 97.76 $\pm$ 0.82 97.76 $\pm$ 0.82 97.44 $\pm$ 0.88
-0.01 0.01 98.8 $\pm$ 0.62 98.4 $\pm$ 0.71 98.8 $\pm$ 0.62 99.2 $\pm$ 0.52 98.4 $\pm$ 0.71 98.4 $\pm$ 0.71
-Top-50 global
-0.0 0.0 91.6 $\pm$ 1.49 92.4 $\pm$ 1.43 93.8 $\pm$ 1.31 90.4 $\pm$ 1.58 93.8 $\pm$ 1.31 94.8 $\pm$ 1.21
-0.01 0.01 89.6 $\pm$ 1.64 91.6 $\pm$ 1.49 96 $\pm$ 1.07 94.4 $\pm$ 1.25 94.4 $\pm$ 1.25 93.2 $\pm$ 1.36
-Table 6: Sequence alignment recall of DNA-ESA - without diversity priors: The various parameters are described in Sec.
-4.
-When compared to the main results presented in Table 1, it is evident that the addition of diversity priors results in an improvement of $\sim$ in recall.
-Similar to the result presented in the main text, performance improves with larger search radius in the vector store (top-$K$), higher quality reads ($Q_{PH}$) and large distance bound ($d_{SW}$).
-10.2 Across read lengths
-In Table 7, the performance of DNA-ESA is reported across read lengths.
-We observe Zero-shot performance at longer read lengths: The model performs better at longer read lengths (even exceeding the read length bound established during training $\mathcal{U}[150,500]$); while evaluating for longer reads, we make sure to guarantee that the reads exist as subsequences of fragments.
-Improving the performance at shorter read lengths is the subject of future work.
-I D $Q=200$ $Q=225$ $Q=250$
-$Q_{PH}\in[30,60]$
-0 0 99 $\pm$ 0.57 98.6 $\pm$ 0.67 99 $\pm$ 0.57
-0.01 0.01 97.4 $\pm$ 0.88 98.8 $\pm$ 0.62 98.6 $\pm$ 0.67
-$Q_{PH}\in[60,90]$
-0 0 98.6 $\pm$ 0.67 98.6 $\pm$ 0.67 98.9 $\pm$ 0.59
-0.01 0.01 98.6 $\pm$ 0.67 99 $\pm$ 0.57 98.6 $\pm$ 0.67
-Table 7: DNA-ESA recall performance across read lengths: Performance of DNA-ESA is higher for longer reads including those lengths on which the model was not trained ($Q\in\mathcal{U}([150,500])$).
-Shorter reads are more challenging for the model potentially due to replicates found across the reference.
-11 Competing interests
-No competing interest is declared.
-12 Author contributions statement
-VR devised the task and framework and assisted with writing.
-MP aided in writing and bridging NLP and Bioinformatics communities.
-LSB contributed to writing and brainstorming.
-PH and KCE realized the framework and assisted with writing.
-SR ran several experiments to improve model performance.
-LM and TG reviewed existing conventional baselines.
+5 5. Limitations and Future Work
+While our first-generation reference-free DNA embedding (RDE) models provide alignment performance that almost matches the performance of traditional aligners (refined over decades), there is room for considerable improvement.
+First, our RDE models have been trained using samples that span only around 2% of the human genome as shown in the Sec-A of the Supporting Material(SM).
+One needs to explore if the embedding properties of our models can be further improved by diversifying and augmenting training data and optimizing over the model parameters.
+The validation of RDE models can be performed only through various genomic tasks.
+In terms of the task of alignment, we plan to explore the following: (i) The current off-the-shelf implementation has a speed of around 10K reads per minute; while the existing aligners can achieve close to 1M reads per minute by trading off alignment accuracy.
+As detailed in SM Sec.
+B, we are considering various optimization strategies, including model compilation, to speed up inference and enhanced parallelization in vector store searches and fragment-read alignment.
+We believe ESA can be made much faster; (ii) Optimization of the training of the RDE models to improve performance for shorter reads.
+As shown in see Table 2, Page 3 of SM the existing model performs well on longer reads; we want to develop RDE models that specialize to aligning shorter reads; and (iii) quantifying how embeddings change as subsequences are edited; this will help in better alignment accuracy.
+It will also help in correlating the geometry of the embedding space to the structure of DNA sequences.
+6 6. Concluding Remarks
+We have introduced a novel Reference-Free DNA Embedding (RDE) framework: it creates an embedding of any given DNA subsequence irrespective of the reference genome from which it is sampled.
+Such models capture similarity among genome subsequences in a purely data-driven manner (without explicitly computing edit-distance measures) and embeds DNA sequences of differing lengths into the same embedding space to facilitate search and assembly.
+Once such RDE models are constructed their expressive power and limitations can be assessed by performing multiple challenging genomic tasks in a flexible manner across different species.
+As a first validation and application of RDE, we developed an Embed-Align-Search (ESA) framework where alignment of reads to a reference genome is achieved by a near-neighbor search in the embedding space: only the fragments of reference genome with embeddings nearest to the embedding of the given read need to be considered.
+Thus, a global search in a giga-bases long reference genome sequence is reduced to a local search in a vector space.
+Detailed experiments showed that the performance of even this first-generation RDE methodology is comparable to that of traditional aligners, which have been refined over decades.
+Future differentiating application of RDE to the alignment task might lie in aligning reads to graph- or pan-genomes liao , where the reference genome has multiple variants.
+Indexing such pangenomes is a difficult task in traditional aligners; however, in ESA one can store the RDE embeddings of the different fragments in the pangenome, and then perform a near-neighbor search in the embedding space.
+Thus, the ESA methodology does not have to be significantly altered, and alignment of reads to pangenomes constitutes an important future task.
+Going forward we aim to answer the following important questions:
+  • How reference-free is our RDE model?
+To benchmark the generalizability of our RDE model we have performed several cross-species alignment tasks as summarized in Tables 3 and 4 in the SM.
+In particular, we have used an RDE trained using human genome to align reads to reference genomes belonging to a wide range of species.
+The results are encouraging and the design of universal RDEs –that would generate effective embeddings of subsequences sampled from the genome of every species – is another direction of future research.
+  • Is the RDE capable of generalizing to other tasks like de novo Genome Assembly?
+One can also harness the RDE space to address the complementary task of de novo Genome Assembly, where given a set of reads one assembles a genome by stitching together the reads, without using any reference.
+Genome assembly is considered to be a much more challenging task than alignment.
+In SM Sec.
+E, we provide initial proofpoints of a genome assembly framework, under reasonable simplifying assumptions about read lengths and overlaps among the reads.
+We show how the RDE model – trained on exemplars from human genome – can effectively embed simulated reads from a very different species, Thermus Aquaticus.
+We find that the read embeddings form a low-dimensional manifold, which can be efficiently searched locally to obtain high-fidelity genome assembly from the simulated reads.
+Our future work on genome assembly will focus on improving RDE models to obtain better embeddings, and also on developing local-search based algorithms that can accurately map low-dimensional manifolds (comprised by the embeddings of reads) into linear DNA sequences.

### v6

@@
 Abstract
-DNA sequence alignment involves assigning short DNA reads to the most probable locations on an extensive reference genome.
-This process is crucial for various genomic analyses, including variant calling, transcriptomics, and epigenomics.
-Conventional methods, refined over decades, tackle this challenge in two steps: genome indexing followed by efficient search to locate likely positions for given reads.
-Building on the success of Large Language Models (LLM) in encoding text into embeddings, where the distance metric captures semantic similarity, recent efforts have explored whether the same Transformer architecture can produce numerical representations for DNA sequences.
-Such models have shown early promise in tasks involving classification of short DNA sequences, such as the detection of coding- vs non-coding regions, as well as the identification of enhancer and promoter sequences.
-Performance at sequence classification tasks does not, however, translate to sequence alignment, where it is necessary to conduct a genome-wide search to align every read successfully, a significantly longer-range task by comparison.
-We bridge this gap by framing the Sequence Alignment task for Transformer models as an “Embed-Search-Align” task.
-In this framework, a novel Reference-Free DNA Embedding RDE model generates representations of reads and fragments of the reference, which are projected into a shared vector space where the read-fragment distance is used as a surrogate for alignment.
-In particular, we introduce: (1) Contrastive loss for self-supervised training of DNA sequence representations, facilitating rich reference-free, sequence-level embeddings, and (2) a DNA vector store to enable search across fragments on a global scale.
+DNA sequence alignment, an important genomic task, involves assigning short DNA reads to the most probable locations on an extensive reference genome.
+Conventional methods tackle this challenge in two steps: genome indexing followed by efficient search to locate likely positions for given reads.
+Building on the success of Large Language Models (LLM) in encoding text into embeddings, where the distance metric captures semantic similarity, recent efforts have encoded DNA sequences into vectors using Transformers and have shown promising results in tasks involving classification of short DNA sequences.
+Performance at sequence classification tasks does not, however, guarantee sequence alignment, where it is necessary to conduct a genome-wide search to align every read successfully, a significantly longer-range task by comparison.
+We bridge this gap by developing a “Embed-Search-Align” (ESA) framework, where a novel Reference-Free DNA Embedding (RDE) Transformer model generates vector embeddings of reads and fragments of the reference in a shared vector space; read-fragment distance metric is then used as a surrogate for sequence similarity.
+ESA introduces: (1) Contrastive loss for self-supervised training of DNA sequence representations, facilitating rich reference-free, sequence-level embeddings, and (2) a DNA vector store to enable search across fragments on a global scale.
 RDE is 99% accurate when aligning 250-length reads onto a human reference genome of 3 gigabases (single-haploid), rivaling conventional algorithmic sequence alignment methods such as Bowtie and BWA-Mem.
-RDE far exceeds the performance of $6$ recent DNA-Transformer model baselines on ESA such as Nucleotide Transformer, Hyena-DNA, and shows task transfer across chromosomes and species.
+RDE far exceeds the performance of $6$ recent DNA-Transformer model baselines such as Nucleotide Transformer, Hyena-DNA, and shows task transfer across chromosomes and species.
 keywords:
@@
 In order to interpret this data, a typical first step is to align the reads to a genome.
-Genomes come in many sizes, but many of the most commonly studied genomes, such as that of humans, are three gigabases long.
-Moreover, most experiments will generate millions of reads.
-The resulting task of aligning short reads to large-size genomes is computationally challenging –akin to finding a pin in a haystack– and many decades of work have led to optimized approaches that can perform these tasks with great efficiency.
+Genomes come in many sizes, and commonly studied genomes range from millions of bases (e.g., bacteria) to billions of bases (e.g., mammals, plants) in length.
+The resulting task of aligning short reads to large-size genomes is computationally challenging—akin to finding a needle in a haystack—and many decades of work have led to optimized approaches that can perform these tasks with great efficiency.
 Since genomes are sequences where the alphabet consists of only a few symbols ($\{A,T,G,C\}$) they can be considered as Limited Alphabet Languages (LAL).
 We ask whether one can design a new paradigm to align reads to genomes by exploiting the Transformer architectures vaswani_attention_2017 adopted for recent language modeling efforts.
-The applications of the Transformer model – and, more generally “Large Language Models” (LLM) – in Bioinformatics applications are still in their infancy, yet hold substantial promise: Transformer models have demonstrated an unprecedented ability to construct powerful numerical representations of sequential data vit ; wav2vec ; gpt3 .
-Numerous DNA Transformer models (ji_dnabert_2021, ; zvyagin_genslms_2022, ; fishman_gena-lm_2023, ; nucleotide, ) have emerged recently, mainly designed for classification tasks in downstream applications.
-However, these models do not explicitly consider a fundamental distinction between Limited Alphabet Languages (LAL) (such as genomes) and natural languages with large vocabulary.
-In LALs, there is a one-to-one correspondence between the precise symbol ordering and the underlying “meaning”.
-If a protein is encoded by one amino acid sequence, a slightly different sequence (with a few edits) would encode for a different protein with a distinct functionality in the cells.
-These models, trained on classification tasks, generate sequence embeddings such that their pairwise distances correspond to class separation.
-As a result, pairs of sequences with very large edit distances between them are mapped to numerical representations that are close by.
-However, in tasks such as Sequence Alignment the objective is quite different: the pairwise representation distance has to closely match the sequence edit distance.
-Indeed as portrayed in Figure 1, current DNA Transformer models without any further refinement are unable to perform Sequence Alignment tasks: the representation space is simply not designed for such an objective.
-This demonstrates the need for a new model which is able to transfer a precise and computationally expensive distance metric, such as the edit distance, over LALs to a continuous representation space.
-We establish a foundation model foundation Reference-free DNA Embedding (RDE) – tailored for embedding DNA sequences.
-The model $h_{\Theta}$ (i.e., parameterized by weights $\Theta$) maps any genome subsequence $F$ into an embedding $h_{\Theta}(F)\in\mathcal{R}^{n}$ (e.g., $n=768)$, such that if $F_{1}$ and $F_{2}$ are similar subsequences of differing lengths – for example, $F_{1}$ is any noisy subsequence of $F_{2}$ or $F_{1}$ and $F_{2}$ have significant overlaps – then the $d(h_{\Theta}(F_{1}),h_{\Theta}(F_{2}))$ is small; where $d(\cdot)$ is a distance metric in the embedding space, such as the cosine distance between two vectors.
+The applications of the Transformer model—and, more generally “Large Language Models” (LLM)—in Bioinformatics applications are still in their infancy, yet hold substantial promise: Transformer models have demonstrated an unprecedented ability to construct powerful numerical representations of sequential data vit ; wav2vec ; gpt3 .
+We establish a foundation model foundation –Reference-free DNA Embedding (RDE)—tailored for embedding DNA sequences.
+The model $h_{\Theta}$ (i.e., parameterized by weights $\Theta$) maps any genome subsequence $F$ into an embedding $h_{\Theta}(F)\in\mathcal{R}^{n}$ (e.g., $n=768)$, such that if $F_{1}$ and $F_{2}$ are similar subsequences of differing lengths—for example, $F_{1}$ is any noisy subsequence of $F_{2}$ or $F_{1}$ and $F_{2}$ have significant overlaps—then the $d(h_{\Theta}(F_{1}),h_{\Theta}(F_{2}))$ is small; where $d(\cdot)$ is a distance metric in the embedding space, such as the cosine distance between two vectors.
 Another requirement of such RDE models is that they should generate embeddings that are reference free, i.e.
@@
 Thus, a global search in a giga-bases long sequence is reduced to a local search in a vector space.
-The simplest sequence alignment task applies to single-end [^1] A DNA fragment is ligated to an adapter and then sequenced from one end only. reads, where a sequencer generates a read of length $Q$
-$r:=(\tilde{b}_{1},\tilde{b}_{2},\dots,\tilde{b}_{Q}),$ (1)
-where $\tilde{b}_{i}\in\{A,T,G,C\}$.
-In practice, these reads come from individual genomes that do not necessarily match the reference and may contain mutations due to base insertions, deletions, and substitutions.
-Thus, it is assumed that this read is a noisy substring taken from a reference genome sequence $\mathcal{R}:=(b_{1},b_{2},\dots,b_{N}),\,b_{i}\in\{A,T,G,C\}$ and $N\gg Q$; for example, for the single-haploid human genome (chm_consort, ), $N\approx 3$ gigabases (gb), and $Q\approx 250$ though reads of much longer lengths are becoming increasingly affordable and accurate giab .
-The alignment task is to find a substring in $\mathcal{R}$
-[…]
-such that the edit distance $d(r,\tilde{r})$ – for example the Smith-Waterman (SW) distance – is minimized.
-The primary objective is to identify the most probable location, $q$, of this read within the reference genome.
 Figure 1: Alignment Recall of Transformer-DNA Baselines by Read Length: Existing Transformer-DNA models were adapted for sequence alignment using mean-/max-pooling.
@@
 3.
-Computational simulators have been developed to generate synthetic reads that have properties of real reads.
-These simulators mimic the read quality and characteristics produced by actual sequencing machines, thus providing a scalable means for validating new alignment approaches (huang_art_2012, ).
 ((a)) Three $20,000$-long nucleotide sequences were selected, one each from Chr.
@@
 Additionally, almost all of the reads are close to their corresponding fragments, making this space viable for the task of Sequence Alignment: by searching for fragments in the neighborhood of an external read represented in the embedding space, one is likely to retrieve a fragment most responsible for the read.
-Conventional sequence alignment methods, such as those used in BWA-MEM, have evolved significantly from the basic SW approach, which computes the similarity between a read $r$ and reference $\mathcal{R}$; using the SW distance on its own, however, is computationally intensive for large genomes.
-Some key advances include: (a) Sharding the reference into smaller fragments for individual searching; (b) Progressive search using phylogenetic trees and distance heuristics for logarithmic complexity scaling; (c) Compression techniques like the Burrows-Wheeler transform (bw_transform, ), reducing the effective search length; and (d) Multi-core/thread implementations and database instantiations for faster computations and improved data recall (minimap, ; bwamem2, ; bowtie, ).
-In this work, we explore an alternative paradigm for aligning a read to a genome, drawing parallels with advancements in Natural Language Processing (NLP).
-Traditional NLP relied on rule-based techniques to diagram sentences into grammatical components, using dependency parse trees to discern semantic relationships.
-However, the introduction of models like the Transformers (vaswani_attention_2017, ) revolutionized this approach.
-Unlike rule-based frameworks, Transformers implicitly identify syntactic and semantic structures tenney , as evidenced by their performance at various NLP tasks like Sentiment Analysis, Entity Recognition, and Question-Answering (devlin_bert_2019, ; reimers_sentence-bert_2019, ; reviewllm, ).
-The flexibility of this architecture has facilitated its application beyond language processing, including in vision (dosovitskiy_image_2020), auditory (audiotransformer, ), and neurological domains (wang_brainbert_2023, ).
-In bioinformatics, researchers are now harnessing Transformers to bypass genome-specific hard-coded rules, a development we contribute to with our novel method for Sequence Alignment.
 1.1 1.1 Transformer models: Written Language to DNA Sequence Alignment
@@
 Traditional DNA modeling efforts have already accommodated mature encoding and hashing techniques initially developed for written language – such as Suffix trees/arrays and Huffman coding (huffman_method_1952, ; manber_suffix_1993, ) – to successfully parse and compress DNA sequences.
-Recognizing such similarities, within the last few years, several Transformer-based models have been developed for DNA sequence analysis.
+Alignment methods such as MinHash and MashMap mashmap have incorporated Locality Sensitive Hashing (LSH) lsh to construct fast, approximate representations of DNA sequences that facilitate the rapid matching of millions of reads onto the reference genome.
+Within the last few years, several Transformer-based models have been developed for DNA sequence analysis.
 Notably, DNABERT-2 (ji_dnabert_2021, ; dna2bert, ), Nucleotide Transformer (nucleotide, ), GenSLM (zvyagin_genslms_2022, ), HyenaDNA (nguyen_hyenadna_2023, ) and GENA-LM (fishman_gena-lm_2023, ) have been designed to discern relationships between short genetic fragments and their functions.
@@
 We leverage a specialized data structure, termed a DNA vector store, that enables efficient storage and search in an embedding space [^1] Codebase available here: https://anonymous.4open.science/r/dna2vec-7E4E/: the entire reference genome is sharded into equal-length overlapping fragments whose embeddings are then uploaded to the vector store.
-For each read, top-K closest fragments can then be searched efficiently in time that scales only logarithimically in the number of fragments.
-These strategies have been explored in NLP: (A) Sequence-to-embedding training using contrastive loss has shown improved performance – over explicit pooling methods – at abstractive semantic tasks such as evidence retrieval lewis and semantic text similarity (gao_simcse_2021, ; chen_simple_2020, ).
+For each read, top-K closest fragments can then be searched efficiently in time that scales only logarithmically in the number of fragments hnsw .
+These strategies have been explored in NLP: (A) Sequence-to-embedding training using contrastive loss has shown improved performance—over explicit pooling methods—at abstractive semantic tasks such as evidence retrieval lewis and semantic text similarity (gao_simcse_2021, ; chen_simple_2020, ).
 (B) Specialized data structures, such as “vector stores” or “vector databases” like FAISS (johnson_billion-scale_2019, ) and Pinecone, use advanced indexing and retrieval algorithms for scalable numerical representation search.
-2.1 2.1 RDE: Designing effective sequence representations
-An optimal sequence encoder model $h$ is such that the corresponding embeddings of any read $r$ and reference fragment $\mathcal{F}$$h(r),h(\mathcal{F})$ respectively – obey the following constraints over a pre-determined distance metric $d$:
-$d\{h(r_{j}),h(\mathcal{F}_{i})\}\geq d\{h(r_{j}),h(\mathcal{F}_{j})\},\quad i\neq j$ (3)
-Subscript $j$ serves to indicate that the read $r_{j}$ is correctly aligned to a fragment reference, $\mathcal{F}_{j}$, (that we call a positive {read, fragment} pair).
-Given another arbitrary fragment $\mathcal{F}_{i}$ to which $r_{j}$ has poor alignment, $\{r_{j},\mathcal{F}_{i}\}$ is a negative {read, fragment} pair.
-Observe that these inequalities constitute the only requirements for the encoder.
-As long as the neighborhood of $r_{j}$ in the representation space contains the representation for $\mathcal{F}_{j}$, it will be recovered in the nearest neighbors (top-$K$ set) and alignment will succeed.
-Equality is observed when $r_{j}$ is a repeat sequence matched equally well to more than one fragment.
-This motivates using self-supervision (hadsell_dimensionality_2006, ; chen_simple_2020, ; gao_simcse_2021, ) where we are only concerned about the relative distances between positive and negative (read, reference fragment) pairs.
-2.2 2.2 RDE: Self-supervision and contrastive loss
-A popular choice for sequence learning using self-supervision involves a contrastive loss setup described by  chen_simple_2020 and gao_simcse_2021 : i.e.
-for a read $r$ aligned to reference fragment $\mathcal{F}_{j}$, the loss $l_{r}$ simultaneously minimizes the distance of $h(r)$ to $h(\mathcal{F}_{j})$ and maximizes the distance to a batch of random fragments of size $B-1$:
-$l_{r}=-\log\frac{e^{-d(h(r),h(\mathcal{F}_{j}))/\tau}}{e^{-d(h(r),h(\mathcal{F}_{j}))/\tau}+\sum^{B-1}_{i=1}e^{-d(h(r),h(\mathcal{F}_{i}))/\tau}}.$ (4)
-Here $\tau$ is a tuneable temperature parameter.
-To stabilize the training procedure and reach a non-trivial solution, the encoder applies different dropout masks to the reads and fragments similar to the method described in prior work chen_simple_2020 .
-Similar setups have been shown to work in written language applications, most notably in Sentence Transformers (reimers_sentence-bert_2019, ; gao_simcse_2021, ; muennighoff_mteb_2023, ), which continue to be a strong benchmark for several downstream tasks requiring pre-trained sequence embeddings.
-2.3 2.3 RDE: Encoder implementation
-RDE uses a Transformer-encoder (devlin_bert_2019, ; vaswani_attention_2017, ), comprising $12$ heads and $6$-layers of encoder blocks.
-The size of the vocabulary is $10,000$.
-Batch size $B$ is set to $16$ with gradient accumulation across $16$ steps.
-Generated numerical representations for each read (and fragment) are projected into a high-dimensional vector space, $h(r)\in\mathbb{R}^{1020}$.
-The learning rate is annealed using one-cycle cosine annealing (smith_super-convergence_2019, ), dropout is set to $0.1$, and $\tau=0.05$.
-During training, the reference fragment $|\mathcal{F}_{i}|\sim\mathcal{U}([800,2000])$ and read $|r_{i}|\sim\mathcal{U}([150,500])$ have variable lengths sampled from a uniform distribution.
-Here, $|x|$ denotes the length of $x$.
-To improve model performance on noisy reads, $1-5\%$ of bases are replaced with another random base in $40\%$ of the reads in a batch.
-Shorter sequences were padded to equal the length of the longest sequence in a batch.
-The distance metric used is Cosine Similarity.
-2.4 2.4 ESA: Task Definition
+2.1 2.1 Task Definition
+The simplest sequence alignment task applies to single-end [^2] A DNA fragment is ligated to an adapter and then sequenced from one end only. reads, where a sequencer generates a read of length $Q$
+$r:=(\tilde{b}_{1},\tilde{b}_{2},\dots,\tilde{b}_{Q}),$ (1)
+where $\tilde{b}_{i}\in\{A,T,G,C\}$.
+In practice, these reads come from individual genomes that do not necessarily match the reference and may contain mutations due to base insertions, deletions, and substitutions.
+Thus, it is assumed that this read is a noisy substring taken from a reference genome sequence $\mathcal{R}:=(b_{1},b_{2},\dots,b_{N}),\,b_{i}\in\{A,T,G,C\}$ and $N\gg Q$; for example, for the single-haploid human genome (chm_consort, ), $N\approx 3$ gigabases (gb), and $Q\approx 250$ though reads of much longer lengths are becoming increasingly affordable and accurate giab .
+The alignment task is to find a substring in $\mathcal{R}$
+[…]
+such that the edit distance $d(r,\tilde{r})$—computed using the Smith-Waterman (SW) algorithm—is minimized.
+The primary objective is to identify the most probable location, $q$, of this read within the reference genome.
 We formulate the problem of Sequence Alignment as minimizing a sequence alignment function, SA, applied to a read $r$ and a reference sequence $\mathcal{R}$ as
@@
 After training, the genome is segmented into overlapping fragments, encoded, and uploaded into the vector store.
+2.2 2.2 RDE: Designing effective sequence representations
+An optimal sequence encoder model $h$ is such that the corresponding embeddings of any read $r$ and reference fragment $\mathcal{F}$$h(r),h(\mathcal{F})$ respectively—obey the following constraints over a pre-determined distance metric $d$:
+$d\{h(r_{j}),h(\mathcal{F}_{i})\}\geq d\{h(r_{j}),h(\mathcal{F}_{j})\},\quad i\neq j$ (5)
+Subscript $j$ serves to indicate that the read $r_{j}$ is correctly aligned to a fragment reference, $\mathcal{F}_{j}$, (that we call a positive {read, fragment} pair).
+Given another arbitrary fragment $\mathcal{F}_{i}$ to which $r_{j}$ has poor alignment, $\{r_{j},\mathcal{F}_{i}\}$ is a negative {read, fragment} pair.
+Observe that these inequalities constitute the only requirements for the encoder.
+As long as the neighborhood of $r_{j}$ in the representation space contains the representation for $\mathcal{F}_{j}$, it will be recovered in the nearest neighbors (top-$K$ set) and alignment will succeed.
+Equality is observed when $r_{j}$ is a repeat sequence matched equally well to more than one fragment.
+This motivates using self-supervision (hadsell_dimensionality_2006, ; chen_simple_2020, ; gao_simcse_2021, ) where we are only concerned about the relative distances between positive and negative (read, reference fragment) pairs.
+2.3 2.3 RDE: Self-supervision and contrastive loss
+A popular choice for sequence learning using self-supervision involves a contrastive loss setup described by  chen_simple_2020 and gao_simcse_2021 : i.e.
+for a read $r$ aligned to reference fragment $\mathcal{F}_{j}$, the loss $l_{r}$ simultaneously minimizes the distance of $h(r)$ to $h(\mathcal{F}_{j})$ and maximizes the distance to a batch of random fragments of size $B-1$:
+$l_{r}=-\log\frac{e^{-d(h(r),h(\mathcal{F}_{j}))/\tau}}{e^{-d(h(r),h(\mathcal{F}_{j}))/\tau}+\sum^{B-1}_{i=1}e^{-d(h(r),h(\mathcal{F}_{i}))/\tau}}.$ (6)
+Here $\tau$ is a tuneable temperature parameter.
+To stabilize the training procedure and reach a non-trivial solution, the encoder applies different dropout masks to the reads and fragments similar to the method described in prior work chen_simple_2020 .
+Similar setups have been shown to work in written language applications, most notably in Sentence Transformers (reimers_sentence-bert_2019, ; gao_simcse_2021, ; muennighoff_mteb_2023, ), which continue to be a strong benchmark for several downstream tasks requiring pre-trained sequence embeddings.
+2.4 2.4 RDE: Encoder implementation
+RDE uses a Transformer-encoder (devlin_bert_2019, ; vaswani_attention_2017, ), comprising $12$ heads and $6$-layers of encoder blocks.
+The size of the vocabulary is $10,000$.
+Batch size $B$ is set to $16$ with gradient accumulation across $16$ steps.
+Generated numerical representations for each read (and fragment) are projected into a high-dimensional vector space, $h(r)\in\mathbb{R}^{1020}$.
+The learning rate is annealed using one-cycle cosine annealing (smith_super-convergence_2019, ), dropout is set to $0.1$, and $\tau=0.05$.
+During training, the reference fragment $|\mathcal{F}_{i}|\sim\mathcal{U}([800,2000])$ and read $|r_{i}|\sim\mathcal{U}([150,500])$ have variable lengths sampled from a uniform distribution.
+Here, $|x|$ denotes the length of $x$.
+To improve model performance on noisy reads, $1-5\%$ of bases are replaced with another random base in $40\%$ of the reads in a batch.
+Shorter sequences were padded to equal the length of the longest sequence in a batch.
+The similarity measure used is Cosine Similarity.
 2.5 2.5 ESA: Search and retrieval
@@
 In our experiments with external read generators (huang_art_2012, ), $Q_{max}=250,|\mathcal{F}_{i}|=1250$.
-Each reference fragment is encoded using the trained RDE model, and the resulting sequence embeddings ($\in\mathbb{R}^{1020}$) – $3$M vectors for a reference of $3$B nucleotides – are inserted into a Pinecone database.
+Each reference fragment is encoded using the trained RDE model, and the resulting sequence embeddings ($\in\mathbb{R}^{1020}$) – $3$M vectors for a reference of $3$B nucleotides – are inserted into a Pinecone [^3] https://www.pinecone.io database.
 Once populated with all the fragments, we are ready to perform the alignment.
@@
 3 demonstrate that even for pure reads, baseline models do not generate adequate representations to perform sequence alignment.
-In this section, RDE and the two best baselines – DB2, mean and HD, max – are evaluated on ESA using reads generated from an external read simulator (ART) – see huang_art_2012 .
+In this section, RDE and the two best baselines—DB2, mean and HD, max—are evaluated on ESA using reads generated from an external read simulator (ART) – see huang_art_2012 .
 ART has served as a reliable benchmark for evaluating other contemporary alignment tools and provides controls to model mutations and variations common in reads generated by Illumina machines.
@@
 For all presented results, the settings are: match_score = -2, mismatch_penalty = +1, open_gap_penalty = +0.5, continue_gap_penalty = +0.1.
-After alignment, we get $q^{*}$ – see (7) – as the estimated location of a read in the genome.
+After alignment, we get $q^{*}$—see (7)—as the estimated location of a read in the genome.
 Let $\hat{q}^{*}$ be its true location.
@@
 We define a bound $d_{SW}$ for classifying whether a (read, retrieved fragment) pair is a successful alignment based on the SW score between the pair ($v^{*}$ – see Eq.
-6) and the optimal SW score as a particular read length $Q$:
+4) and the optimal SW score as a particular read length $Q$:
 $d_{SW}=\frac{v^{*}-mQ}{(n-m)Q},$
@@
 The validation of RDE models can be performed only through various genomic tasks.
-In terms of the task of alignment, we plan to explore the following: (i) The current off-the-shelf implementation has a speed of around 10K reads per minute; while the existing aligners can achieve close to 1M reads per minute by trading off alignment accuracy.
+In terms of the task of alignment, we plan to explore the following: (i) The current off-the-shelf implementation of our ESA algorithm has a speed of around 10K reads per minute.
+The existing aligners such as Bowtie are faster, and can achieve close to 1M reads per minute; additional speedups can be achieved by trading off alignment accuracy vargas .
 As detailed in SM Sec.
@@
 Detailed experiments showed that the performance of even this first-generation RDE methodology is comparable to that of traditional aligners, which have been refined over decades.
-Future differentiating application of RDE to the alignment task might lie in aligning reads to graph- or pan-genomes liao , where the reference genome has multiple variants.
-Indexing such pangenomes is a difficult task in traditional aligners; however, in ESA one can store the RDE embeddings of the different fragments in the pangenome, and then perform a near-neighbor search in the embedding space.
-Thus, the ESA methodology does not have to be significantly altered, and alignment of reads to pangenomes constitutes an important future task.
-Going forward we aim to answer the following important questions:
-  • How reference-free is our RDE model?
+Future application of RDE to the alignment task might lie in aligning reads to pan-genomes liao , where the diversity in the genomic content – modeled as allelic paths in a reference genome graph – can be exhaustively mapped onto the described embedding space (as in Fig.
+2).
+Conveniently, the proposed nearest-neighbor search would align a given read to fragments; the only difference being that multiple fragments can belong to the same reference location but different reference paths.
+The ESA methodology would not need to be modified beyond this.
+An important question we want to answer going forward is: How reference-free is our RDE model?
 To benchmark the generalizability of our RDE model we have performed several cross-species alignment tasks as summarized in Tables 3 and 4 in the SM.
 In particular, we have used an RDE trained using human genome to align reads to reference genomes belonging to a wide range of species.
-The results are encouraging and the design of universal RDEs –that would generate effective embeddings of subsequences sampled from the genome of every species – is another direction of future research.
-  • Is the RDE capable of generalizing to other tasks like de novo Genome Assembly?
+The results are encouraging and the design of universal RDEs—that would generate effective embeddings of subsequences sampled from the genome of every species—is another direction of future research.
 One can also harness the RDE space to address the complementary task of de novo Genome Assembly, where given a set of reads one assembles a genome by stitching together the reads, without using any reference.
@@
 In SM Sec.
-E, we provide initial proofpoints of a genome assembly framework, under reasonable simplifying assumptions about read lengths and overlaps among the reads.
-We show how the RDE model – trained on exemplars from human genome – can effectively embed simulated reads from a very different species, Thermus Aquaticus.
+F, we provide initial proofpoints of a genome assembly framework, under reasonable simplifying assumptions about read lengths and overlaps among the reads.
+We show how the RDE model—trained on exemplars from human genome—can effectively embed simulated reads from a very different species, Thermus Aquaticus.
 We find that the read embeddings form a low-dimensional manifold, which can be efficiently searched locally to obtain high-fidelity genome assembly from the simulated reads.